### v1

[…]
1 Introduction
Monte Carlo (MC) simulation is a foundational tool in computational finance and other domains where closed-form solutions for complex stochastic models are unavailable. Its principal virtues are unbiasedness and the ability to rigorously quantify uncertainty through confidence intervals. By relying on random sampling, MC provides well-understood statistical guarantees and error assessments. However, Monte Carlo methods face significant challenges when applied to complex, nested, or path-dependent settings. As Glasserman (2013a) notes,“The slow convergence of Monte Carlo - its $O(1/\sqrt{n})$ error after $n$ replications - is a persistent obstacle to its wider use.” In other words, while MC excels at delivering unbiased estimates with reliable uncertainty quantification, it often becomes a “slow evaluation”. For example, in the context of derivative pricing, the challenge is further compounded for path-dependent options. Hull (2018) points out that “for path-dependent options, the entire price path must be simulated, significantly increasing computational complexity.” Similarly, Broadie and Glasserman (1996) state that “in many financial simulations, the value at each time step is a function of the previous step, creating a sequential dependency that hinders straightforward parallelization.”
In recent years, machine learning (ML) prediction methods have emerged as a fast alternative for approximating expensive functions, including prices of options and of other financial derivatives (Ferguson and Liang (2018), Bayer and Stemper (2019)). Neural networks and other ML techniques can learn complex mappings quickly, and once trained, can generate predictions at negligible marginal cost - what practitioners often call ”fast evaluation” (Horváth et al. (2021)). However, their black-box nature and lack of inherent statistical control make it difficult to dissect error and impossible to ensure unbiasedness Hutchinson et al. (1994). As a result, pure ML approaches often yield biased estimates with no guaranteed measure of error, making them potentially unsuitable for high-stakes applications like risk management where reliable uncertainty quantification is crucial Buehler et al. (2019).
This tension between MC’s reliability and ML’s efficiency motivates the exploration of classical variance reduction techniques, particularly control variates (CV). Control variates provide a systematic approach to improving MC efficiency by exploiting auxiliary functions with known expectations ( Glasserman (2013b), Asmussen and Glynn (2007)). When one finds a quantity that is highly correlated with the target while also possessing an analytically known mean, then the variance reduction could be substantial (Boyle et al. (1997)). However, as in many path-dependent exotic derivatives, under many complex models — effective control variates are simply unavailable (see, for instance, Duffie and Glynn (1995)). As noted in Glasserman (2013b),“some exotic payoffs are so irregular that finding an effective control variate becomes a significant challenge in itself.”
Our work seeks to bridge these two approaches, offering a hybrid framework that leverages the best of both worlds. We propose a novel control variate (CV)-based methodology, termed Prediction Enhanced Monte Carlo (PEMC), which extends the classical MC framework by incorporating modern ML techniques while preserving the unbiasedness and rigorous error quantification of MC. In the context of option pricing, unlike traditional ML applications in pricing derivatives, where models are trained to directly predict option prices, we train a certain conditional expectation as CVs, where we break away from two major characteristics of traditional CVs:
  • Known mean requirement: Unlike traditional CVs that demand the auxiliary function have a known mean (a challenging condition for complex or irregular payoffs), our approach does not rely on this assumption. Instead, PEMC utilizes ML-optimized predictors that are simple to estimate and still provide effective variance reduction.
  • Per-run variance reduction: While traditional CV techniques are designed to reduce variance on a per-sample basis—pairing each sample with a corresponding control variate—PEMC takes a broader approach. It focuses on achieving variance reduction across the entire simulation scheme, rather than targeting individual sample variance reduction.
Through a careful combination of classical MC principles and ML innovation, our approach eliminates bias from the black-box nature of ML models while also enjoying the unbiasedness and variance reduction provided by the traditional CV methods. By combining these elements, PEMC provides a principled and scalable framework that addresses the shortcomings of both classical CVs and ML-based methods, achieving reliable and interpretable results even in high-dimensional or irregular settings.
1.1 Our Contribution: A Modernized View of Control Variates
We introduce Prediction-Enhanced Monte Carlo (PEMC), a novel framework for evaluating Monte Carlo, which can be viewed as a modernized view of control variates: rather than searching for a single closed-form control variate with a known mean, we train a flexible predictor that acts as a control variate over the entire scheme. Crucially, the construction maintains unbiasedness and allows for valid statistical inference—just as in MC—while leveraging the representational power of ML models to approximate conditional expectations that can be evaluated efficiently. The key innovation of PEMC lies in relaxing the classical requirement that control variates must have a known mean. This is achieved through a two-stage estimation process: the first stage uses a small number of full simulations to estimate the difference between the target and the predictor, while the second stage computes the predictor average using a large number of computationally inexpensive simulations.
This design makes the full simulation cost modest, despite harnessing a large number of cheap samples to drive down variance of the predictor. Consequently, we achieve a form of variance reduction not just on a per-path basis (as the traditional CV), but scheme-wide: the entire simulation procedure, not a given sample design, is optimized to reduce the effective scheme variance vs scheme cost tradeoff, not the per sample variance vs per sample cost tradeoff. This differs from the defining characteristic of traditional CV methods where one control variate is applied per run. Our PEMC approach modernizes this idea, offering a “scheme-wide” variance reduction that can adapt to complex models and broad ranges of parameters.
Compared to purely machine learning-based methods, PEMC transforms the role of ML in Monte Carlo tasks. Rather than relying on ML as a direct surrogate predictor for the quantity of interest—which often introduces biased approximations with unquantifiable errors—PEMC uses ML to construct a robust, unbiased estimator augmented by modernized control variates. This approach enables fast predictions while maintaining Monte Carlo as the de facto standard of evaluation.
1.2 Literature Review
The literature on enhancing MC with ML approximators is extensive, but existing ML-related methods for control variates and variance reduction often impose stringent conditions that limit their practicality. Techniques such as the reproducing Stein kernel approach (Oates et al. (2017), Li and Zhang (2023)), regularized least squares for CV construction (Portier and Segers (2018), South et al. (2022), Leluc et al. (2021)), and adaptive CV schemes (Henderson and Glynn (2002), Henderson and Simon (2004), Kim and Henderson (2007)), as well as $\mathscr{L}^{2}$ function approximation frameworks (Maire (2003)), can in theory deliver powerful improvements—including even “supercanonical” convergence rates. But these methods usually require strict assumptions to ensure that the control variate’s mean is known and that it can be easily integrated into the simulation. In many practical settings, these assumptions fail, thus limiting the application scope of these methods. Even though some approaches can theoretically achieve supercanonical convergence rates, their applicability is limited by these constraints.
In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively in scenarios where auxiliary means are unknown. Unlike the aforementioned methods, PEMC aligns more closely with the recently proposed Prediction-Powered Inference (PPI) approach (Angelopoulos et al. (2023)), particularly in its hybrid methodology that integrates valid, unbiased and classical statistical principles with modern machine learning (ML) techniques. This aligns with recent trends in the evaluation of generative models Zrnic and Candès (2024), such as large language models (LLMs), where fast and interpretable metrics are increasingly critical. In this context, PPI-type of approaches highlight the value of combining human annotations and automatic predictions to yield unbiased estimates of model performance, even in low-label settings. This innovation is particularly valuable in high-dimensional or irregular settings, such as LLM evaluation, where annotated data may be sparse, but unannotated or simulation-based samples are plentiful Boyeau et al. (2024), Eyre and Madras (2024).
Traditional variance reduction and numerical approximation techniques have a long and established history in computational finance and stochastic modeling. One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by Giles (2015). MLMC exploits a hierarchy of discretizations (e.g., finer and coarser time steps in an SDE simulation) to reduce variance at a fraction of the cost of a naive Monte Carlo approach. Its basic premise involves coupling simulations at different resolution levels so that differences between levels have lower variance, ultimately leading to significant computational savings. Despite its solid theoretical foundations, MLMC can be challenging to implement in practice: constructing suitable couplings between the fine and coarse approximations, ensuring stability, and tuning parameters to achieve the promised complexity gains are all nontrivial tasks. However, adopting a coupling perspective provides valuable insights into how PEMC manages the relationship between the target and the predictor.
The advent of machine learning (ML) has spurred a wave of research aiming to accelerate and improve financial computation, including option pricing. Techniques such as neural network surrogates can approximate pricing functions or Greeks, drastically cutting down on runtime (Hutchinson et al. (1994), Buehler et al. (2019)). However, a major drawback is the lack of rigorous error controls and biases introduced during training.
Our paper is also related to the recent stream of literature in causal inference, where methods aim to mitigate challenges suc as selection bias and confounding. Doubly robust estimators serve as a key example (Bang and Robins (2005)). These estimators maintain consistency even if only one of two specified models—either the outcome model or the propensity score model—is correctly specified. Their inherent idea of robustness through additional predictor resonates with the philosophy behind control variates, as well as PEMC, where pairing a complex estimator with a suitably chosen secondary construct can yield improved stability and reduced variance.
The rest of the paper is organized as follows. In Section 2, we provide a comprehensive background on PEMC. In Section 3, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical illustrations. In Section 4, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework. We conclude the paper in Section 5 with discussion on broader implications and potential directions for future research.
2 Prediction-Enhanced Monte Carlo: Framework
In this section, we describe how to apply the “Prediction-Enhanced Monte Carlo” Framework for a general simulation evaluation. While this framework has broad applicability, for clarity and concreteness, we will present it in the context of options pricing. This choice of notation and presentation aligns with our subsequent examples and reflects a significant area of application. However, it’s important to note the principles and methodology we discuss here can be readily adapted to a wide range of problems beyond financial derivatives.
To effectively illustrate our proposed framework, facilitate a comprehensive comparison with traditional methods, and enhance the reader’s understanding, we employ Asian option pricing as a running example throughout in this section. Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period. As mentioned in Boyle et al. (1997) and Dingec et al. (2015), closed-form solutions for the pricing of arithmetic average Asian options are generally not available, but arithmetic average Asian options serve as an exemplary case where control variate methods, based on closed-form pricing of the geometric average Asian option under well-known models, can drastically reduce variance and enhance the efficiency of Monte Carlo simulations. However, as we will demonstrate, our new framework achieves the same level of variance reduction without requiring explicit knowledge of the construction of a control variate or its closed-form mean. This significantly expands the applicability and effectiveness of control variate techniques in arbitrarily complex models.
2.1 Problem Setup
The quantity of interest is an evaluation that can be expressed via:
Option $\text{Option Price}=\mathbb{E}_{\text{general risk neutral measure}}[f_{\text{payoff}}(\bm{Y})].$ (1)
Here $\bm{Y}$ represents a potentially high-dimensional random quantity that encapsulated all the randomness generated under a general, accessible sampling mechanism, denoted here by a general risk neutral measure $\mathbb{E}$. For example, if one is required to generate a discretized sample path of a certain stochastic differential equation (SDE) using some numerical scheme, $\bm{Y}$ then represents a complete sample path generated and the measure induced by that discretized SDE would be the general risk neutral measure. Typically, the more complex the simulation scheme is, the higher the dimension of $\bm{Y}$ will be, as more sophisticated simulation scheme often requires more intermediate steps or additional variables (e.g., an order of $O(T/\Delta t)$, when $T$ is the time horizon and $\Delta t$ is the discretization time step). Then, $f_{\text{payoff}}$, denoted by the term payoff function, represents the proper function (functional) that transform or summarize $\bm{Y}$ into a sample for the quantity of interest. The motivation for our framework arises, when the complexity and generality of the payoff function $f_{\text{payoff}}$ leads to a highly complicated dependence on the entirety of the random variables $\bm{Y}$. This complication may stem from factors such as path dependency, nested or multilevel simulations, or intricate interactions within the model. Consequently, drawing valid samples from the distribution of $f_{\text{payoff}}(\bm{Y})$ becomes both challenging and computationally intensive.
To fully specify the framework, we need to define the several parameters governing both the risk-neutral measure and the payoff function. We formally list the key definitions as:
  1. 1.
    Model Parameters $\bm{\theta}_{\text{model}}$: These are the parameters that specify the simulation or the stochastic model of the underlying process $\bm{Y}$ (e.g., asset value time series). For instance, in a Heston model (Heston 1993), the price process $S_{t}$ and the instantaneous volatility process $\nu_{t}$ are jointly modeled:
    $dS_{t}=$ $rS_{t}dt+\sqrt{\nu}_{t}S_{t}dW_{t}^{S}$
    $d{\nu}_{t}=$ $\kappa(\eta-\nu_{t})dt+\delta\sqrt{\nu_{t}}dW_{t}^{\nu}$ (2)
    The model parameter $\bm{\theta}_{\text{model}}$ is thus 5 dimensional $\bm{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. It specifies the risk neutral drift $r$, the long-run average variance $\eta$, the volatility of volatility $\delta$, the correlation $\rho$ of between $W^{S}$ and $W^{\nu}$, and the mean reversion rate $\kappa$. In options pricing, these parameters are typically calibrated to fit market data and then used to generate sample paths $\bm{Y}:=\{S_{t}\}_{t}$. Thus $\bm{\theta}_{\text{model}}$ governs and describes $\mathbb{E}_{\text{ risk neutral measure}}$ that generates $\bm{Y}$.
  2. 2.
    Simulation Parameters $\bm{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$. The distinction between $\bm{\theta}_{\text{simulation}}$ and $\bm{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are not calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed). For example, when one simulates the Heston model above using Euler discretization scheme, [^1] As mentioned in Dingec et al. (2015), Euler scheme was the primary method for simulation of Heston model until the exact simulation of Broadie and Kaya (2006) is developed. However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential (Andersen 2008) or Euler scheme remains prevalent (see, e.g., Lord et al. (2010)). we use $\bm{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$. The choice of these parameters can impact both the accuracy and efficiency of the simulation. While not typically part of the model calibration process, these parameters also play a crucial role in describing risk-neutral measure.
  3. 3.
    Payoff Function Parameters $\bm{\theta_{\text{payoff}}}$: This parameter $\bm{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\bm{\theta}_{\text{payoff}}}$. In options pricing, this specifies the details of the contract. For example, given the full price path $\bm{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
    $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
    where $\bm{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. Notice we have simply omitted the discount factor $e^{-rT}$.
Under these definitions, we can summarize the standard Monte Carlo approach as follows: one calibrates $\bm{\theta}_{\text{model}}$, specify $\bm{\theta}_{\text{simulation}}$ and generate $n$ independent paths $\{\bm{Y}_{i}\}_{i\in[n]}$ according to $\mathbb{E}_{\text{ risk neutral measure}}$ given by $\bm{\theta}_{\text{model}}$ and $\bm{\theta}_{\text{simulation}}$. Finally, one averages $\frac{1}{n}\sum_{i}f_{\bm{\theta}_{\text{payoff}}}(\bm{Y}_{i})$.
Remark Having specified the distinct roles of $\bm{\theta}_{\text{model}}$, $\bm{\theta}_{\text{simulation}}$ and $\bm{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\bm{\theta}$ when no ambiguity arises.
2.2 Prediction-Enhanced Monte Carlo: Motivation
As we have seen, Asian options are path-dependent derivative securities whose payoffs depend on the average prices of the underlying asset. Closed-form solutions for their prices are generally unavailable, numerous studies have proposed new numerical methods for pricing them (see Dingec et al. (2015)). While Partial Differential Equation (PDE)-based methods and Monte Carlo simulation are both generic methods for pricing options, applying PDE methods to particular models often requires special effort, primarily due to the curse of dimensionality. Moreover, for our discussion here, it has been noted that no PDE method performs well for Asian options under the Heston model Dingec et al. (2015). Therefore, across various models, simulation becomes the common choice. The principal drawback of simulation, as strongly highlighted in the literature, is its slow convergence rate. Employing variance reduction methods is therefore essential to improve the precision of simulation results. However, designing a successful variance reduction method requires exploiting the special structures of the derivative security and the underlying model. As a result, control variate (CV) techniques are often not available for pricing options under most of the complicated models used in practice. In fact, the pricing of arithmetic Asian option under Black–Scholes geometric Brownian motion (GBM) model is a main exception where the closed-form expression of the geometric Asian option allows one to effectively deploy control variate techniques. However, as soon as one extends beyond the most basic GBM model to more complex models such as the Heston model, the closed-form solutions for the geometric Asian options are no longer available and the applicability of control variate in these contexts is severely limited (see Dingec et al. (2015)).
More generally in the context of options pricing, while widely used, Monte Carlo methods face challenges when applied to complex, path-dependent derivatives. In such simulations, the state at each time step is determined by the state at the previous step, creating a sequential dependency that precludes straightforward parallelization or speedup. Moreover, for the complex risk neutral measures or the irregular payoff functions encountered in many exotic options pricing problems, effective control variate or variance reduction techniques are frequently unavailable. As a result, the slow convergence of Monte Carlo - its $O(1/\sqrt{n})$ error after $n$ replications - becomes an acute bottleneck and achieving high accuracy often implies a undesirable large number of replications in practice.
To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework is designed to provide a robust and easily accessible mechanism for designing control variates. This novel approach aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable. Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that Glasserman (2013b) highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
2.3 Prediction-Enhanced Monte Carlo: Prediction Model
In this section, we present our PEMC framework, which is centered around an efficient prediction model. In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is essentially compatible with any black-box model. However, as we will demonstrate, the squared error loss—a function that is both convex and differentiable—combined with the need for a flexible model family capable of expressing complex dependencies and being straightforward to train using off-the-shelf, easily accessible solvers, makes neural networks an ideal choice.
When applying the PEMC framework, an efficient prediction model refers to one that can process highly parallelizable and computationally efficient input $\bm{X}$, which serves as a strong predictor or good feature Hastie et al. (2009) for predicting $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$. In particular, given the whole path $\bm{Y}$, we want to find some input feature $\bm{X}:=\phi(\bm{Y})$ that is some typically low-dimensional transformation $\phi$ of $\bm{Y}$ with:
  • Property 1: $\bm{X}$ can be regarded as a reasonable predictor for $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$. For example, if $\bm{X}$ exhibits a significant correlation with $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$, it may serve as the basis for a reasonable linear regressor, although our framework far exceeds linear predictors.
  • Property 2: While the the joint distribution of $d$ is hard to obtain, the marginal distribution of $\bm{X}$ allows for efficient and parallelizable simulation.
The selection of an appropriate $\bm{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks. This process necessitates domain expertise and a thoughtful evaluation of the specific financial instrument and underlying model. The challenge lies in striking a balance between the predictive capability of $\bm{X}$ for $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$ and its computational efficiency. In subsequent sections, we will explore the intricacies of this selection process in greater detail, examining its impact on the performance of the PEMC framework through illustrative examples. For now, we proceed under the assumption that a suitable $\bm{X}$ has been identified.
We now focus on the pre-training phase of our prediction model. Unlike many machine learning or deep learning tasks, where the quantity and quality of data are critical challenges—and often a core difficulty due to the expense or scarcity of labeled data—in the Monte Carlo setting, we benefit from an effectively unlimited supply of clean, labeled data that can be generated as needed. The data generation and model training process can be outlined through the following key steps:
  1. 1.
    We first define a parameter space $\bm{\Theta}$ that encompasses all combinations of $\bm{\theta}:=(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}})$ that are of practical interest. This space should cover the range of parameters that would likely be encountered in real-world pricing scenarios on a daily, weekly or monthly basis, depending on the model update frequency.
  2. 2.
    We then draw samples of $\bm{\theta}_{i}$ from $\bm{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage. For each sampled parameter set, we generate one pair  [^2] You could also generate multiple pair of $(\bm{Y}(\bm{\theta}_{i}),\bm{X}(\bm{\theta}_{i}))$ under the same sampled $\bm{\theta}_{i}$. of $\bm{Y}(\bm{\theta}_{i})$ and its corresponding $\bm{X}(\bm{\theta}_{i}):=\phi(\bm{Y}(\bm{\theta}_{i}))$. It’s important to note that $\bm{X}$ and $\bm{Y}$ are coupled in this generation process.
  3. 3.
    For each generated pair, we compute the payoff $f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$, which serves as a label in our training data. The corresponding feature vector is comprised of $(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$.
  4. 4.
    This process is repeated $N_{\text{train}}$ times to generate training dataset which consists of $N_{\text{train}}$ pairs of
    $\textit{label}_{i}:=$ $(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$
    $\textit{feature}_{i}:=$ $f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$
    for $i\in[N_{\text{train}}]$.
  5. 5.
    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and save the resulting model ${g}$:
    $\min_{g}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-g(\text{feature}_{i}))^{2}.$ (3)
1:procedure DataGeneration($N_{\text{train}},\bm{\Theta}$)
[…]
3:     for $i=1$ to $N_{\text{train}}$ do
4:         Sample parameters $\bm{\theta}_{i}\sim\Theta$ uniformly
5:         Generate $\bm{X}(\bm{\theta}_{i}),\bm{Y}(\bm{\theta}_{i})\sim\mathbb{E}_{\text{ risk neutral measure}(\bm{\theta}_{i})}$
6:         $\textit{label}_{i}\leftarrow f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$
7:         $\textit{feature}_{i}\leftarrow(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$
[…]
9:     end for
[…]
11:end procedure
12:
[…]
14:     Initialize model $g$
15:     $N_{\text{train}}\leftarrow\text{length}(\textit{features})$
16:     Minimize: $\min_{g}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-g(\text{feature}_{i}))^{2}$
17:     return trained model $g$
18:end procedure
Algorithm 1 Prediction Model Training
The essential steps are summarized in Algorithm 1. While the data generation process can be time-intensive due to the complexity of $\bm{Y}$, it is conducted offline during the pre-training phase. This approach minimizes its impact on real-time pricing applications by effectively trading memory usage for computational speed during execution.
This pre-training setup presents an ideal scenario for machine learning applications. Unlike many real-world problems where data scarcity is a concern, access to the data-generating process in PEMC allows for the creation of virtually unlimited, clean data. This abundance greatly facilitates the training of complex models. It is important to also note that PEMC does not require $g$ to take specific forms. However, the convex and differentiable nature of the loss function (i.e., squared loss), the large supply of data (i.e., from Monte Carlo simulation), and the need for expressiveness of the model family make neural networks (NNs) an ideal choice.
When training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer Kingma and Ba (2014), batch normalization Ioffe and Szegedy (2015), and skip connections He et al. (2016)—can be employed to improve performance. In fact, recent advancements in machine learning technologies have made the training process remarkably accessible and efficient. The availability of open-source, user-friendly frameworks like PyTorch Paszke et al. (2019), TensorFlow Abadi et al. (2016), and JAX Bradbury et al. (2018)—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets. These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) LeCun et al. (1998), Residual Networks (ResNets) He et al. (2016), and Transformers Vaswani et al. (2017), as well as various Stochastic Gradient Descent (SGD) optimization algorithms Bottou (2010). In applications, we only utilized well-established and widely recognized neural network structures that are now considered common. These models can typically be implemented using just a few lines of code, as they have been incorporated into existing packages and extensively optimized.
Finally, regarding the choice loss function, we note that the goal of the prediction model $g$ is to produce an approximation of the underlying regression function:
$g(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X}(\bm{\theta}))$
$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}f_{\bm{\theta}_{\text{payoff}}}(\bm{Y}(\bm{\theta}))\bigg{|}(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X}(\bm{\theta}))\bigg{]}$
As we will see, the quality of this approximation will impact the effectiveness of variance reduction, which we shall explore in detail in subsequent sections.
2.4 Prediction-Enhanced Monte Carlo: Evaluation
Having completed the training phase of PEMC, we now turn our attention to the evaluation procedure. This stage employs the pre-trained model to enhance the efficiency of Monte Carlo simulations, and we will demonstrate how control variates are cleverly constructed to achieve variance reduction. The evaluation process proceeds as follows:
  1. 1.
    A specific parameter set $\bm{\theta}:=(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}})\in\bm{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing. These parameters can be determined through combination of calibration, customization, client agreement or external observation.
  2. 2.
    Generate $n$ pairs of $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$. Here we have suppressed $\bm{Y}$’s dependency on $\bm{\theta}$, as in the evaluation stage only a fixed $\bm{\theta}$ is considered.
  3. 3.
    Based on the marginal distribution of $\bm{X}$, independently generate $N$ additional samples of $(\bm{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$, directly according to its marginal distribution. Importantly, these samples are independent of the previous data $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$.
  4. 4.
    Utilize the pre-trained model ${g}$ to evaluate the PEMC estimator:
    $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$ (4)
For notational simplicity, we have written ${g}(\bm{X})$ instead of ${g}(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X})$, and $f(\bm{Y})$ instead of $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$. A key aspect of this procedure is the relationship between $N$ and $n$. Typically, we choose $N$ to be one, or several orders of magnitude larger than $n$. The optimal choice of $N$ versus $n$, will be discussed Section 3. The PEMC evaluation step is summarized in Algorithm 2.
1:procedure Evaluation($\bm{\theta},n,N,g$)
2:     Generate i.i.d. $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$ from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$
3:     Generate i.i.d. $(\bm{\tilde{X}}_{j})_{j\in[N]}$ independently from its marginal in $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$
4:     Compute $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$
5:     return $PEMC$
6:end procedure
Algorithm 2 Evaluation
2.5 PEMC: An Example on Asian Options
To demonstrate the PEMC framework, we return to our ongoing example on arithmetic Asian call option under the Heston model. It is worth noting that while the geometric Asian option is a natural control variate in this context, its closed-form expression is not available under the Heston model. To use control variates here, one must resort to sophisticated methods, e.g, numerically approximating the characteristic function Fusai and Meucci (2008) or employing new framewrok for conditional Monte Carlo Dingec et al. (2015).
This section serves as a user manual, guiding readers through the process of implementing PEMC for pricing an exotic option. When pricing an arithmetic Asian call option under the Heston model, the option payoff is given by:
$P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
where $\bm{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. The Heston model is described in (1) which requires $\bm{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. For this example, let’s say we use a neural network (NN) as the prediction model. Then, the PEMC framework could proceed as follows:
  1. 1.
    Define the parameter space $\bm{\Theta}$ that encompasses realizations of
    $\bm{\theta}:=(\underbrace{r,\eta,\delta,\rho,\kappa}_{\bm{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0},\Delta t,T}_{\bm{\theta}_{\text{simulation}}},\underbrace{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}}_{\bm{\theta}_{\text{payoff}}})$
    which are of practical interest. The choice of $\bm{\Theta}$ also depends on the model capacity and the model updating frequency. We shall discuss this in subsequent sections.
  2. 2.
    Uniformly sample $\bm{\theta}$ from $\bm{\Theta}$. This step is straightforward if the parameter space is a Cartesian product of intervals. For each sampled $\bm{\theta}$, generate process $\bm{Y}:=(S_{t},\nu_{t})_{t}$ using a Euler scheme with step size $\Delta t$, according to the Heston model specified by $\bm{\theta}$.
  3. 3.
    During the sampling of $\bm{Y}$, we collect $\bm{X}$ to be the sum of Brownian increment
    $W_{T}^{S}:=\sum_{j=1}^{T/\Delta t}\Delta W^{S}_{j\Delta t}\quad\text{ and }\quad W_{T}^{\nu}:=\sum_{j=1}^{T/\Delta t}\Delta W^{\nu}_{j\Delta t},$
    simulated during each step of the Euler scheme. This makes the marginal of $\bm{X}$ simply a two-dimensional Gaussian: $\frac{\bm{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
  4. 4.
    Save
[…]
  5. 5.
    Repeat steps 2-4 $N_{\text{train}}$ times to generate dataset $(\textit{feature}_{i},\textit{label}_{i})_{i\in[N_{\text{train}}]}$ of size $N_{\text{train}}$.
  6. 6.
    Train a NN with weights $\bm{w}$ to minimize the MSE loss:
    $\min_{{\bm{w}}}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-NN_{\bm{w}}(\text{feature}_{i}))^{2}$
    and use $\hat{w}$ to approximate
    $NN_{\bm{\hat{w}}}\approx\mathbb{E}\bigg{[}(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K)^{+}\bigg{|}{r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu}\bigg{]}.$
  7. 7.
    At evaluation, a specific $\bm{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4). We also generate $N$ i.i.d. $\sqrt{T}\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$ samples and store them as $(\tilde{\textit{feature}})_{j\in[N]}$.
  8. 8.
    Output $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(\textit{label}_{i}-NN_{\bm{\hat{w}}}(\textit{feature}_{i}))+\frac{1}{N}\sum_{j=1}^{N}NN_{\bm{\hat{w}}}(\tilde{\textit{feature}}_{j})$
2.6 Implementation Details in PEMC
Having walked through the Asian option example under the PEMC framework, we discuss several important implementation details. These considerations are critical for successfully applying PEMC in practice.
2.6.1 Parameter Space $\Theta$
When determining the parameter space $\bm{\Theta}$, a key consideration is how frequently the model will be updated. To ensure the performance of PEMC, it is desirable that the training data encompass all practical scenarios the model is expected to encounter within the designated update period. For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$. If one updates the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses about the bounds within which these parameters are likely to fluctuate over the coming month. For example, one could postulate the underlying price would mostly be fluctuating within a ±$\textpm 5-6\%$ range, i.e., around 490-550, based on the historical volatility of SPY. This expectation specifies an interval for $S_{0}$ in $\bm{\Theta}$. Similarly, one might set intervals for each parameter in the same manner, e.g., interval for $r\in(4.0\%,5.0\%)$, anticipating that the interest rates change would not exceed half a percentage point. Finally, we note that while these parameters are jointly distributed and could be calibrated together—especially considering that many market conditions are intertwined and correlated—this is often unnecessary. Instead, one can define intervals for each parameter individually and enlarge the uncertainty set, since the coverage of practical scenarios is more important in PEMC.
This process is conceptually similar to the selection of uncertainty sets in robust optimization Ben-Tal et al. (2008) or distributionally robust optimization Esfahani and Kuhn (2018), where the goal is to include realizations of parameters with high likelihood. However, unlike robust optimization approaches, which often favor data-dependent uncertainty sets, our focus here is also guided by the expertise of financial engineers. The training parameter space is designed to reflect realistic and practical scenarios derived from domain knowledge, rather than being strictly driven by statistical guarantees. In applications, we construct reasonable $\bm{\Theta}$ for practical models, such as forward curves in HJM modeling Glasserman (2013b) or local stochastic volatility 2D grids Gatheral (2006), among others.
2.6.2 The Design of $\Large{X}$: Feature Engineering
In our Asian option example, we designed the feature $\bm{X}$ as the sum of two correlated Brownian increments. As discussed in Section 2.3, the choice was motivated by two key properties:
  • Predictability: The sum of Brownian motion increments explains away a large portion of the variance driving the movement of SDE.
  • Ease of Generation: The sum of Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
Given these reasons, it is natural to use such sums as $\bm{X}$ for all related SDE driving by Brownian motion. Indeed, this is the approach we adopt in our applications, and it has proven to be both robust and effective. For cases where the path is very long, one can divide the sum into chunks of smaller increments. However, we found that using just one increment typically gives sufficient variance reduction.
However, the choice of Brownian increments, while effective, is certainly not the only one. As with most ML tasks, feature engineering that leverages domain expertise can yield better results. We illustrate with one such example, the pricing of a floating strike lookback call option, under the Heston model(1). The option payoff is given by:
$f(\{S_{t}\}_{t})=\max(S_{T}-\min_{0\leq t\leq T}S_{t},0)$
where $S_{t}$ is the asset price process and $T$ is the option maturity. A good choice of feature could be:
$\bm{X}=(W_{T}^{S},\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}).$
The rationale behind the construction of $\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}$ term is grounded in the domain knowledge of the Heston model. Specifically, one who is familiar with the role of $\eta$ as the mean-reverting level of variance process $\nu_{t}$ may approximate $\nu_{t}\approx\eta$ (i.e., assume fixed volatility) and simplify the asset price process to a geometric Brownian motion:
$S_{t}\approx S_{0}e^{(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}\text{ and }\min_{0\leq t\leq T}S_{t}\approx S_{0}e^{\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}.$
This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\bm{X}]_{1}}-S_{0}e^{[\bm{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$. This corresponds directly to Property 1 in the construction of $\bm{X}$ in 2.3. For Property 2, the marginal distribution of $\bm{X}$ permits efficient and parallelizable simulation. In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\bm{X}]_{1},[\bm{X}]_{2})$ Karatzas and Shreve (1991). This result makes it straightforward to sample these components directly.
2.6.3 Eliminating Data Storage Overheads
In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively. By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage. This difference stems from two main considerations.
First, the volume of training data required to achieve a well-performing prediction model can be large. Storing all of it would be both expensive and unnecessary. In PEMC, the training data can be produced directly via Monte Carlo simulation, ensuring an effectively unlimited supply. Second, this flexibility allows for a more efficient workflow. Data can be generated “on the fly” and processed in streaming fashion. For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training. In this manner, we have successfully trained models in the applications using data on the order of $10^{7}–10^{8}$.
3 Analysis and Performance
In this section, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical considerations. All properties of PEMC discussed here pertain to the evaluation stage, where $\bm{\theta}$ is fixed. Henceforth, we omit the explicit dependence on $\bm{\theta}$ in the notation (e.g., writing $\bm{Y}$ in place of $\bm{Y}(\bm{\theta})$). We begin by establishing the unbiasedness of our estimator.
3.1 Bias Analysis
Theorem 3.1 (Unbiasedness)
The PEMC estimator in equation (4), is unbiased:
$\mathbb{E}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})\bigg{]}=\mathbb{E}[f(\bm{Y})]$
Proof 3.2
[…]
Given the unbiased nature of PEMC, the analysis of PEMC as a Monte Carlo method primarily hinges on its variances and computational costs.
3.2 Variance Analysis
To facilitate the analysis, we first introduce some necessary notation.
Definition 3.3
Given function $f$ that takes input $\bm{Y}$ and $g$ that takes input $\bm{X}$ (so both $f,g$ are considered fixed here), we denote
$\sigma_{f-g}^{2}:=\mathrm{Var}(f(\bm{Y})-g(\bm{X})),\text{ }\sigma^{2}_{f}:=\mathrm{Var}(f(\bm{Y})),\text{ and }\sigma^{2}_{g}:=\mathrm{Var}(g(\bm{X})).$
Similarly, we denote cost of generating a coupled sample $f(\bm{Y})-g(\bm{X})$ as $c_{f-g}$, the cost of generating $f(\bm{Y})$ as $c_{f}$ and the cost of generating $g(\bm{X})$ as $c_{g}$.
Lemma 3.4
Using the notations from 3.3, we have
$\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))=\frac{\sigma^{2}_{f}+\sigma^{2}_{g}-\sigma^{2}_{f-g}}{2}.$
We make two remarks here. First, since we generally obtain $\bm{X}$ in the same process during the generation of $\bm{Y}$, thus typically in either sense of the cost, we have $c_{f-g}\approx c_{f}$. Second, the term cost is intentionally left ambiguous because its interpretation varies depending on the context. In some cases, cost refers to the sample size, while in others it represents the wall-clock time required to generate and evaluate the sample. Furthermore, in certain scenarios, both time and computational resources must be jointly considered and optimized, under parallelization and other speedup, making the concept of cost more instance dependent. Therefore, for now, we use cost as an umbrella term to capture various meanings under different contexts.
Lemma 3.5
The Variance of the PEMC estimator in equation (4) is
$\mathrm{Var}(PEMC)=\frac{1}{n}\sigma_{f-{g}}^{2}+\frac{1}{N}\sigma_{g}^{2}.$
Proof 3.6
[…]
Having established the unbiasedness and the variance form of the PEMC estimator, we could turn to inference. In particular, using consistent estimates of variance, the application of the central limit theorem enables the construction of valid asymptotic confidence intervals. Moreover, a range of inferential techniques may be employed—such as deriving high-probability bounds or other non-asymptotic guarantees—the asymptotic confidence interval construction suffices to illustrate the point. The approach follows closely from results in prediction-powered inference Angelopoulos et al. (2023).
Theorem 3.7 (Asymptotic Confidence Intervals)
In the set up of the evaluation algorithm 2, given the prediction model $g$, $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$ and $(\bm{\tilde{X}}_{j})_{j\in[N]}$, we define
$\hat{\sigma}^{2}_{f-g}=\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}f(\bm{Y}_{i})-g(\bm{X}_{i})-\Bigl{(}\frac{1}{n}\sum_{i^{\prime}=1}^{n}f(\bm{Y}_{i^{\prime}})-g(\bm{X}_{i^{\prime}})\Bigr{)}\Bigr{)}^{2},$
and
$\hat{\sigma}^{2}_{g}=\frac{1}{N}\sum_{j=1}^{N}\bigl{(}g(\bm{\tilde{X}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}g(\bm{\tilde{X}}_{j^{\prime}})\bigr{)}^{2},$
be the respective sample variance. Let $z_{1-\alpha/2}$ denote the $(1-\alpha/2)$-quantile of the standard normal distribution. Then, the interval
$\Bigg{(}PEMC-z_{1-\alpha/2}\sqrt{\frac{\hat{\sigma}^{2}_{f-g}}{n}+\frac{\hat{\sigma}^{2}_{g}}{N}},\quad PEMC+z_{1-\alpha/2}\sqrt{\frac{\hat{\sigma}^{2}_{f-g}}{n}+\frac{\hat{\sigma}^{2}_{g}}{N}}\text{ }\Bigg{)},$
is an asymptotically valid $1-\alpha$ confidence interval for $\mathbb{E}[f(\bm{Y})]$, as $n,N\to\infty$.
Proof 3.8
[…]
3.3 Variance Reduction
Having analyzed the bias and variance of the PEMC estimator, we now turn to a fundamental practical question: Under a fixed cost budget—whether measured in terms of the number of samples, computational time, or other limited resources—when does PEMC outperform standard Monte Carlo (MC)? Our aim is to understand how PEMC’s variance reduction scales with its cost, thereby identifying regimes where PEMC emerges as more efficient than standard MC. To this end, we first determine the optimal allocation of the sample sizes $n$ (the number of expensive full simulations) and $N$ (the number of cheap, feature-only evaluations) within the PEMC framework, given a total resource budget $B$. This investigation provides a guideline for parameter selection and reveals conditions under which PEMC’s gains are maximized.
Lemma 3.9
Assuming $n,N\in\mathbb{R}^{+}$ and $c_{g},c_{f-g}\in\mathbb{R}^{+}$, the optimal allocation between $n$ and $N$ for PEMC, for any positive budget $B$, follows as
$\frac{n}{N}=\frac{\sigma_{f-g}}{\sigma_{g}}\cdot\sqrt{\frac{c_{g}}{c_{f-g}}}\quad.$ (6)
Proof 3.10
Proof By relaxing the constraints to $n,N\in\mathbb{R}^{+}$, the optimization problem
$\underset{n>0,N>0}{\text{min}}\quad$ $\frac{1}{n}\sigma_{f-{g}}^{2}+\frac{1}{N}\sigma_{g}^{2}.$
s.t. $nc_{f-g}+Nc_{g}\leq B.$
[…]
In practice, $c_{f-g}\gg c_{g}$ often holds which suggests one should set $N\gg n$ according to Lemma 3.9. While Lemma 3.9 relies on treating $n,N\in\mathbb{R}^{+}$ as continuous—ignoring the integer constraints $n,N\in\mathbb{N}^{+}$— this continuous approximation still provides valuable guidance when $n,N\geq 1$. In fact, in practice, we found the ratio suggested by Lemma 3.9 guides near-optimal parameter selection. However, the standard MC corresponds to the case $N=0$, a scenario not covered by Lemma 3.9, and thus requires separate consideration. Building on Lemma 3.9, we can estimate an upper bound on the variance reduction PEMC can achieve relative to standard MC under these idealized assumptions.
Lemma 3.11
Assume $c_{f-g}=c_{f}$. In the same setup as Lemma 3.9, the variance ratio between PEMC under the optimal allocation and standard MC follows as
$\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}=\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}\bigg{(}1+\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{g}}{c_{f}}}\bigg{)}+\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}\bigg{(}\frac{\sigma_{f-g}}{\sigma_{g}}\cdot\sqrt{\frac{c_{g}}{c_{f}}}+\frac{c_{g}}{c_{f}}\bigg{)}.$ (7)
Proof 3.12
[…]
At this point, a natural question arises: how can we gauge $\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}$ or $\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}$ in practice? How are these ratios tied to the quality of the predictive model $g$ in PEMC? As we shall see, the extent of variance reduction that PEMC can deliver is linked to how well $g$ is trained. To illustrate this, we first consider an ideal scenario where $g$ is trained to the optimum, i.e., the true regression function.
Lemma 3.13
Suppose $c_{f-g}=c_{f}$, $f(\bm{Y})$ is square-integrable and $g=\mathbb{E}[f(\bm{Y})\,|\,\bm{X}]$. Define $\rho:=\mathrm{corr}(f,g)$ and $c:=\frac{c_{g}}{c_{f}}$. Then we have $\rho=\frac{\sigma_{g}}{\sigma_{f}}$ and the $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}$ in Lemma 3.11 becomes
$r(\rho,c):=$ $(1-\rho^{2})\bigg{(}1+\frac{\rho}{\sqrt{1-\rho^{2}}}\cdot\sqrt{c}\bigg{)}+\rho^{2}\bigg{(}\frac{\sqrt{1-\rho^{2}}}{\rho}\cdot\sqrt{c}+c\bigg{)}.$ (8)
Proof 3.14
Proof When $g=\mathbb{E}[f(\bm{Y})\,|\,\bm{X}]$ is the true regressor, i.e.,
[…]
Lemma 3.13 says that the variance reduction of PEMC relative to standard MC can be approximated by the function $r(\rho,c)$ in (8), where $\rho=\sigma_{g}/\sigma_{f}$ captures the predictive power of $g(\bm{X})$ for $f(\bm{Y})$, and $c=c_{g}/c_{f}$ quantifies the relative cost of evaluating $g(\bm{X})$ versus generating full samples for $f(\bm{Y})$. In Figure 3.3, we visualize the variance reduction function $r(\rho,c)$ for $\rho,c\in(0,1)$. On the left is a contour plot of $r(\rho,c)$ where the red line traces the curve $r(\rho,c)=1$, corresponding to the “break-even” boundary where PEMC’s variance matches that of MC. Regions below this line, $\{(\rho,r)\in[0,1]^{2}|r(\rho,c)<1\}$, represent regimes of $\rho$ and $c$ where PEMC achieves variance reduction, i.e., where $\rho$ is sufficiently large and $c$ is sufficiently small. On the right, the graph of $r(\rho,c)$ as a function of $\rho$ when $c=0.001$ provides a clear benchmark: a correlation $\rho=0.5$ yields approximately 22.2% variance reduction, while $\rho=0.7$ yields about 45.8%. In our PEMC applications, we design feature $\bm{X}$ and predictive model $g$ so that, for most $\bm{\theta}\in\Theta$ during evaluation, we can safely gauge that $c$ falls between $10^{-2}$ to $10^{-3}$ and $\rho$ exceeds 0.5. In these regions, Lemma 3.9 suggests choosing $N/n$ in the range of 5 to 20 for near-optimal performance. In our experiments, a ratio of $N=10n$ proved effective.
[…]
Variance Reduction Function $r(\rho,c)$. Left figure: A contour map of the variance reduction ratio $r(\rho,c)$ as a function of the correlation $\rho=\sigma_{g}/\sigma_{f}$ and relative cost $c=c_{g}/c_{f}$ The red curve indicates the curve $r(\rho,c)=1$ and the regimes where PEMC attains variance reduction over MC lie below the line. Right figure: A graph of $r(\rho,c)$ as a function of $\rho$ when $c=10^{-3}$.
3.3.1 Learning Theory Estimates
The statements in Lemma 3.9 assume that $g$ represents the true regression function $\mathbb{E}[f|\bm{X}]$. In practice, however, $g$ is obtained through a learning procedure applied to finite training samples, and thus will only approximate the true conditional expectation. Consequently, any practical implementation of PEMC must account for the approximation and estimation errors inherent in the training process. Moreover, as the parameter $\bm{\theta}$ varies, the correlation $\rho$ between $g(\bm{X})$ and $f(\bm{Y})$, the complexity of the underlying model and payoff function also changes.
To rigorously quantify these effects, one can invoke tools from statistical learning theory, such as uniform convergence bounds, VC-dimension, or Rademacher complexities (Vapnik 1998, Bartlett and Mendelson 2005, Mohri et al. 2018). These quantities, by relating the complexity of the function class used to represent $g$ and the available training sample size $N_{\text{train}}$, lead to error bounds that ensure, with high probability, that the trained predictor $g$ is close to the true regression function within a controlled margin. Although numerous, more refined approaches exist, we avoid delving deeply into them so as not to distract from our main theme. Instead, we adopt a standard, off-the-shelf, almost ‘placeholder’-type result detailed in the Appendix, that illustrates the typical form of guarantees one could expect from PEMC.
Lemma 3.15
Under the regularity conditions specified in the Appendix, for any $\epsilon,\delta>0$, there exists a sufficiently large sample size $N_{\text{train}}$ and a suitably chosen neural network class from which we select and train a predictor ${g}$ on $N_{\text{train}}$ samples, such that
$\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}\leq r(\rho,c)+\epsilon$
for a randomly evaluated $\theta\sim\Theta$ with probability at least $1-\delta$.
The simplified version captures the essential idea: under suitable regularity conditions, one can achieve the result of Lemma 3.9 within an arbitrarily small margin, as both the sample size $N_{\text{train}}$ and the complexity of the hypothesis class increase.
3.4 Control Variate Coefficient
In this section we explore a connection between the PEMC framework and traditional Control Variate (CV) methods. In (4), note that we could introduce a free parameter $a$ to create a variant:
$PEMC(a):=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-ag(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}ag(\bm{\tilde{X}}_{j})$ (9)
The introduction of $a$ is analogous to CV approaches, where the parameter $a$ is chosen to minimize variance. In PEMC, this viewpoint may appear somewhat redundant since $ag$ is also a valid prediction model, which suggests $a$ would be chosen implicitly during the training of the prediction model. Nevertheless, suppose $g$ is fixed. Then $\frac{1}{n}\sum_{i=1}^{n}-g(\bm{X}_{i})+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$ can be viewed as a readily available, zero-mean control variate. We can then choose $a$ which minimizes the variance. The optimal $a^{*}$ then is:
$a^{*}=\frac{\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))}{(n/N+1)\mathrm{Var}(g(\bm{X}))}.$
In the ideal scenario, we have $N\gg n$ and $g(\bm{X})=\mathbb{E}[f(\bm{Y})|\bm{X}]$, which would imply $\mathrm{Cov}(f(\bm{Y})-g(\bm{X}),g(\bm{X}))=0$ and $\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))=\mathrm{Var}(g(\bm{X}))$. This would further imply $a^{*}=\frac{1}{1+n/N}\approx 1$. Consequently, we can justify PEMC’s default choice of $a=1$, which saves us the trouble of estimating $a^{*}$, and gives PEMC a degree of robustness.
3.5 Variance Reduction of PEMC: An Example on Asian Options
We revisit the Asian option example and conduct experiments using the arithmetic Asian call option under the Geometric Brownian Motion (GBM) model. We choose the GBM model instead of the Heston model (1) to facilitate a direct comparison between PEMC, standard Monte Carlo (MC), and the traditional Control Variate (CV) method on the extent of variance reduction. This is because the geometric Asian option, which serves as a well-established control variate, is analytically tractable within the GBM framework but lacks a closed-form solution under the Heston model (Dingec et al. 2015).
This experiment aims to demonstrate the typical level of variance reduction that PEMC achieves. Our focus is strictly limited to variance reduction, because the GBM model actually allows for highly efficient and parallelized generation of both asset paths and corresponding payoffs, making the simulation of $\bm{Y}$ computationally straightforward. In other words, one could not fully leverage PEMC’s strengths in GBM model and this toy experiment only serves to provide insights into PEMC’s effectiveness in a controlled setting where a known CV exists. Practical applications and actual use-cases of PEMC will be provided in Section 4.
3.5.1 Experimental Setup
We implement the PEMC procedure detailed in Section 2.5, utilizing a neural network (NN) as the predictive model. The asset price follows the Geometric Brownian Motion (GBM) model:
$dS_{t}=$ $rS_{t}dt+\sigma S_{t}dW_{t}^{S}.$ (10)
The payoff calculation $P_{A}(\{S_{t}\}_{t})=(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0)^{+}$ with $n_{D}=252$, where simulations are carried out using daily time increments (i.e., $\Delta t=1$). The parameter space $\bm{\theta}:=(r,S_{0},\sigma,K)\in\bm{\Theta}$ is: $r\in[0.01,0.03]$, $S_{0}\in[80,120]$, $\sigma\in[0.05,0.25]$, and $K\in[90,110]$. For our feature $\bm{X}$, we experiment with the whole sum of Brownian increments $W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim(\bm{X})=1$. We also tried a more granular representation where we partition 252 Brownian increments evenly into 14 parts, each summing 18 consecutive increments: $[\bm{X}]_{i}=\sum_{j=18(i-1)+1}^{18i}\Delta W^{S}_{j}$ for $i\in[14]$, yielding $\dim(\bm{X})=14$. The neural network is trained on a dataset comprising $N_{\text{train}}=1.28\times 10^{6}$ samples, generated by uniformly sampling $\bm{\theta}\sim\bm{\Theta}$.
The neural network architecture consists of three components: a $\bm{\theta}$ network branch, a $\bm{X}$ network branch, and a combined network. This architecture incorporates several modern deep learning practices, including Batch Normalization (Ioffe and Szegedy 2015), Rectified Linear Unit (ReLU) activations (Nair and Hinton 2010), and skip connections (He et al. 2016) to enhance training stability and model performance. The $\bm{\theta}$ network branch processes the 4-dimensional input $\bm{\theta}$ using two fully connected layers with Batch Normalization and ReLU activation functions, outputting a 10-dimensional feature vector for subsequent processing. The $\bm{X}$ network branch uses two fully connected layers with width $\max(32,2\dim(\bm{X}))$ neurons. Finally, the combined network integrates outputs from both branches through concatenation and employs skip connections featured in ResNet architectures to preserve information flow. The concatenated features pass through two additional fully connected layers with Batch Normalization and ReLU activation to produce the final prediction for PEMC. The network is implemented using PyTorch (Paszke et al. 2019) and trained using the Adam optimizer (Kingma and Ba 2014) with a learning rate of $1\times 10^{-3}$, incorporating dropout layers (Srivastava et al. 2014) with a rate of 0.5 after each hidden layer to prevent overfitting.
3.5.2 Evaluation
For comparison, we implement PEMC, MC and CV method with geometric Asian option
$P_{G}(\{S_{t}\}_{t})=((\prod_{i=1}^{n_{D}}S_{t_{i}})^{\frac{1}{{n_{D}}}}-K,0)^{+},$
for the evaluation at $\bm{\theta}=(r,S_{0},\sigma,K)=(0.02,100,0.2,100)$. The CV estimator is given by
$\text{CV}=\frac{1}{n}\sum_{i=1}^{n}\left(P_{A}(\{S_{t}^{(i)}\}_{t})-P_{G}(\{S_{t}^{(i)}\}_{t})\right)+P_{G}^{\text{exact}}(\bm{\theta}),$
[…]
Table 1: Root Mean Squared Error from 300 Experiments
Method $n=1000$ $n=4000$ $n=9000$
Monte Carlo (MC) 0.2376 0.1173 0.0854
PEMC ($\dim\bm{X}=1$) 0.1509 0.0809 0.0481
PEMC ($\dim\bm{X}=14$) 0.0781 0.0397 0.0261
Geometric CV 0.0099 0.0051 0.0036
[…]
Performance of Estimators for Asian Options. Comparison of estimator across Monte Carlo (MC), PEMC, and Geometric Control Variate (CV) based on 300 experiments. Top: Mean squared errors (MSE) plot as a function of $n$ for (left) PEMC with $\dim\bm{X}=1$ and (right) PEMC with $\dim\bm{X}=14$, both compared against MC and Geometric CV. Bottom: Corresponding boxplots of the 300 estimates.
Given that all estimators are unbiased, our analysis focuses exclusively on variance-related metrics. The experimental results demonstrate the superior performance of the PEMC framework compared to the standard Monte Carlo (MC) estimator. In particular, in all sample sizes, the PEMC framework delivers visible variance reduction over MC: PEMC’s basic variant with $\dim\bm{X}=1$ achieves a 30-40% reduction in root mean squared error (RMSE) relative to MC, while the more sophisticated variant with $\dim\bm{X}=14$ attains an impressive 65-70% reduction in RMSE over MC. However, the Geometric CV emerges as the most efficient estimator by far, with RMSE an order of magnitude smaller than MC.
To better understand this hierarchy in estimator performance, it is helpful to recognize PEMC not as a specific estimator, but rather as a framework that latches on and enhance an existing baseline. In our current example, we chose standard MC as the baseline, and design our PEMC directly upon MC to improve it. This also illustrates a common state of practice: in most complex scenarios, sophisticated baselines like CV are not available, and practitioners must rely on some variants of naive MC methods. Then, PEMC can be applied to these baselines to achieve variance reduction. However, PEMC is a flexible framework designed to enhance general simulation baselines, including sophisticated ones, not just standard Monte Carlo. In particular, when a control variate method is available, such as in our current example, PEMC can be built upon it to potentially achieve even greater variance reduction. To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\bm{\theta})$ (Geometric CV baseline), and let the model $g$ learn
$g(\bm{\theta},\bm{X}(\bm{\theta}))\approx$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\bm{\theta})\bigg{|}(\bm{\theta},\bm{X}(\bm{\theta}))\bigg{]}$
$=$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})\bigg{|}(\bm{\theta},\bm{X}(\bm{\theta}))\bigg{]}+P_{G}^{\text{exact}}(\bm{\theta}).$ (11)
This formulation suggests we can train $g$ on the difference of $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})$ as labels and then add back the closed-form $P_{G}^{\text{exact}}(\bm{\theta})$ to $g$. We implemented this approach, which we call “Boost PEMC,” using the basic PEMC variant where $\bm{X}=W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim\bm{X}=1$. As shown in Table 2 and Figure 3.5.2, this Boost PEMC indeed achieves variance reduction over the sophisticated Geometric CV estimator, delivering a 35-40% reduction in RMSE, remarkably consistent with the relative performance gain we observed when applying PEMC to the MC baseline in Figure 3.5.2.
Table 2: Root Mean Squared Error from 300 Experiments
Method $n=1000$ $n=4000$ $n=9000$
Boost PEMC ($\dim\bm{X}=1$) 0.0065 0.0031 0.0021
Geometric CV 0.0099 0.0051 0.0036
[…]
Performance of CV-based Estimators for Asian Options. Comparison of estimator performance for Boost PEMC and Geometric Control Variate (CV) based on 300 experiments. Left: Mean squared errors (MSE) plot as a function of $n$. Right: Corresponding boxplots of 300 estimates.
This experiment with Asian options under the GBM model serves as a controlled study to demonstrate PEMC’s effectiveness as a variance reduction framework. In the following section, we go beyond this over-simplified setting and apply PEMC to practical applications in complex derivative pricing.
4 Applications
After establishing the theoretical foundations of PEMC and demonstrating its effectiveness in controlled environments, we now apply it to real-world financial scenarios where the complexity of models often renders standard variance reduction techniques difficult or unavailable. In this section, we examine PEMC in practice through two complex pricing problems. In particular, we examine variance swaps pricing under stochastic (local) volatility models (Dupire et al. 1994, Derman and Kani 1994, Gatheral 2006), and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework (Carmona 2007). In both cases, the path-dependent nature of the contracts and the complexity of the model dynamics make traditional control variates difficult.
4.1 Variance Swaps Pricing
Variance swaps are financial derivatives that enable investors to trade future realized volatility against current implied volatility (Bayer et al. 1999). Unlike traditional options, which derive their value from the underlying asset’s price, variance swaps are based on the variance of the asset’s returns over a specified period (Derman and Fantazzini 1999). This unique structure allows for pure exposure to volatility, making variance swaps valuable tools for risk management and speculative strategies (Gatheral 2006, Protter 2010).
In this subsection, we present an application of the PEMC framework in pricing a variance swaps, with the underlying asset paths assumed to be generated according to a Heston model (1) and a Stochastic Local Volatility (SLV) model (Ren et al. 2007), respectively. These models represent two distinct approaches to stochastic volatility modeling in practical derivatives pricing: the Heston model offers a parsimonious parametric form with relatively few parameters to calibrate, while the SLV model provides greater flexibility through a data-intensive, non-parametric specification of the volatility surface. By examining both cases, we demonstrate PEMC’s versatility across the spectrum of model complexity.
4.1.1 Variance swaps under a Heston model
Under the Heston model (1), and the procedure follows from the illustration in Section 2.5. Given $\bm{\theta}_{\text{model}}:=(r,\kappa,\eta,\delta,\rho)$, the asset price $\{S_{t}\}_{t}$ is simulated by Euler’s scheme with fixed $T$ and $\Delta t$:
$\nu_{t+\Delta t}$ $\leftarrow\kappa(\eta-\nu_{t})\Delta t+\delta\sqrt{\nu_{t}}\Delta W^{\nu}_{t},$
$S_{t+\Delta t}$ $\leftarrow S_{t}\exp\Big{(}\big{(}r-\tfrac{1}{2}\nu_{t+\Delta t}^{2}\big{)}\Delta t+\sqrt{\nu_{t+\Delta t}}\Delta W^{S}_{t}\Big{)}.$
where $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$. For each asset path $\{S_{t}\}_{t}$, the payoff is
$f_{\bm{\theta}_{\text{payoff}}}(\{S_{t}\}_{t}):=\sqrt{\frac{252\times\sum_{t=1}^{T/\Delta t}\log(S_{t}/S_{t-1})^{2}}{T/\Delta t}}-K.$ (12)
Note that due to the linear form of strike $K$ in the payoff, it suffices to effectively removes strike $K$ from training. We also conveniently omit any discount factor. The parameters space $\bm{\Theta}$, as well as the evaluation $\bm{\theta}$, is summarized in Table 4.
Table 3: Parameter Setup in Heston Model
mode $S_{0}$ $r$ $\nu_{0}$ $\kappa$ $\eta$ $\delta$ $\rho$ $T$, $\Delta t$
$N_{\text{train}}=3,000,000$ $[50,150]$ $[0.01,0.05]$ $[0.1,0.375]^{2}$ $[1.5,4.5]$ $[0.1,0.3]^{2}$ $[0.1,1.0]$ $[-0.2,-0.9]$ $1$, $1/252$
evaluation $100$ $0.02$ $0.25^{2}$ $3.0$ $0.2^{2}$ $0.6$ $-0.4$
The feature $\bm{X}$ is constructed by
$W_{T}^{S}:=\sum_{j=1}^{T/\Delta t}\Delta W^{S}_{j\Delta t}\quad\text{ and }\quad W_{T}^{\nu}:=\sum_{j=1}^{T/\Delta t}\Delta W^{\nu}_{j\Delta t},$
which are jointly distributed as $\sqrt{T}\cdot\mathcal{N}(0,[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$ just as in Section 2.5. For the prediction model training in PEMC algorithm 1, a basic Residual Network (ResNet)-style architecture (He et al. 2016) is adopted and trained on $N_{\text{train}}=3\times 10^{6}$ samples. The network consists of two Multi-Layer Perceptron (MLP) blocks connected through skip connections. The first MLP block processes $(\bm{\theta},\bm{X})$ together and pass them through hidden layers with dimension 512, eventually producing a 256-dimensional intermediate representation. This output is then fed into the second MLP block, which maintains the same hidden dimension of 256 throughout its layers. The complete architecture specification is provided in Table 4. At evaluation stage (i.e., Algorithm 2), we use $\bm{\theta}$ in Table 4 and set $N=10n$ as before.
Table 4: Hyper-parameter Setup for Neural Network
1st MLP block 2nd MLP Block
hidden dim: 512; output dim: 256 hidden dim: 256
Following our evaluation methodology from Section 3.5, we first establish a ground truth value by averaging $5\times 10^{7}$ MC samples. We then implement both standard MC and PEMC estimators (with $N=10n$) using sample sizes $n\in\{1000,2000,4000,8000,10000,20000\}$, repeating each experiment 200 times to assess sampling variability. As shown in Table 5, PEMC achieves a 30-50% reduction in root mean squared error compared to standard MC, consistent with the variance reduction levels we observed in Section 3.5. This demonstrates that PEMC’s effectiveness as a variance reduction technique carries over to more complex models and derivative products.
Table 5: RMSE from 200 Experiments for Pricing Variance Swap under the Heston Model
Method $n=1000$ $n=2000$ $n=4000$ $n=6000$ $n=8000$ $n=10000$ $n=20000$
Monte Carlo (MC) 0.0996 0.0659 0.0517 0.0391 0.0366 0.0290 0.0218
PEMC 0.0599 0.0387 0.0263 0.0233 0.0213 0.0179 0.0126
4.1.2 Variance swaps under an SLV model
Having demonstrated PEMC’s effectiveness for the Heston model in Section 4.1.1, we turn to the more sophisticated Stochastic Local Volatility (SLV) framework. While parametric models like Heston offer mathematical elegance and computational efficiency, they often lack the flexibility to fully capture market dynamics (Gatheral 2006). SLV models emerged as a hybrid approach that combines the market-implied local volatility surface with stochastic volatility dynamics (Guyon and Henry-Labordère 2014), providing practitioners with greater calibration flexibility and more accurate price reproduction across strike-maturity ranges (Ren et al. 2007).
The increased modeling power of SLV, however, comes with significant computational challenges. It is known that SLV has no closed-form formula for even vanilla options (Tataru and Fisher 2010). Moreover, unlike the Heston model’s parsimonious parameter set, SLV requires handling a full volatility surface discretized on a dense 2D grid, effectively turning our parameter space $\bm{\Theta}$ for PEMC into a high-dimensional object. To efficiently process this grid-structured volatility data, we adopt a Convolutional Neural Network (CNN) architecture (LeCun et al. 1998) for our PEMC implementation, effectively feeding the local volatility surface as an image into PEMC. The CNNs are particularly well-suited for this task as they naturally exploit the spatial relationships in the volatility surface, similar to their success in image processing tasks (He et al. 2016, Krizhevsky et al. 2012).
Under the SLV model, we adopt the following SDEs for simulating the asset price $\{S_{t}\}_{t}$:
$\mathrm{d}S_{t}$ $=rS_{t}\mathrm{d}t+\sigma(S_{t},t)e^{\nu_{t}}S_{t}\mathrm{d}W_{t}^{S},$ (13)
$\mathrm{d}\nu_{t}$ $=\kappa(\eta_{t}-\nu_{t})\mathrm{d}t+\delta\mathrm{d}W_{t}^{\nu},$ (14)
with $\langle\mathrm{d}W_{t}^{S},\mathrm{d}W_{t}^{\nu}\rangle=\rho$ and $\eta_{t}:=-\frac{\delta^{2}}{2\kappa}(1+e^{-2\kappa t}).$ Here $\sigma(\cdot,\cdot):\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is 2D function representing the local volatility surface and $\exp(\nu_{t})$ is a stochastic multiplier with $\exp(\nu_{0})=1$ and $\mathbb{E}[e^{2\nu_{t}}]=1$. In practice, local volatility surfaces are calibrated to and stored as discrete two-dimensional grids indexed by asset prices (spot) and time. During simulation, these discrete values are interpolated as needed to obtain volatilities at arbitrary price-time points (Coleman et al. 2001). For our PEMC implementation, we treat this discrete grid as part of our input parameter $\bm{\theta}_{\text{model}}$, reflecting how the market-calibrated surfaces would be used in practice. While local volatility models (Dupire et al. 1994, Derman and Kani 1994) and their calibration to market data constitute an extensive research area in their own right, our focus here is on the PEMC implementation where, a calibrated local volatility has been given for evaluation, regardless of the calibration method used to obtain it.
Following practical conventions but also simplifying for the sake of illustration, we store $\sigma(\cdot,\cdot)$ on a $|\mathcal{S}|\times|\mathcal{T}|$ grid, where $\mathcal{S}$ contains $|\mathcal{S}|$ equally-spaced price points in $[S_{\text{surface}}^{\min},S_{\text{surface}}^{\max}]$ and $\mathcal{T}$ contains $|\mathcal{T}|$ equally-spaced time points in $[t_{\text{surface}}^{\min},t_{\text{surface}}^{\max}]$. At each point on the grid, we store the value of local volatility according to Carmona (2007):
$\sigma_{base}^{2}(x,t)=\frac{\sum_{i=0}^{2}p_{i}\tau_{i}e^{-x^{2}/(2t\tau_{i}^{2})-t\tau^{2}_{i}/8}}{\sum_{i=0}^{2}(p_{i}/\tau_{i})e^{-x^{2}/(2t\tau_{i}^{2})-t\tau_{i}^{2}/8}},\quad\text{with}~{}~{}~{}p_{0}:=1-p_{1}-p_{2},~{}~{}x:=\log(S_{t}/S_{0}),$ (15)
for $p_{0}=0.3,p_{1}=0.5,p_{2}=0.2,\tau_{0}=0.4,\tau_{1}=0.3,\tau_{2}=0.6$ as in Figure 2 in Carmona (2007). While this analytical form (15) is used in our data-generating process, it is important to note that PEMC treats the surface as any market-calibrated volatility surface - accessing it only through its discrete grid values. When sampling $\bm{\theta}\sim\bm{\Theta}$ to produce $\sigma$ grid, we add a $\mathcal{N}(0,\xi^{2})$ noise independently to all the $|\mathcal{S}|\times|\mathcal{T}|$ points in the grid, on top of their baseline value $\sigma_{base}$ (15). A path is then generated by Euler’s scheme:
$\nu_{t+\Delta t}$ $\leftarrow\nu_{t}+\kappa(\eta_{t}-\nu_{t})\Delta t+\delta\Delta W^{\nu}_{t},$
$S_{t+\Delta t}$ $\leftarrow S_{t}\exp\Big{(}\big{(}r-\tfrac{1}{2}\tilde{\nu}_{t+\Delta t}^{2}\big{)}\Delta t+\tilde{\nu}_{t+\Delta t}\Delta W^{S}_{t}\Big{)},\quad\text{where}~{}~{}\tilde{\nu}_{t+\Delta t}=\widehat{\sigma}(S_{t},t)e^{\nu_{t+\Delta t}},$
with $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}(\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$, and $\widehat{\sigma}(S_{t},t)$ obtained by interpolations of the grid $\sigma$. The input $(\bm{\theta},\bm{X})$ can be represented as
$\text{feature}_{i}=\{\underbrace{\{\sigma^{(s,t)}\}_{s\in\mathcal{S},t\in\mathcal{T}}}_{\text{surface info }},\underbrace{r,\delta,\kappa,\rho,\mu}_{\bm{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0}}_{\bm{\theta}_{\text{simulation}}},\underbrace{K}_{\bm{\theta}_{\text{payoff}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}(\bm{\theta})}\}.$
The parameters space $\bm{\Theta}$, as well as the evaluation $\bm{\theta}$, is summarized in Table 6.
Table 6: Parameter Setup in SLV Model
mode $\xi$ $S_{0}$ $r$ $\kappa$ $\lambda$ $\rho$ $T$, $\Delta t$
$N_{\text{train}}=3,000,000$ $0.02$ $[50,150]$ $0.02$ $[1.5,4.5]$ $[0.1,1.0]$ $[-0.2,-0.9]$ $1$, $1/252$
evaluation 00 $100$ $0.02$ $3.0$ $0.5$ $-0.5$
To handle the 2D grid of high-dimensional volatility surface data, we design a two-branch neural network architecture, which is illustrated in Figure 1. The first branch processes the discretized volatility surface $\sigma^{2}(x,t)$ using a CNN architecture inspired by VGG (Simonyan and Zisserman 2014), which has become a standard choice for image processing tasks and is readily available in modern deep learning packages like PyTorch (Paszke et al. 2019). This branch then consists of two 2D convolutional layers interspersed with ReLU activations, followed by a MaxPool2d operation for dimensionality reduction. The surface features are then flattened through a fully connected layer to produce an embedding. The second branch handles the remaining model parameters in $(\bm{\theta},\bm{X})$ through a series of fully connected layers with dropout regularization, batch normalization, and ReLU activations, ultimately producing another embedding. Finally, the two separate embeddings are then fed into a “Synthesizer” module, which combines the information through additional fully connected layers with dropout and ReLU activations to produce the final prediction. This architecture choice is motivated by the proven effectiveness of CNNs in handling grid-structured data (LeCun et al. 1998, He et al. 2016), and particularly the VGG architecture’s success in extracting hierarchical features while maintaining relative simplicity (Simonyan and Zisserman 2014). The details of NN architecture are summarized in Table 7.
Figure 1: Neural Network architecture for the SLV Model.
Table 7: Neural Network Architecture Parameters
CNN Branch Feed-forward Branch
kernel size: 3, stride: 1, padding: 1
max pooling (kernel: 2, stride: 2, padding: 0)
hidden dim: 512
output dim: 128
Using the same experimental setup as before, we evaluate PEMC’s performance in the SLV setting across sample sizes $n\in\{1000,2000,4000,8000,10000,20000\}$ with $N=10n$, benchmarking against a ground truth computed from $5\times 10^{7}$ MC samples. The results, based on 200 independent experiments and shown in Table 8, reveal that PEMC’s effectiveness persists even in this more complex setting. Despite the added complexity of handling high-dimensional volatility surfaces, PEMC achieves a 30-40% reduction in mean squared error compared to standard MC, demonstrating its robustness as a variance reduction technique across different model frameworks.
Table 8: RMSE from 200 Experiments for Pricing Variance Swap under the SLV Model
Method $n=1000$ $n=2000$ $n=4000$ $n=6000$ $n=8000$ $n=10000$ $n=20000$
Monte Carlo (MC) 0.0206 0.0145 0.0101 0.0075 0.0064 0.0065 0.0047
PEMC 0.0130 0.0088 0.0061 0.0055 0.0040 0.0040 0.0027
4.2 Swaptions in HJM Models
Interest rate derivatives play a crucial role in financial markets, with swaptions being particularly important instruments for managing interest rate risk (Brigo and Mercurio 2006). A swaption gives its holder the right to enter into an interest rate swap at a future date, providing flexibility in hedging future interest rate exposures (Hull 2018). The pricing of these instruments, however, presents significant computational challenges due to the high-dimensional nature of interest rate modeling (Andersen and Piterbarg 2010). In this subsection, we demonstrate PEMC’s application to swaption pricing under the Heath-Jarrow-Morton (HJM) framework (Heath et al. 1992). The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models. For illustration purposes, we focus on a one-factor specification with exponential volatility structure (Glasserman 2013b), though the framework readily extends to multi-factor cases. This model choice represents a balance between analytical tractability and practical relevance, allowing us to demonstrate PEMC’s effectiveness in a high-dimensional setting while maintaining computational feasibility.
4.2.1 Swaptions and HJM Model
A swaption is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date. In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite. Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$. The value of this swap at time $t^{\prime}_{0}$ is:
$V_{t^{\prime}_{0}}=C\left(R\sum_{l=1}^{n_{p}}B(t^{\prime}_{0},t^{\prime}_{l})\Delta t^{\prime}+B(t^{\prime}_{0},t^{\prime}_{n_{p}})-1\right),$
where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}$. A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$. The payoff of the swaption is simly
$\max(0,V_{t^{\prime}_{0}}),$
[…]
$df(t,T)=\mu(t,T)\,dt+\sigma(t,T)^{\top}dW(t),$
where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion. In contrast to short-rate models (e.g., Vasicek (Vasicek 1977) or Cox-Ingersoll-Ross (Cox et al. 1985)), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates (Heath et al. 1992). The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures (Brigo and Mercurio 2006, Andersen and Piterbarg 2010). However, the model’s generality also leads to the need for sophisticated numerical methods for simulation Glasserman (2013b). A key property of the HJM model is the no-arbitrage condition (Glasserman 2013b), which specifies the drift completely by the volatility:
[…]
Thus, in the HJM framework, the model is fully specified by defining the initial forward rate curve $f(0,T)$ and the structure of the volatility $\sigma(t,T)$. In our experiment we used a simple one factor HJM for illustration.
4.2.2 PEMC for HJM
Just as with the local volatility surface case, in practice $\sigma(t,T)$ cannot be predefined parametrically and must be calibrated from market data of caps, floors, and swaptions, yielding a discrete grid of values. However, for demonstration purposes, we employ a classical exponential decay specification as our baseline model inspired from Glasserman (2013b):
$\sigma_{base}(t,T)=\sigma_{0}\exp(-\alpha_{\sigma}(T-t))$ (17)
with $\sigma_{0}$ and $\alpha_{\sigma}$ as part of $\bm{\theta}$. Similarly, for a baseline initial forward curve, we use:
$f_{base}(0,T)=f_{0}+c_{f}(1-\exp(-\alpha_{f}T)).$ (18)
with $f_{0},c_{f}$, and $\alpha_{f}$ part of the $\bm{\theta}$. While this analytical form serves as our data-generating process, PEMC accesses it only through its discrete grid values with added noise. This approach mirrors our treatment of the local volatility surface in the previous section, where we used the parametric form in Carmona (2007) solely as a realistic baseline for generating training data.
Indeed, while HJM and these analytical forms (17)-(18) are formulated in continuous time, in practice we need to implement numerical discretization. Following the scheme in Glasserman (2013b), one discretizes both the time axis and the set of maturities i.e., time steps $\mathcal{T}_{t}=\{t_{0},t_{1},\ldots,t_{N_{T}}\}$ and maturity points $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{M}}\}$ forming a time-maturity grid. In our experiment, for simplicity, we assume they share one grid $\mathcal{T}$ of size $|\mathcal{T}|$. Then, when sampling $\bm{\theta}\sim\bm{\Theta}$, one first sample $\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}$, then one sample the grids $\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}$ and $\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}$ from (17) and (18) with added noise $\mathcal{N}(0,(\frac{\sigma_{0}}{2(t_{j}+5)})^{2})$ and $\mathcal{N}(0,(\frac{1}{100(t_{j}+5)})^{2})$ respectively, on top of their baseline values. The noise level $\xi=\frac{\sigma_{0}}{2(t+5)}$. Paths are then generated using the discretization scheme described in Glasserman (2013b) (e.g., ${f}(t_{i},t_{j})={f}(t_{i-1},t_{j})\;+\;{\mu}(t_{i-1},t_{j})(t_{i}-t_{i-1})\;+\;\sqrt{t_{i}-t_{i-1}}{\sigma}(t_{i-1},t_{j})\mathcal{N}(0,1)$ where ${\mu}(t_{i-1},t_{j})$ is the discretized drift term determined through (16)). We refer interested readers to Glasserman (2013b) for the complete derivation and implementation details of this standard simulation scheme.
For PEMC, our the parameter $\bm{\theta}\sim\bm{\Theta}$ includes the volatility parameters $(\sigma_{0},\alpha_{\sigma})$ in (17) and initial forward curve parameters $(f_{0},c,\alpha_{f})$ in (18). The swaption’s fixed rate is set as $R=\exp(-\frac{\sum_{i}^{n-1}f(0,t^{\prime}_{i})}{T_{\text{final}}-t^{\prime}_{0}})$, reflecting the common practice where swap rates are typically determined in reference to the prevailing forward rate curve rather than being arbitrarily chosen. Here the parameters are notional amount $C$, start time $t^{\prime}_{0}$, payment interval $\Delta t^{\prime}$, and number of payments $n_{p}$. The simulation uses time step $\Delta t$ up to final maturity $T_{\text{final}}$. During training, parameters are sampled uniformly from the ranges specified in Table 9 where the evaluation $\bm{\theta}$ is also listed.
Table 9: Parameter Setup in HJM Model
Mode $\sigma_{0}$ $\alpha_{\sigma}$ $f_{0}$ $c_{f}$ $\alpha_{f}$ $R$ $C$ $t^{\prime}_{0}$ $\Delta t^{\prime}$ $n_{p}$ $\Delta t$ $T^{*}$
$N_{\text{train}}=3,000,000$ $[0.01,0.03]$ $[0.001,0.9]$ $[0.01,0.03]$ $[0.01,0.05]$ $[0.001,0.9]$ $\exp(-\frac{\sum_{i=0}^{n-1}f(0,t^{\prime}_{i})}{T^{*}-t^{\prime}_{0}})$ $100$ $5$ $1$ $20$ $1/52$ $25$
evaluation $0.0015$ $100$ $0.02$ $3.0$ $0.5$
The input for PEMC is, similar as before:
$\text{feature}=\{\underbrace{\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}}_{\text{volatility structure}},\underbrace{\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}}_{\bm{\theta}_{\text{model}}},\underbrace{\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}}_{{\text{initial forward curve}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}}\}.$
To handle both the two-dimensional grid of volatility structure $\sigma(t,T)$ and the one dimensional grid of initial forward curve $f(0,T)$, we design a three-branch neural network architecture, illustrated in Figure 2. The first branch, labeled“2D Function Encoder”, processes the 2d volatility structure grid using a CNN architecture with two 2D convolutional layers, each followed by batch normalization. The branch concludes with an average pooling operation and produces an embedding. The second branch processes the initial forward curve $f(0,T)$ grid through a “1D Function Encoder” utilizing 1D convolutional layers - a natural choice for sequential data (LeCun et al. 1998) - followed by batch normalization and average pooling to produce another embedding. The third branch handles the remaining input through fully connected layers with batch normalization. Finally, these three separate embeddings are then fed into a “Synthesizer” module that combines the information through multiple fully connected layers with batch normalization, ultimately producing the final prediction. This architecture effectively leverages both the spatial structure of the volatility surface through 2D CNNs (Simonyan and Zisserman 2014), the sequential nature of the forward curve through 1D CNNs (Oord et al. 2016), and the scalar parameters through standard deep learning techniques (He et al. 2016). The complete network architecture is detailed in Table 10.
Figure 2: Neural network architecture for modeling Swaption payoff.
2D function branch parameters 1D function branch parameters Vector feature branch parameters
Kernel size: (1, 3) Stride: (1, 3) Padding: 0 AvgPool2d kernel size: (2, 2) AvgPool2d stride: (2, 2) AvgPool2d padding: 0 Kernel size: 10 Stride: 3 Padding: 0 AvgPool1d kernel size: 2 AvgPool1d stride: 2 AvgPool1d padding: 0 Hidden dim: 512 Output dim: 128
Feed-forward Synthesizer Parameters
Hidden dim: 128
Output dim: 1
Table 10: Hyper-parameter setup for the neural network
Finally, following the same evaluation methodology, we assess PEMC’s performance in the HJM swaption pricing setting across sample sizes $n\in\{1000,3000,5000,7000,9000,11000\}$ with $N=10n$, using $5\times 10^{7}$ MC samples to establish the ground truth value. As presented in Table 11, the outcomes from 200 independent experiments confirm that PEMC remains highly effective within the context of interest rate derivatives. Despite the added complexity of managing both volatility structures and forward rate curves, PEMC achieves a 45-50% reduction in RMSE to standard Monte Carlo methods. This performance aligns with the variance reduction levels observed in our earlier examples. The boxplot is shown in Figure 4.2.2. The consistent effectiveness of PEMC across various financial instruments and modeling frameworks further underscores its versatility as a robust variance reduction technique.
Table 11: Root Mean Squared Error from 200 Experiments
Method $n=1000$ $n=3000$ $n=5000$ $n=7000$ $n=9000$ $n=11000$
Monte Carlo (MC) 0.0096 0.0062 0.0048 0.0039 0.0035 0.0029
PEMC 0.0055 0.0028 0.0024 0.0019 0.0018 0.0015
Boxplots for HJM Experiments.
4.3 Discussions and Extensions
We conclude with several practical observations and potential extensions of the PEMC framework.
4.3.1 Evaluation Metric
First, while training the neural network estimator $g$ using mean squared error (MSE) loss is common, it is not always clear how to interpret the resulting MSE score. Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem. This ambiguity makes it challenging to determine when the network is sufficiently trained. To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$. If the network approximates this expectation well, then the sample average of $g(\bm{X})$ should be close to the sample average of $f(\bm{Y})$ over a given dataset. One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages. If $\mathbb{E}[g(\bm{X})]\approx\mathbb{E}[f(\bm{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation. Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in machine learning workflows. More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard Monte Carlo.
4.3.2 XVA, Greeks and Quasi-Monte Carlo
Beyond pricing exotic options, this approach naturally extends to other computationally intensive Monte Carlo settings in quantitative finance. Notably, adjustments like Credit valuation adjustment (CVA) and more complex XVAs often require vast simulation runs, sometimes taking days to complete (Green et al. 2015, Andersen et al. 2017). Integrating PEMC with a well-trained neural network can cut these computational times drastically while preserving accuracy.
Moreover, in the pricing of exotic derivatives and other path-dependent instruments, automatic differentiation (AD) techniques are often used to compute Greeks efficiently. By leveraging modern deep learning frameworks—such as PyTorch or TensorFlow—that support automatic differentiation natively, we can easily differentiate the trained network $g$ with respect to input parameters (Giles and Glasserman 2006). This enables quick and accurate sensitivity analysis without resorting to finite-difference approximations or nested simulations, which could lead to another direction of extension for PEMC.
Finally, although we have focused on standard Monte Carlo sampling, combining PEMC with Quasi Monte Carlo (QMC) techniques may offer further variance reduction benefits (Caflisch 1998). However, it is important to note that QMC methods introduce a low-discrepancy bias, thus forgoing the unbiasedness property of pure Monte Carlo. Balancing unbiasedness with the additional variance reduction from QMC is an intriguing direction for future research.
5 Conclusion
In this paper, we introduced the Prediction-Enhanced Monte Carlo (PEMC) framework, a novel approach that leverages machine learning-based predictors to serve as flexible, data-driven control variates. By doing so, we preserve the unbiasedness of Monte Carlo estimators while significantly enhancing their efficiency in complex settings where traditional variance reduction techniques such as control variates are difficult or impossible to construct.
We demonstrated the practical value of PEMC through applications: pricing variance swaps under both a Heston and a Stochastic Local Volatility model, and pricing swaptions under the Heath-Jarrow-Morton framework. These examples, each introducing their own modeling and computational challenges, highlight PEMC’s versatility and effectiveness.
While we have demonstrated the power of the PEMC framework on an options pricing application, we believe PEMC represents a new frontier for variance reduction, blending modern machine learning methodologies with classical Monte Carlo simulation.
A promising avenue for future research is to explore the connection between PEMC and meta-learning, which seeks to develop models capable of learning how to construct effective control variates from related problems. This approach reduces the need for problem-specific tuning, as demonstrated by Finn et al. (2017). Additionally, the recent surge in consistency models and diffusion-based generative techniques in machine learning (Ho et al. (2020), Song et al. (2021)) opens up new possibilities. While their direct application to pricing and control variates remains largely unexplored, these flexible generative models could, in theory, produce advanced proposal distributions or control variates. Nonetheless, significant methodological and theoretical gaps must be bridged before such methods can be integrated seamlessly and rigorously into financial computation pipelines.
Another intriguing direction lies in leveraging recent advancements in causal inference and robust statistics, particularly methods designed to handle shifts in underlying data distributions (see Li and Lam (2020), Bodnar et al. (2022), Ellickson et al. (2023), Kang and Schafer (2007), Hernán and Robins (2020)). For instance, studies such as Angelopoulos et al. (2023), Li and Lam (2020), Quionero-Candela et al. (2009), Bickel et al. (2009), Sugiyama and Kawanabe (2009) propose robust regression techniques under covariate shift. These advances suggest that methodologies initially developed for causal inference—designed to ensure stability against model misspecification and distributional drift—could inform the creation of more versatile and robust variance reduction schemes. Incorporating ideas from doubly robust estimators into the ML-based control variate framework could pave the way for hybrid methods like PEMC to better handle uncertain or dynamic environments.

### v2

@@
-
+Prediction-Enhanced Monte Carlo: A Machine Learning View on Control Variate
+Abstract.
+For many complex simulation tasks spanning areas such as healthcare, engineering, and finance, Monte Carlo (MC) methods are invaluable due to their unbiased estimates and precise error quantification.
+Nevertheless, Monte Carlo simulations often become computationally prohibitive, especially for nested, multi-level, or path-dependent evaluations lacking effective variance reduction techniques.
+While machine learning (ML) surrogates appear as natural alternatives, naïve replacements typically introduce unquantifiable biases.
+We address this challenge by introducing Prediction-Enhanced Monte Carlo (PEMC), a framework that leverages modern ML models as learned predictors, using cheap and parallelizable simulation as features, to output unbiased evaluation with reduced variance and runtime.
+PEMC can also be viewed as a “modernized" view of control variates, where we consider the overall computation-cost-aware variance reduction instead of per-replication reduction, while bypassing the closed-form mean function requirement and maintaining the advantageous unbiasedness and uncertainty quantifiability of Monte Carlo.
+We illustrate PEMC’s broad efficacy and versatility through three examples: first, equity derivatives such as variance swaps under stochastic local volatility models; second, interest rate derivatives such as swaption pricing under the Heath-Jarrow-Morton (HJM) interest-rate model.Finally, we showcase PEMC in a socially significant context—ambulance dispatch and hospital load balancing—where accurate mortality rate estimates are key for ethically sensitive decision-making.
+Across these diverse scenarios, PEMC consistently reduces variance while preserving unbiasedness, highlighting its potential as a powerful enhancement to standard Monte Carlo baselines.
+Equal Contribution. 1Morgan Stanley. 2Department of IEOR, Columbia University. The Work was done during the internship at Morgan Stanley in the Machine Learning Research group.
+     Corresponding author: fengpei.li@morganstanley.com
 […]
-[…]
-1 Introduction
-Monte Carlo (MC) simulation is a foundational tool in computational finance and other domains where closed-form solutions for complex stochastic models are unavailable.
-Its principal virtues are unbiasedness and the ability to rigorously quantify uncertainty through confidence intervals.
-By relying on random sampling, MC provides well-understood statistical guarantees and error assessments.
-However, Monte Carlo methods face significant challenges when applied to complex, nested, or path-dependent settings.
-As Glasserman (2013a) notes,“The slow convergence of Monte Carlo - its $O(1/\sqrt{n})$ error after $n$ replications - is a persistent obstacle to its wider use.” In other words, while MC excels at delivering unbiased estimates with reliable uncertainty quantification, it often becomes a “slow evaluation”.
-For example, in the context of derivative pricing, the challenge is further compounded for path-dependent options.
-Hull (2018) points out that “for path-dependent options, the entire price path must be simulated, significantly increasing computational complexity.” Similarly, Broadie and Glasserman (1996) state that “in many financial simulations, the value at each time step is a function of the previous step, creating a sequential dependency that hinders straightforward parallelization.”
-In recent years, machine learning (ML) prediction methods have emerged as a fast alternative for approximating expensive functions, including prices of options and of other financial derivatives (Ferguson and Liang (2018), Bayer and Stemper (2019)).
-Neural networks and other ML techniques can learn complex mappings quickly, and once trained, can generate predictions at negligible marginal cost - what practitioners often call ”fast evaluation” (Horváth et al.
-(2021)).
-However, their black-box nature and lack of inherent statistical control make it difficult to dissect error and impossible to ensure unbiasedness Hutchinson et al.
-(1994).
-As a result, pure ML approaches often yield biased estimates with no guaranteed measure of error, making them potentially unsuitable for high-stakes applications like risk management where reliable uncertainty quantification is crucial Buehler et al.
-(2019).
-This tension between MC’s reliability and ML’s efficiency motivates the exploration of classical variance reduction techniques, particularly control variates (CV).
-Control variates provide a systematic approach to improving MC efficiency by exploiting auxiliary functions with known expectations ( Glasserman (2013b), Asmussen and Glynn (2007)).
-When one finds a quantity that is highly correlated with the target while also possessing an analytically known mean, then the variance reduction could be substantial (Boyle et al.
-(1997)).
-However, as in many path-dependent exotic derivatives, under many complex models — effective control variates are simply unavailable (see, for instance, Duffie and Glynn (1995)).
-As noted in Glasserman (2013b),“some exotic payoffs are so irregular that finding an effective control variate becomes a significant challenge in itself.”
-Our work seeks to bridge these two approaches, offering a hybrid framework that leverages the best of both worlds.
-We propose a novel control variate (CV)-based methodology, termed Prediction Enhanced Monte Carlo (PEMC), which extends the classical MC framework by incorporating modern ML techniques while preserving the unbiasedness and rigorous error quantification of MC.
-In the context of option pricing, unlike traditional ML applications in pricing derivatives, where models are trained to directly predict option prices, we train a certain conditional expectation as CVs, where we break away from two major characteristics of traditional CVs:
-  • Known mean requirement: Unlike traditional CVs that demand the auxiliary function have a known mean (a challenging condition for complex or irregular payoffs), our approach does not rely on this assumption.
-Instead, PEMC utilizes ML-optimized predictors that are simple to estimate and still provide effective variance reduction.
-  • Per-run variance reduction: While traditional CV techniques are designed to reduce variance on a per-sample basis—pairing each sample with a corresponding control variate—PEMC takes a broader approach.
-It focuses on achieving variance reduction across the entire simulation scheme, rather than targeting individual sample variance reduction.
-Through a careful combination of classical MC principles and ML innovation, our approach eliminates bias from the black-box nature of ML models while also enjoying the unbiasedness and variance reduction provided by the traditional CV methods.
-By combining these elements, PEMC provides a principled and scalable framework that addresses the shortcomings of both classical CVs and ML-based methods, achieving reliable and interpretable results even in high-dimensional or irregular settings.
-1.1 Our Contribution: A Modernized View of Control Variates
-We introduce Prediction-Enhanced Monte Carlo (PEMC), a novel framework for evaluating Monte Carlo, which can be viewed as a modernized view of control variates: rather than searching for a single closed-form control variate with a known mean, we train a flexible predictor that acts as a control variate over the entire scheme.
-Crucially, the construction maintains unbiasedness and allows for valid statistical inference—just as in MC—while leveraging the representational power of ML models to approximate conditional expectations that can be evaluated efficiently.
-The key innovation of PEMC lies in relaxing the classical requirement that control variates must have a known mean.
-This is achieved through a two-stage estimation process: the first stage uses a small number of full simulations to estimate the difference between the target and the predictor, while the second stage computes the predictor average using a large number of computationally inexpensive simulations.
-This design makes the full simulation cost modest, despite harnessing a large number of cheap samples to drive down variance of the predictor.
-Consequently, we achieve a form of variance reduction not just on a per-path basis (as the traditional CV), but scheme-wide: the entire simulation procedure, not a given sample design, is optimized to reduce the effective scheme variance vs scheme cost tradeoff, not the per sample variance vs per sample cost tradeoff.
-This differs from the defining characteristic of traditional CV methods where one control variate is applied per run.
-Our PEMC approach modernizes this idea, offering a “scheme-wide” variance reduction that can adapt to complex models and broad ranges of parameters.
-Compared to purely machine learning-based methods, PEMC transforms the role of ML in Monte Carlo tasks.
-Rather than relying on ML as a direct surrogate predictor for the quantity of interest—which often introduces biased approximations with unquantifiable errors—PEMC uses ML to construct a robust, unbiased estimator augmented by modernized control variates.
-This approach enables fast predictions while maintaining Monte Carlo as the de facto standard of evaluation.
-1.2 Literature Review
-The literature on enhancing MC with ML approximators is extensive, but existing ML-related methods for control variates and variance reduction often impose stringent conditions that limit their practicality.
-Techniques such as the reproducing Stein kernel approach (Oates et al.
-(2017), Li and Zhang (2023)), regularized least squares for CV construction (Portier and Segers (2018), South et al.
-(2022), Leluc et al.
-(2021)), and adaptive CV schemes (Henderson and Glynn (2002), Henderson and Simon (2004), Kim and Henderson (2007)), as well as $\mathscr{L}^{2}$ function approximation frameworks (Maire (2003)), can in theory deliver powerful improvements—including even “supercanonical” convergence rates.
-But these methods usually require strict assumptions to ensure that the control variate’s mean is known and that it can be easily integrated into the simulation.
-In many practical settings, these assumptions fail, thus limiting the application scope of these methods.
-Even though some approaches can theoretically achieve supercanonical convergence rates, their applicability is limited by these constraints.
-In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively in scenarios where auxiliary means are unknown.
-Unlike the aforementioned methods, PEMC aligns more closely with the recently proposed Prediction-Powered Inference (PPI) approach (Angelopoulos et al.
-(2023)), particularly in its hybrid methodology that integrates valid, unbiased and classical statistical principles with modern machine learning (ML) techniques.
-This aligns with recent trends in the evaluation of generative models Zrnic and Candès (2024), such as large language models (LLMs), where fast and interpretable metrics are increasingly critical.
-In this context, PPI-type of approaches highlight the value of combining human annotations and automatic predictions to yield unbiased estimates of model performance, even in low-label settings.
-This innovation is particularly valuable in high-dimensional or irregular settings, such as LLM evaluation, where annotated data may be sparse, but unannotated or simulation-based samples are plentiful Boyeau et al.
-(2024), Eyre and Madras (2024).
-Traditional variance reduction and numerical approximation techniques have a long and established history in computational finance and stochastic modeling.
-One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by Giles (2015).
+Li et al.
+Key words: Monte Carlo methods, Exotic Options Pricing, Control Variates, Variance Swaps, Swaptions.
+1. Introduction
+Monte Carlo (MC) is a fundamental tool for evaluating complex stochastic models where closed-form solutions are unavailable.
+Its main strengths—unbiasedness and rigorous uncertainty quantification—make it essential in high-stakes, high-complexity applications like healthcare resource allocation and financial risk management, where fairness or regulatory compliance are crucial.
+For example, transparent and bias-aware methods in healthcare are vital to ensure equitable outcomes and prevent disparities in care access [110, 76, 41].
+Similarly, Monte Carlo estimation is widely used in financial risk management to implement hedging strategies, rebalance portfolios, and establish reliable error bounds, as even small biases in financial simulations can accumulate over time and lead to significant errors [30].
+Despite these advantages, classical MC faces significant challenges in nested, multilevel, or path-dependent simulations that are difficult to parallelize.
+Complex stochastic dynamics can make each sample path extremely costly to generate—sometimes taking hours for one sample (e.g., credit valuation adjustment [62])—and effective variance reduction methods, such as control variates, are often unavailable [9].
+Consequently, MC’s inherent $\mathcal{O}(1/\sqrt{n})$ convergence rate often prohibits real-time decision-making with adequate precision.
+This is also problematic in exotic option pricing, where sample paths are generated sequentially under complex dynamics, making parallelization unavailable [60].
+The issue is further exacerbated when many evaluation tasks need to be repeatedly conducted in a timely manner.
+In application domains such as option pricing in computational finance, such a computational barrier has been pointed out prominently, e.g., [73, Chapter 21] or [30, Chapter 4].
+Conversely, machine learning (ML) based predictive modeling has emerged as a fast alternative for approximating MC evaluations, including pricing financial derivatives [74, 52, 15].
+Neural networks and other ML techniques can learn complex mappings and, once trained, can generate predictions at negligible marginal cost - what practitioners often call “fast evaluation" [72].
+However, the black-box nature and the lack of inherent statistical control lead to unquantifiable biases, rendering them unsuitable for tasks requiring rigorous error quantification, e.g., high-stakes applications in risk management [74, 32].
+Thus, MC and ML represent opposite ends of the spectrum: MC is reliable and uncertainty quantifiable but intrinsically slow, while ML is computationally fast but falls short in statistical guarantees.
+Given this trade-off between the speed of ML and the reliability of MC, a natural question arises: can we design a method that combines both advantages while avoiding their respective pitfalls?
+In this paper, we answer affirmatively by introducing the Prediction-Enhanced Monte Carlo (PEMC) framework.
+PEMC leverages predictive ML modeling to replace a large portion of costly simulations with cheap/parallelizable samples from ML models, reducing variance on a scheme-wide level.
+Moreover, despite relying on black-box predictive models, PEMC maintains unbiasedness, balancing computational savings with rigorous statistical guarantees.
+1.1. Contribution
+The PEMC innovates in two aspects.
+First, PEMC integrates ML predictive models into any MC baseline while preserving unbiasedness and error quantification of MC.
+Second, we show how the predictive modeling of ML combined with cheap, parallelizable samples will result in a scheme-wide reduction of variance and costly samples, enhancing real-time efficiency for complex MC simulations.
+PEMC method can also be viewed as a modernized variant of the control variate (CV) technique, a classical approach that reduces Monte Carlo variance via auxiliary outputs whose known mean offsets random fluctuations [60, 9].
+Unfortunately, the classical CV requirement of a closed-form mean of the a correlated auxiliary output easily becomes prohibitive in many modern applications or complicated stochastic systems.
+In such cases, there simply is no readily available analytic expression for the auxiliary variable’s expectation, effectively barring the use of a CV.
+The PEMC framework bridges the above gap and substantially enlarge the scope of CV to many ML candidates.
+The key idea is the following: In the traditional CV framework, the known mean of CV is used to achieve a per-replication variance reduction while maintaining unbiasedness, i.e., the variance is reduced even if we run just one simulation replication.
+This is a powerful ability, but as mentioned above it is achieved at the expense of severely narrowing the applicability of CV by requiring closed-form mean function.
+However, if the CV itself is computationally cheap in the sense of being estimatable with fast simulation, then we can afford to run plenty of additional MC separately from the original MC to estimate its mean.
+This would preserve unbiasedness and overall estimation variance relative to the total computation cost, we could achieve a scheme-wide efficiency improvement.
+That is, in the PEMC framework, we consider scheme-wide instead of per-replication variance reduction, through which we relax the crucial requirement of known CV mean, and this in turn allows us to substantially expand the CV methodology to cover modern ML models.
+The alignment of our framework with the use of ML predictions prompts us to regard PEMC as a modernized view of CV.
+We demonstrate PEMC’s efficacy across multiple MC applications.
+First, we consider two exotic financial derivative-pricing scenarios: (i) variance swap pricing under a stochastic local volatility model, a fundamental instrument for volatility trading due to its payout on realized variance; and (ii) swaption pricing within the Heath-Jarrow-Morton (HJM) interest rate framework, a sophisticated yet widely used derivative by financial institutions for hedging and speculation.
+Both instruments are highly liquid, exchange-traded products but pose significant computational challenges due to the infeasibility of traditional variance reduction techniques, as they lack analytic or easily computable expected values.
+Additionally, we illustrate PEMC’s broad applicability through an ambulance dispatch and hospital load-balancing simulation, where unbiased mortality rate estimates critically inform ethical and efficient resource allocation.
+These examples collectively highlight PEMC’s capability to leverage machine learning predictions to effectively manage computational complexity while maintaining unbiased estimation and reliable error quantification.
+1.2. Literature Review
+The literature on enhancing Monte Carlo (MC) simulations with machine learning (ML) approximators is extensive, yet many existing ML-based methods for control variates (CV) and variance reduction impose restrictive conditions that limit their practical applicability.
+Approaches such as reproducing Stein kernels [100, 88], Neural CV derived from Stein’s identity [129, 82], regularized least squares for CV construction [106, 121, 85], adaptive CV schemes [66, 67, 79], and $\mathscr{L}^{2}$ function approximation frameworks [93] have demonstrated theoretical promise with potentially rapid convergence rates.
+Recent theoretical work by [18] provides important insights into when regression-adjusted control variates can achieve optimal performance, showing that their effectiveness depends critically on function smoothness.
+However, these methods typically require knowledge of the control variate mean or impose specific structural assumptions that can be difficult to check.
+As a result, while these existing methods offer strong theoretical foundations, their applicability in complex financial modeling scenarios can be sometimes limited.
+Multilevel Monte Carlo (MLMC) [58], another prominent variance reduction approach, has shown effectiveness through coupling simulations at different discretization levels but is often challenging to implement, requiring careful design for stable coupling [107, 19, 112, 57, 115].
+Our Prediction-Enhanced Monte Carlo (PEMC) framework addresses these limitations comprehensively by eliminating the necessity of known CV means, enabling the direct use of modern ML predictors as control variates, while rigorously preserving unbiasedness.
+In a different line of literature, the work of Emsermann and Simon [49] on quasi control variates and the work of [102] on control variates using Estimated Means (CVEMs) share notable similarities with our approach, as it relies on an estimated mean (rather than a priori known, and thus not a traditional control variate) of a quantity that is expected to exhibit a strong correlation with the random variable of interest.
+However, our methodology distinguishes itself in several critical ways.
+First, our primary motivation centers on integrating modern ML capabilities to train a suitable control variate, rather than relying on a fixed quantity observed to be correlated with the target (which may not always be obvious or even available, especially for our options pricing examples).
+In other words, we can create an effective control variate from an easily estimable quantity, even if its naive form does not exhibit strong correlation with the target.
+This approach circumvents the challenges of identifying a closed-form mean and strictly relying on high correlation-based pairing.
+Second, instead of adopting an adaptive algorithm as they do, we focus on a pre-training approach, demonstrating that our performance advantages would not emerge under their adaptive scheme.
+Third, our analysis is not anchored exclusively in asymptotic arguments.
+We provide a formulation that can be interpreted in both finite-sample and asymptotic regimes, offering more direct and flexible theoretical guarantees.
+In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively even when auxiliary means are unknown.
+Unlike previously discussed methods, PEMC aligns closely with the recently proposed Prediction-Powered Inference (PPI) approach [6], particularly in its hybrid methodology that integrates classical, unbiased statistical principles with modern machine learning (ML) techniques.
+This hybrid strategy is becoming increasingly important in evaluating generative models such as large language models (LLMs), where rapid, interpretable, and unbiased metrics are essential [135, 22, 51].
+PPI-type approaches combine sparse human annotations with abundant ML-generated labels, facilitating accurate model performance evaluation even in low-label scenarios.
+Additionally, these methodologies intersect with a rich literature in causal inference, specifically doubly robust estimators [10, 55, 113] and orthogonal statistical learning [36, 54].
+Doubly robust methods construct additional robustness layers by pairing sophisticated predictive models with auxiliary estimators, ensuring that minor errors in one model component do not significantly bias the overall estimation.
+This philosophy closely mirrors the logic of control variates, and by extension, the core principles of our PEMC framework.
+Traditional variance reduction and numerical approximation techniques also have a long and established history in computational finance and stochastic modeling.
+One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by [58].
 MLMC exploits a hierarchy of discretizations (e.g., finer and coarser time steps in an SDE simulation) to reduce variance at a fraction of the cost of a naive Monte Carlo approach.
@@
 Despite its solid theoretical foundations, MLMC can be challenging to implement in practice: constructing suitable couplings between the fine and coarse approximations, ensuring stability, and tuning parameters to achieve the promised complexity gains are all nontrivial tasks.
-However, adopting a coupling perspective provides valuable insights into how PEMC manages the relationship between the target and the predictor.
-The advent of machine learning (ML) has spurred a wave of research aiming to accelerate and improve financial computation, including option pricing.
-Techniques such as neural network surrogates can approximate pricing functions or Greeks, drastically cutting down on runtime (Hutchinson et al.
-(1994), Buehler et al.
-(2019)).
-However, a major drawback is the lack of rigorous error controls and biases introduced during training.
-Our paper is also related to the recent stream of literature in causal inference, where methods aim to mitigate challenges suc as selection bias and confounding.
-Doubly robust estimators serve as a key example (Bang and Robins (2005)).
-These estimators maintain consistency even if only one of two specified models—either the outcome model or the propensity score model—is correctly specified.
-Their inherent idea of robustness through additional predictor resonates with the philosophy behind control variates, as well as PEMC, where pairing a complex estimator with a suitably chosen secondary construct can yield improved stability and reduced variance.
+Nevertheless, the class of MLMC methods and their adoption of a coupling perspective provide extremely valuable insights into how PEMC manages the relationship between the target and the predictor [107, 19, 112, 57, 115].
+In computational finance, the emergence of machine learning (ML) has also sparked significant research efforts aimed at enhancing the speed and accuracy of financial computations, including option pricing.
+By using techniques such as neural network surrogates, these methods can approximate pricing functions or the Greeks at a fraction of the computational cost of traditional numerical schemes [74, 32, 119, 114, 84].
+Early work by hutchinson1994nonparametric [74], for instance, used neural networks to relax stringent modeling assumptions, while more recent approaches [32, 119, 114, 84] demonstrate that deep architectures can effectively handle high-dimensional and path-dependent pricing problems.
+However, these purely data-driven approximations typically lack the rigorous error controls afforded by more classical methods.
+As a result, biases introduced during training and potential overfitting can lead to systematic mispricings, particularly under market regimes not well-represented in the training data.
+Such issues underscore the need for frameworks that combine the computational advantages of ML with robust theoretical guarantees to ensure reliable and unbiased pricing outcomes.
+On the other hand, within option pricing, even though both MC and PDE-based methods are widely used, it is known that PDE methods generally cannot be used to price options (i.e., payoff depends on the price path).
+While special cases allow for PDE formulations through state-space augmentation or boundary condition adjustments, these remain as exceptions rather than the rule [116].
 The rest of the paper is organized as follows.
@@
 In Section 4, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical illustrations.
-In Section 4, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework.
+In Section 5, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the HJM framework.
 We conclude the paper in Section 6 with discussion on broader implications and potential directions for future research.
-2 Prediction-Enhanced Monte Carlo: Framework
-In this section, we describe how to apply the “Prediction-Enhanced Monte Carlo” Framework for a general simulation evaluation.
-While this framework has broad applicability, for clarity and concreteness, we will present it in the context of options pricing.
-This choice of notation and presentation aligns with our subsequent examples and reflects a significant area of application.
-However, it’s important to note the principles and methodology we discuss here can be readily adapted to a wide range of problems beyond financial derivatives.
-To effectively illustrate our proposed framework, facilitate a comprehensive comparison with traditional methods, and enhance the reader’s understanding, we employ Asian option pricing as a running example throughout in this section.
+2. Prediction-Enhanced Monte Carlo
+We provide an overview and motivational explanation of our PEMC framework. Suppose we want to produce an unbiased evaluation of the quantity
+$\mu(\boldsymbol{\theta})=\mathbb{E}_{\boldsymbol{\theta}}[f_{\boldsymbol{\theta}}(\boldsymbol{Y})]$ (1)
+for arbitrary, potentially high-dimensional parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{d_{\theta}}$, where $\Theta$ is the parameter space.
+Here, $\boldsymbol{Y}\in\mathbb{R}^{d_{y}}$ represents a high-dimensional random vector governed by $\mathbb{E}_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}}$ is the evaluation function $\mathbb{R}^{d_{y}}\to\mathbb{R}$.
+In general, both $f_{\boldsymbol{\theta}}$ and $\mathbb{E}_{\boldsymbol{\theta}}$ can depend on $\boldsymbol{\theta}$.
+The goal is to produce an unbiased estimate of $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, while the primary bottleneck stems from costly and time-consuming simulation of $\boldsymbol{Y}$.
+When $\boldsymbol{Y}$ can be simulated and $f$ evaluated, then standard MC amounts to generating many $f(\boldsymbol{Y}_{i})$, $i=1,\ldots,n$ and outputting their average.
+While this approach is demanding when $f$ and $\boldsymbol{Y}$ are expensive, it is unbiased and its statistical error can be quantified straightforwardly via standard confidence intervals.
+To reduce computation, one can approximate $\mu(\boldsymbol{\theta})$ with an ML regression model, say $\hat{\mu}(\boldsymbol{\theta})$, which can be evaluated more cheaply than simulating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$’s.
+This natural approach, however, would face challenges at multiple levels when reliability and error quantification is important: From a theoretical standpoint, ML generalization bounds typically assess training errors only within the model class, i.e., $\hat{\mu}(\boldsymbol{\theta})-\mu^{*}(\boldsymbol{\theta})$ where $\mu^{*}(\cdot)$ is the best in-class model that can be different from $\mu(\cdot)$.
+Moreover, even assuming $\mu^{*}(\cdot)$ is a “rich" class, these bounds still only assess errors in an aggregate, worst-case fashion, i.e., high-probability bounds treating $\boldsymbol{\theta}$ as a random covariate.
+On the practical front, training-testing split or cross-validation succumbs to similar issues, in that it cannot measure bias against $\mu(\cdot)$, nor at the level of each $\boldsymbol{\theta}$.
+That is, in terms of reliability and uncertainty handling, MC appears superior by providing accurate, straightforward and instance-specific error assessment on each estimate of $\mu(\boldsymbol{\theta})$.
+None of these advantages are available for ML despite its computational edge.
+An idea to leverage the respective advantages of MC and ML is to use control variate (CV), a classical variance reduction method.
+CV aims to reduce the error of each MC run, consequently the required replication size, by utilizing information from “auxiliary” simulation outputs.
+Specifically, in generating $f(\boldsymbol{Y}_{i})$, suppose we can also generate auxiliary output $\boldsymbol{X}_{i}$ and evaluate $g(\boldsymbol{X}_{i})$ for some function $g$.
+Then we can use $f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i})+\mathbb{E}[g(\boldsymbol{X})]$ as an unbiased replicate for $\mathbb{E}[f(\boldsymbol{Y})]$.
+In particular, if $f(\boldsymbol{Y})$ and $g(\boldsymbol{X})$ are highly correlated, this replicate will have a smaller variance than $f(\boldsymbol{Y}_{i})$.
+Naturally, one can think of using ML to train a good predictor $g(\boldsymbol{X})$ for $f(\boldsymbol{Y})$, which would be highly correlated with $f(\boldsymbol{Y})$ and serve as an effective CV.
+However, taking aside the question of how to construct $\boldsymbol{X}$ and $g$, the CV method requires knowledge of $\mathbb{E}[g(\boldsymbol{X})]$, the closed-from mean on the auxiliary output.
+This knowledge is the key convertible into variance reduction.
+Yet, it is seldom available and has apparently severely limited the historical applicability of CV [61] and arguably even the more recent ML-inspired approaches [129, 121, 88].
+We introduce Prediction-Enhanced Monte Carlo (PEMC), a framework that fulfills these objectives.
+PEMC leverages the classical control variate idea but generalizes its variance-reduction notion to a scheme-wide, modernized view, allowing for effective integration of MC with modern ML.
+Concretely, PEMC aims to estimate (1) by
+$PEMC(\boldsymbol{\theta}):=\frac{1}{n}\sum_{i=1}^{n}\big{(}f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})\big{)}+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\theta},\boldsymbol{\tilde{X}}_{j}).$ (2)
+Here, $\boldsymbol{X}$ is an auxiliary variable, coupled with $\boldsymbol{Y}$, that serves as features for predicting $f(\boldsymbol{Y})$ and is also chosen to have a marginal distribution that is both efficiently simulatable and highly parallelizable, allowing large number $N$ independent samples of $\tilde{\boldsymbol{X}}$ with the same marginal as $\boldsymbol{X}$ to be generated at low cost.
+The function $g$ is a prediction model trained via squared-error loss using MC sample under $(\boldsymbol{X},\boldsymbol{Y})\sim\mathbb{P}_{\boldsymbol{\theta}}$ for various $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, which aims to approximate the function $g\bigl{(}\boldsymbol{\theta},\boldsymbol{X}\bigr{)}\approx\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}f_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{Y}\bigr{)}\big{|}\boldsymbol{\theta},\boldsymbol{X}\bigr{]}$.
+As we will see, the selection of an appropriate $\boldsymbol{X}$ is a pivotal component of PEMC, analogous to feature selection in common ML tasks.
+The key lies in selecting $\boldsymbol{X}$ to be both a feature for predicting $f(\boldsymbol{Y})$ and an efficiently simulatable, parallelizable random variable.
+Finding a good $\boldsymbol{X}$ often requires domain knowledge.
+In particular, given the whole path $\boldsymbol{Y}$, one can choose $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ to be some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ (e.g., ingredients for constructing $\boldsymbol{Y}$).
+For instance, for SDEs, one can define $\boldsymbol{X}$ as the sum of driving Brownian increments that generate $\boldsymbol{Y}$.
+This choice could capture a substantial portion of the SDE’s variance, making $\boldsymbol{X}$ a good feature for predicting $f(\boldsymbol{Y})$, while its Gaussianity allows it to be generated cheaply and in parallel.
+We explore this in detail in subsequent sections.
+Many implementation details remain unspecified in the basic description above.
+For instance, how exactly to train g, how to select the sample sizes $n$ and $N$, and how to adapt PEMC to a conditional setting where $\theta\mapsto\boldsymbol{X}(\theta)$, $\boldsymbol{Y}(\theta)$, and $\mathbb{E}_{\theta}$ all depend on some parameter $\theta$.
+The next sections will address these questions in depth, providing both theoretical justification and practical guidelines for deploying PEMC in diverse scenarios.
+3. Prediction-Enhanced Monte Carlo: Framework via Example
+In this section, we employ Asian option pricing as a running example on how to apply PEMC.
 Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period.
-As mentioned in Boyle et al.
-(1997) and Dingec et al.
-(2015), closed-form solutions for the pricing of arithmetic average Asian options are generally not available, but arithmetic average Asian options serve as an exemplary case where control variate methods, based on closed-form pricing of the geometric average Asian option under well-known models, can drastically reduce variance and enhance the efficiency of Monte Carlo simulations.
-However, as we will demonstrate, our new framework achieves the same level of variance reduction without requiring explicit knowledge of the construction of a control variate or its closed-form mean.
-This significantly expands the applicability and effectiveness of control variate techniques in arbitrarily complex models.
-2.1 Problem Setup
+As mentioned in [23] and [44], closed-form solutions for the pricing of arithmetic average Asian options are generally not available.
+3.1. Problem Setup
 The quantity of interest is an evaluation that can be expressed via:
-Option $\text{Option Price}=\mathbb{E}_{\text{general risk neutral measure}}[f_{\text{payoff}}(\bm{Y})].$ (1)
-Here $\bm{Y}$ represents a potentially high-dimensional random quantity that encapsulated all the randomness generated under a general, accessible sampling mechanism, denoted here by a general risk neutral measure $\mathbb{E}$.
-For example, if one is required to generate a discretized sample path of a certain stochastic differential equation (SDE) using some numerical scheme, $\bm{Y}$ then represents a complete sample path generated and the measure induced by that discretized SDE would be the general risk neutral measure.
-Typically, the more complex the simulation scheme is, the higher the dimension of $\bm{Y}$ will be, as more sophisticated simulation scheme often requires more intermediate steps or additional variables (e.g., an order of $O(T/\Delta t)$, when $T$ is the time horizon and $\Delta t$ is the discretization time step).
-Then, $f_{\text{payoff}}$, denoted by the term payoff function, represents the proper function (functional) that transform or summarize $\bm{Y}$ into a sample for the quantity of interest.
-The motivation for our framework arises, when the complexity and generality of the payoff function $f_{\text{payoff}}$ leads to a highly complicated dependence on the entirety of the random variables $\bm{Y}$.
-This complication may stem from factors such as path dependency, nested or multilevel simulations, or intricate interactions within the model.
-Consequently, drawing valid samples from the distribution of $f_{\text{payoff}}(\bm{Y})$ becomes both challenging and computationally intensive.
-To fully specify the framework, we need to define the several parameters governing both the risk-neutral measure and the payoff function. We formally list the key definitions as:
-  1. 1.
-    Model Parameters $\bm{\theta}_{\text{model}}$: These are the parameters that specify the simulation or the stochastic model of the underlying process $\bm{Y}$ (e.g., asset value time series).
-For instance, in a Heston model (Heston 1993), the price process $S_{t}$ and the instantaneous volatility process $\nu_{t}$ are jointly modeled:
+Option $\text{Option Price}(\boldsymbol{\theta})=\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}[f_{\text{payoff}(\boldsymbol{\theta})}(\boldsymbol{Y})],$ (3)
+For Asian option, we can specify the framework as:
+  1. (1)
+    Model Parameters $\boldsymbol{\theta}_{\text{model}}$: These are the parameters that specify the simulation or the stochastic model of the underlying process $\boldsymbol{Y}$ (e.g., asset value time series).
+For instance, in a Heston model [69], the price process $S_{t}$ and the instantaneous volatility process $\nu_{t}$ are jointly modeled:
     $dS_{t}=$ $rS_{t}dt+\sqrt{\nu}_{t}S_{t}dW_{t}^{S}$
     $d{\nu}_{t}=$ $\kappa(\eta-\nu_{t})dt+\delta\sqrt{\nu_{t}}dW_{t}^{\nu}$ (4)
-    The model parameter $\bm{\theta}_{\text{model}}$ is thus 5 dimensional $\bm{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$.
+    The model parameter $\boldsymbol{\theta}_{\text{model}}$ is thus 5 dimensional $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$.
 It specifies the risk neutral drift $r$, the long-run average variance $\eta$, the volatility of volatility $\delta$, the correlation $\rho$ of between $W^{S}$ and $W^{\nu}$, and the mean reversion rate $\kappa$.
-In options pricing, these parameters are typically calibrated to fit market data and then used to generate sample paths $\bm{Y}:=\{S_{t}\}_{t}$.
-Thus $\bm{\theta}_{\text{model}}$ governs and describes $\mathbb{E}_{\text{ risk neutral measure}}$ that generates $\bm{Y}$.
-  2. 2.
-    Simulation Parameters $\bm{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$.
-The distinction between $\bm{\theta}_{\text{simulation}}$ and $\bm{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are not calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed).
-For example, when one simulates the Heston model above using Euler discretization scheme, [^1] As mentioned in Dingec et al.
-(2015), Euler scheme was the primary method for simulation of Heston model until the exact simulation of Broadie and Kaya (2006) is developed.
-However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential (Andersen 2008) or Euler scheme remains prevalent (see, e.g., Lord et al.
-(2010)).
-we use $\bm{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$.
+In options pricing, these parameters are typically calibrated to fit market data and then used to generate sample paths $\boldsymbol{Y}:=\{S_{t}\}_{t}$.
+Thus $\boldsymbol{\theta}_{\text{model}}$ governs and describes $\mathbb{E}_{\text{ risk neutral measure}}$ that generates $\boldsymbol{Y}$.
+  2. (2)
+    Simulation Parameters $\boldsymbol{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$.
+The distinction between $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed).
+For example, when one simulates the Heston model above using Euler discretization scheme, [^1] As mentioned in [44], Euler scheme was the primary method for simulation of Heston model until the exact simulation of [31] is developed.
+However, due to its significant computational demands, the exact method is less commonly used in practice and simpler, biased simulation method such as Quadratic Exponential [3] or Euler scheme remains prevalent (see, e.g., [91]).
+we use $\boldsymbol{\theta}_{\text{simulation}}$ to specify the initial stock price $S_{0}$, the initial volatility $\nu_{0}$, the time horizon $T$, and discretization time step $\Delta t$.
 The choice of these parameters can impact both the accuracy and efficiency of the simulation.
 While not typically part of the model calibration process, these parameters also play a crucial role in describing risk-neutral measure.
-  3. 3.
-    Payoff Function Parameters $\bm{\theta_{\text{payoff}}}$: This parameter $\bm{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\bm{\theta}_{\text{payoff}}}$.
+  3. (3)
+    Payoff Function Parameters $\boldsymbol{\theta_{\text{payoff}}}$: This parameter $\boldsymbol{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\boldsymbol{\theta}_{\text{payoff}}}$.
 In options pricing, this specifies the details of the contract.
-For example, given the full price path $\bm{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
+For example, given the full price path $\boldsymbol{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
     $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
-    where $\bm{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
+    where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
 Notice we have simply omitted the discount factor $e^{-rT}$.
-Under these definitions, we can summarize the standard Monte Carlo approach as follows: one calibrates $\bm{\theta}_{\text{model}}$, specify $\bm{\theta}_{\text{simulation}}$ and generate $n$ independent paths $\{\bm{Y}_{i}\}_{i\in[n]}$ according to $\mathbb{E}_{\text{ risk neutral measure}}$ given by $\bm{\theta}_{\text{model}}$ and $\bm{\theta}_{\text{simulation}}$.
-Finally, one averages $\frac{1}{n}\sum_{i}f_{\bm{\theta}_{\text{payoff}}}(\bm{Y}_{i})$.
-Remark Having specified the distinct roles of $\bm{\theta}_{\text{model}}$, $\bm{\theta}_{\text{simulation}}$ and $\bm{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\bm{\theta}$ when no ambiguity arises.
-2.2 Prediction-Enhanced Monte Carlo: Motivation
-As we have seen, Asian options are path-dependent derivative securities whose payoffs depend on the average prices of the underlying asset.
-Closed-form solutions for their prices are generally unavailable, numerous studies have proposed new numerical methods for pricing them (see Dingec et al.
-(2015)).
-While Partial Differential Equation (PDE)-based methods and Monte Carlo simulation are both generic methods for pricing options, applying PDE methods to particular models often requires special effort, primarily due to the curse of dimensionality.
-Moreover, for our discussion here, it has been noted that no PDE method performs well for Asian options under the Heston model Dingec et al.
-(2015).
-Therefore, across various models, simulation becomes the common choice.
-The principal drawback of simulation, as strongly highlighted in the literature, is its slow convergence rate.
-Employing variance reduction methods is therefore essential to improve the precision of simulation results.
-However, designing a successful variance reduction method requires exploiting the special structures of the derivative security and the underlying model.
-As a result, control variate (CV) techniques are often not available for pricing options under most of the complicated models used in practice.
-In fact, the pricing of arithmetic Asian option under Black–Scholes geometric Brownian motion (GBM) model is a main exception where the closed-form expression of the geometric Asian option allows one to effectively deploy control variate techniques.
-However, as soon as one extends beyond the most basic GBM model to more complex models such as the Heston model, the closed-form solutions for the geometric Asian options are no longer available and the applicability of control variate in these contexts is severely limited (see Dingec et al.
-(2015)).
-More generally in the context of options pricing, while widely used, Monte Carlo methods face challenges when applied to complex, path-dependent derivatives.
-In such simulations, the state at each time step is determined by the state at the previous step, creating a sequential dependency that precludes straightforward parallelization or speedup.
-Moreover, for the complex risk neutral measures or the irregular payoff functions encountered in many exotic options pricing problems, effective control variate or variance reduction techniques are frequently unavailable.
-As a result, the slow convergence of Monte Carlo - its $O(1/\sqrt{n})$ error after $n$ replications - becomes an acute bottleneck and achieving high accuracy often implies a undesirable large number of replications in practice.
-To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework is designed to provide a robust and easily accessible mechanism for designing control variates.
-This novel approach aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable.
-Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that Glasserman (2013b) highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
-2.3 Prediction-Enhanced Monte Carlo: Prediction Model
-In this section, we present our PEMC framework, which is centered around an efficient prediction model.
-In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is essentially compatible with any black-box model.
-However, as we will demonstrate, the squared error loss—a function that is both convex and differentiable—combined with the need for a flexible model family capable of expressing complex dependencies and being straightforward to train using off-the-shelf, easily accessible solvers, makes neural networks an ideal choice.
-When applying the PEMC framework, an efficient prediction model refers to one that can process highly parallelizable and computationally efficient input $\bm{X}$, which serves as a strong predictor or good feature Hastie et al.
-(2009) for predicting $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$.
-In particular, given the whole path $\bm{Y}$, we want to find some input feature $\bm{X}:=\phi(\bm{Y})$ that is some typically low-dimensional transformation $\phi$ of $\bm{Y}$ with:
-  • Property 1: $\bm{X}$ can be regarded as a reasonable predictor for $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$.
-For example, if $\bm{X}$ exhibits a significant correlation with $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$, it may serve as the basis for a reasonable linear regressor, although our framework far exceeds linear predictors.
-  • Property 2: While the the joint distribution of $d$ is hard to obtain, the marginal distribution of $\bm{X}$ allows for efficient and parallelizable simulation.
-The selection of an appropriate $\bm{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks.
-This process necessitates domain expertise and a thoughtful evaluation of the specific financial instrument and underlying model.
-The challenge lies in striking a balance between the predictive capability of $\bm{X}$ for $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$ and its computational efficiency.
-In subsequent sections, we will explore the intricacies of this selection process in greater detail, examining its impact on the performance of the PEMC framework through illustrative examples.
-For now, we proceed under the assumption that a suitable $\bm{X}$ has been identified.
-We now focus on the pre-training phase of our prediction model.
-Unlike many machine learning or deep learning tasks, where the quantity and quality of data are critical challenges—and often a core difficulty due to the expense or scarcity of labeled data—in the Monte Carlo setting, we benefit from an effectively unlimited supply of clean, labeled data that can be generated as needed.
-The data generation and model training process can be outlined through the following key steps:
-  1. 1.
-    We first define a parameter space $\bm{\Theta}$ that encompasses all combinations of $\bm{\theta}:=(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}})$ that are of practical interest.
-This space should cover the range of parameters that would likely be encountered in real-world pricing scenarios on a daily, weekly or monthly basis, depending on the model update frequency.
-  2. 2.
-    We then draw samples of $\bm{\theta}_{i}$ from $\bm{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage.
-For each sampled parameter set, we generate one pair  [^2] You could also generate multiple pair of $(\bm{Y}(\bm{\theta}_{i}),\bm{X}(\bm{\theta}_{i}))$ under the same sampled $\bm{\theta}_{i}$.
-of $\bm{Y}(\bm{\theta}_{i})$ and its corresponding $\bm{X}(\bm{\theta}_{i}):=\phi(\bm{Y}(\bm{\theta}_{i}))$.
-It’s important to note that $\bm{X}$ and $\bm{Y}$ are coupled in this generation process.
-  3. 3.
-    For each generated pair, we compute the payoff $f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$, which serves as a label in our training data.
-The corresponding feature vector is comprised of $(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$.
-  4. 4.
+Note that the discrete sampling defied by the payoff does not necessarily determine the Euler discretization in time and in fact the latter is usually much more frequent for accuracy.
+Remark Having specified the distinct roles of $\boldsymbol{\theta}_{\text{model}}$, $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\boldsymbol{\theta}$ when no ambiguity arises.
+3.2. Prediction-Enhanced Monte Carlo: Motivation
+As we have seen, Asian options are path-dependent derivatives whose payoffs depend on the average prices of the underlying asset.
+Generally, closed-form solutions for these options are unavailable, prompting numerous studies to propose various numerical methods for their pricing.
+In particular, the arithmetic Asian option under the geometric Brownian motion (GBM) model is one of the rare cases where a closed-form geometric Asian option price exists and can serve as a control variate.
+Beyond simple models like GBM, closed-form CV solutions are unavailable [44].
+To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable.
+Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that [60] highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
+In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is compatible with any black-box model.
+However, we want to find some input feature $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ that is some typically low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ with:
+  • Property 1: $\boldsymbol{X}$ can be regarded as a reasonable predictor for $f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y})$.
+  • Property 2: In contrast to that of $\boldsymbol{Y}$, the marginal distribution of $\boldsymbol{X}$ allows for efficient and parallelizable simulation.
+The selection of an appropriate $\boldsymbol{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks.
+This process necessitates domain expertise.
+For now, we proceed under the assumption that a suitable $\boldsymbol{X}$ has been identified.
+3.3. Prediction-Enhanced Monte Carlo: Prediction Model
+We now focus on the pre-training phase of our prediction model. The data generation and model training process can be outlined through the following key steps:
+  1. (1)
+    We first define a parameter space $\boldsymbol{\Theta}$ that encompasses all combinations of $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})$ that are of practical interest.
+This space covers the range of parameters that would likely be encountered in real-world pricing scenarios on a daily, weekly or monthly basis, depending on the model update frequency.
+  2. (2)
+    We then draw samples of $\boldsymbol{\theta}_{i}$ from $\boldsymbol{\Theta}$, using uniform sampling or any measure that ensures comprehensive coverage.
+For each sampled parameter set, we generate one pair  [^2] You could also generate multiple pair of $(\boldsymbol{Y}(\boldsymbol{\theta}_{i}),\boldsymbol{X}(\boldsymbol{\theta}_{i}))$ under the same sampled $\boldsymbol{\theta}_{i}$.
+of $\boldsymbol{Y}(\boldsymbol{\theta}_{i})$ and its corresponding $\boldsymbol{X}(\boldsymbol{\theta}_{i}):=\phi(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$.
+It’s important to note that $\boldsymbol{X}$ and $\boldsymbol{Y}$ are coupled in this generation process.
+  3. (3)
+    For each generated pair, we compute the payoff $f_{\boldsymbol{\theta}_{i,\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$, which serves as a label in our training data.
+The corresponding feature vector is comprised of $(\boldsymbol{\theta}_{i,\text{model}},\boldsymbol{\theta}_{i,\text{simulation}},\boldsymbol{\theta}_{i,\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}_{i}))$.
+  4. (4)
     This process is repeated $N_{\text{train}}$ times to generate training dataset which consists of $N_{\text{train}}$ pairs of
-    $\textit{label}_{i}:=$ $(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$
-    $\textit{feature}_{i}:=$ $f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$
+    $\textit{feature}_{i}:=$ $(\boldsymbol{\theta}_{i,\text{model}},\boldsymbol{\theta}_{i,\text{simulation}},\boldsymbol{\theta}_{i,\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}_{i}))$
+    $\textit{label}_{i}:=$ $f_{\boldsymbol{\theta}_{i,\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$
     for $i\in[N_{\text{train}}]$.
-  5. 5.
-    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and save the resulting model ${g}$:
+  5. (5)
+    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and saved:
     $\min_{g}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-g(\text{feature}_{i}))^{2}.$ (5)
-1:procedure DataGeneration($N_{\text{train}},\bm{\Theta}$)
+1:procedure DataGeneration($N_{\text{train}},\boldsymbol{\Theta}$)
 […]
 3:     for $i=1$ to $N_{\text{train}}$ do
-4:         Sample parameters $\bm{\theta}_{i}\sim\Theta$ uniformly
-5:         Generate $\bm{X}(\bm{\theta}_{i}),\bm{Y}(\bm{\theta}_{i})\sim\mathbb{E}_{\text{ risk neutral measure}(\bm{\theta}_{i})}$
-6:         $\textit{label}_{i}\leftarrow f_{\bm{\theta}_{i,\text{payoff}}}(\bm{Y}(\bm{\theta}_{i}))$
-7:         $\textit{feature}_{i}\leftarrow(\bm{\theta}_{i,\text{model}},\bm{\theta}_{i,\text{simulation}},\bm{\theta}_{i,\text{payoff}},\bm{X}(\bm{\theta}_{i}))$
+4:         Sample parameters $\boldsymbol{\theta}_{i}\sim\Theta$ uniformly
+5:         Generate $\boldsymbol{X}(\boldsymbol{\theta}_{i}),\boldsymbol{Y}(\boldsymbol{\theta}_{i})\sim\mathbb{E}_{\text{ risk neutral measure}(\boldsymbol{\theta}_{i})}$
+6:         $\textit{label}_{i}\leftarrow f_{\boldsymbol{\theta}_{i,\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$
+7:         $\textit{feature}_{i}\leftarrow(\boldsymbol{\theta}_{i,\text{model}},\boldsymbol{\theta}_{i,\text{simulation}},\boldsymbol{\theta}_{i,\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}_{i}))$
 […]
@@
 The essential steps are summarized in Algorithm 1.
-While the data generation process can be time-intensive due to the complexity of $\bm{Y}$, it is conducted offline during the pre-training phase.
+While the data generation process can be time-intensive due to the complexity of $\boldsymbol{Y}$, it is conducted offline during the pre-training phase.
 This approach minimizes its impact on real-time pricing applications by effectively trading memory usage for computational speed during execution.
 This pre-training setup presents an ideal scenario for machine learning applications.
-Unlike many real-world problems where data scarcity is a concern, access to the data-generating process in PEMC allows for the creation of virtually unlimited, clean data.
-This abundance greatly facilitates the training of complex models.
 It is important to also note that PEMC does not require $g$ to take specific forms.
 However, the convex and differentiable nature of the loss function (i.e., squared loss), the large supply of data (i.e., from Monte Carlo simulation), and the need for expressiveness of the model family make neural networks (NNs) an ideal choice.
-When training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer Kingma and Ba (2014), batch normalization Ioffe and Szegedy (2015), and skip connections He et al.
-(2016)—can be employed to improve performance.
+For training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer [80], batch normalization [75], and skip connections [64]—can be employed to improve performance.
 In fact, recent advancements in machine learning technologies have made the training process remarkably accessible and efficient.
-The availability of open-source, user-friendly frameworks like PyTorch Paszke et al.
-(2019), TensorFlow Abadi et al.
-(2016), and JAX Bradbury et al.
-(2018)—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets.
-These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) LeCun et al.
-(1998), Residual Networks (ResNets) He et al.
-(2016), and Transformers Vaswani et al.
-(2017), as well as various Stochastic Gradient Descent (SGD) optimization algorithms Bottou (2010).
+The availability of open-source, user-friendly frameworks like PyTorch [103], TensorFlow [1], and JAX [24]—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets.
+These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) [83], Residual Networks (ResNets) [64], and Transformers [128], as well as various Stochastic Gradient Descent (SGD) optimization algorithms [21].
 In applications, we only utilized well-established and widely recognized neural network structures that are now considered common.
@@
 Finally, regarding the choice loss function, we note that the goal of the prediction model $g$ is to produce an approximation of the underlying regression function:
-$g(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X}(\bm{\theta}))$
-$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}f_{\bm{\theta}_{\text{payoff}}}(\bm{Y}(\bm{\theta}))\bigg{|}(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X}(\bm{\theta}))\bigg{]}$
+$g(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))$
+$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))\bigg{|}(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}$
 As we will see, the quality of this approximation will impact the effectiveness of variance reduction, which we shall explore in detail in subsequent sections.
-2.4 Prediction-Enhanced Monte Carlo: Evaluation
+3.4. Prediction-Enhanced Monte Carlo: Evaluation
 Having completed the training phase of PEMC, we now turn our attention to the evaluation procedure.
@@
 The evaluation process proceeds as follows:
-  1. 1.
-    A specific parameter set $\bm{\theta}:=(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}})\in\bm{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing.
-These parameters can be determined through combination of calibration, customization, client agreement or external observation.
-  2. 2.
-    Generate $n$ pairs of $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$.
-Here we have suppressed $\bm{Y}$’s dependency on $\bm{\theta}$, as in the evaluation stage only a fixed $\bm{\theta}$ is considered.
-  3. 3.
-    Based on the marginal distribution of $\bm{X}$, independently generate $N$ additional samples of $(\bm{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$, directly according to its marginal distribution.
-Importantly, these samples are independent of the previous data $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$.
-  4. 4.
+  1. (1)
+    A specific parameter $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})\in\boldsymbol{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing.
+  2. (2)
+    Generate $n$ pairs of $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$.
+Here we have suppressed $\boldsymbol{Y}$’s dependency on $\boldsymbol{\theta}$, as in the evaluation stage only one fixed $\boldsymbol{\theta}$ is considered.
+  3. (3)
+    Based on the marginal distribution of $\boldsymbol{X}$, independently generate $N$ additional samples of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$, directly according to its marginal distribution.
+Importantly, these samples are independent of the previous data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$.
+  4. (4)
     Utilize the pre-trained model ${g}$ to evaluate the PEMC estimator:
-    $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$ (4)
-For notational simplicity, we have written ${g}(\bm{X})$ instead of ${g}(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X})$, and $f(\bm{Y})$ instead of $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$.
+    $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})$ (6)
+For notational simplicity, we have written ${g}(\boldsymbol{X})$ instead of ${g}(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X})$, and $f(\boldsymbol{Y})$ instead of $f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y})$.
 A key aspect of this procedure is the relationship between $N$ and $n$.
@@
 The PEMC evaluation step is summarized in Algorithm 2.
-1:procedure Evaluation($\bm{\theta},n,N,g$)
-2:     Generate i.i.d. $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$ from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$
-3:     Generate i.i.d. $(\bm{\tilde{X}}_{j})_{j\in[N]}$ independently from its marginal in $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$
-4:     Compute $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$
+1:procedure Evaluation($\boldsymbol{\theta},n,N,g$)
+2:     Generate i.i.d. $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$
+3:     Generate i.i.d. $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ independently from its marginal in $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$
+4:     Compute $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})$
 5:     return $PEMC$
@@
 Algorithm 2 Evaluation
-2.5 PEMC: An Example on Asian Options
-To demonstrate the PEMC framework, we return to our ongoing example on arithmetic Asian call option under the Heston model.
-It is worth noting that while the geometric Asian option is a natural control variate in this context, its closed-form expression is not available under the Heston model.
-To use control variates here, one must resort to sophisticated methods, e.g, numerically approximating the characteristic function Fusai and Meucci (2008) or employing new framewrok for conditional Monte Carlo Dingec et al.
-(2015).
-This section serves as a user manual, guiding readers through the process of implementing PEMC for pricing an exotic option.
-When pricing an arithmetic Asian call option under the Heston model, the option payoff is given by:
+3.5. PEMC: An Example on Asian Options
+We now return to our example on arithmetic Asian option under the Heston model:
 $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
-where $\bm{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
-The Heston model is described in (1) which requires $\bm{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$.
+where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
+The Heston model is described in (1) which requires $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$.
 For this example, let’s say we use a neural network (NN) as the prediction model.
 Then, the PEMC framework could proceed as follows:
-  1. 1.
-    Define the parameter space $\bm{\Theta}$ that encompasses realizations of
-    $\bm{\theta}:=(\underbrace{r,\eta,\delta,\rho,\kappa}_{\bm{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0},\Delta t,T}_{\bm{\theta}_{\text{simulation}}},\underbrace{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}}_{\bm{\theta}_{\text{payoff}}})$
-    which are of practical interest. The choice of $\bm{\Theta}$ also depends on the model capacity and the model updating frequency. We shall discuss this in subsequent sections.
-  2. 2.
-    Uniformly sample $\bm{\theta}$ from $\bm{\Theta}$.
+  1. (1)
+    Define the parameter space $\boldsymbol{\Theta}$ that encompasses realizations of
+    $\boldsymbol{\theta}:=(\underbrace{r,\eta,\delta,\rho,\kappa}_{\boldsymbol{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0},\Delta t,T}_{\boldsymbol{\theta}_{\text{simulation}}},\underbrace{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}}_{\boldsymbol{\theta}_{\text{payoff}}})$
+    which are of practical interest.
+  2. (2)
+    Uniformly sample $\boldsymbol{\theta}$ from $\boldsymbol{\Theta}$.
 This step is straightforward if the parameter space is a Cartesian product of intervals.
-For each sampled $\bm{\theta}$, generate process $\bm{Y}:=(S_{t},\nu_{t})_{t}$ using a Euler scheme with step size $\Delta t$, according to the Heston model specified by $\bm{\theta}$.
-  3. 3.
-    During the sampling of $\bm{Y}$, we collect $\bm{X}$ to be the sum of Brownian increment
+For each sampled $\boldsymbol{\theta}$, generate process $\boldsymbol{Y}:=(S_{t},\nu_{t})_{t}$ using a discretization scheme with step size $\Delta t$, with the Heston model specified by $\boldsymbol{\theta}$.
+  3. (3)
+    During the sampling of $\boldsymbol{Y}$, we collect $\boldsymbol{X}$ to be the sum of Brownian increment
     $W_{T}^{S}:=\sum_{j=1}^{T/\Delta t}\Delta W^{S}_{j\Delta t}\quad\text{ and }\quad W_{T}^{\nu}:=\sum_{j=1}^{T/\Delta t}\Delta W^{\nu}_{j\Delta t},$
     simulated during each step of the Euler scheme.
-This makes the marginal of $\bm{X}$ simply a two-dimensional Gaussian: $\frac{\bm{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
-  4. 4.
+This makes the marginal of $\boldsymbol{X}$ simply a two-dimensional Gaussian: $\frac{\boldsymbol{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
+  4. (4)
     Save
@@
 […]
-  5. 5.
+  5. (5)
     Repeat steps 2-4 $N_{\text{train}}$ times to generate dataset $(\textit{feature}_{i},\textit{label}_{i})_{i\in[N_{\text{train}}]}$ of size $N_{\text{train}}$.
-  6. 6.
-    Train a NN with weights $\bm{w}$ to minimize the MSE loss:
-    $\min_{{\bm{w}}}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-NN_{\bm{w}}(\text{feature}_{i}))^{2}$
+  6. (6)
+    Train a NN with weights $\boldsymbol{w}$ to minimize the MSE loss:
+    $\min_{{\boldsymbol{w}}}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-NN_{\boldsymbol{w}}(\text{feature}_{i}))^{2}$
     and use $\hat{w}$ to approximate
-    $NN_{\bm{\hat{w}}}\approx\mathbb{E}\bigg{[}(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K)^{+}\bigg{|}{r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu}\bigg{]}.$
-  7. 7.
-    At evaluation, a specific $\bm{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4).
+    $NN_{\boldsymbol{\hat{w}}}\approx\mathbb{E}\bigg{[}(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K)^{+}\bigg{|}{r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu}\bigg{]}.$
+  7. (7)
+    At evaluation, a specific $\boldsymbol{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4).
 We also generate $N$ i.i.d.
 $\sqrt{T}\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$ samples and store them as $(\tilde{\textit{feature}})_{j\in[N]}$.
-  8. 8.
-    Output $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(\textit{label}_{i}-NN_{\bm{\hat{w}}}(\textit{feature}_{i}))+\frac{1}{N}\sum_{j=1}^{N}NN_{\bm{\hat{w}}}(\tilde{\textit{feature}}_{j})$
-2.6 Implementation Details in PEMC
+  8. (8)
+    Output $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(\textit{label}_{i}-NN_{\boldsymbol{\hat{w}}}(\textit{feature}_{i}))+\frac{1}{N}\sum_{j=1}^{N}NN_{\boldsymbol{\hat{w}}}(\tilde{\textit{feature}}_{j})$
+3.6. Implementation Details in PEMC
 Having walked through the Asian option example under the PEMC framework, we discuss several important implementation details.
 These considerations are critical for successfully applying PEMC in practice.
-2.6.1 Parameter Space $\Theta$
-When determining the parameter space $\bm{\Theta}$, a key consideration is how frequently the model will be updated.
+3.6.1. Parameter Space $\Theta$
+When determining the parameter space $\boldsymbol{\Theta}$, a key consideration is how frequently the model will be updated.
 To ensure the performance of PEMC, it is desirable that the training data encompass all practical scenarios the model is expected to encounter within the designated update period.
-For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$.
-If one updates the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses about the bounds within which these parameters are likely to fluctuate over the coming month.
-For example, one could postulate the underlying price would mostly be fluctuating within a ±$\textpm 5-6\%$ range, i.e., around 490-550, based on the historical volatility of SPY.
-This expectation specifies an interval for $S_{0}$ in $\bm{\Theta}$.
-Similarly, one might set intervals for each parameter in the same manner, e.g., interval for $r\in(4.0\%,5.0\%)$, anticipating that the interest rates change would not exceed half a percentage point.
-Finally, we note that while these parameters are jointly distributed and could be calibrated together—especially considering that many market conditions are intertwined and correlated—this is often unnecessary.
-Instead, one can define intervals for each parameter individually and enlarge the uncertainty set, since the coverage of practical scenarios is more important in PEMC.
-This process is conceptually similar to the selection of uncertainty sets in robust optimization Ben-Tal et al.
-(2008) or distributionally robust optimization Esfahani and Kuhn (2018), where the goal is to include realizations of parameters with high likelihood.
+For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$ and the user of PEMC chooses to update the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses or confidence intervals about the bounds within which these parameters are likely to fluctuate over the coming month.
+This process is conceptually similar to the selection of uncertainty sets in robust optimization [16] or distributionally robust optimization [50], where the goal is to include realizations of parameters with high likelihood.
 However, unlike robust optimization approaches, which often favor data-dependent uncertainty sets, our focus here is also guided by the expertise of financial engineers.
 The training parameter space is designed to reflect realistic and practical scenarios derived from domain knowledge, rather than being strictly driven by statistical guarantees.
-In applications, we construct reasonable $\bm{\Theta}$ for practical models, such as forward curves in HJM modeling Glasserman (2013b) or local stochastic volatility 2D grids Gatheral (2006), among others.
-2.6.2 The Design of $\Large{X}$: Feature Engineering
-In our Asian option example, we designed the feature $\bm{X}$ as the sum of two correlated Brownian increments. As discussed in Section 2.3, the choice was motivated by two key properties:
-  • Predictability: The sum of Brownian motion increments explains away a large portion of the variance driving the movement of SDE.
-  • Ease of Generation: The sum of Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
-Given these reasons, it is natural to use such sums as $\bm{X}$ for all related SDE driving by Brownian motion.
-Indeed, this is the approach we adopt in our applications, and it has proven to be both robust and effective.
-For cases where the path is very long, one can divide the sum into chunks of smaller increments.
-However, we found that using just one increment typically gives sufficient variance reduction.
+In applications, we construct reasonable $\boldsymbol{\Theta}$ for practical models, such as forward curves in HJM modeling [60] or local stochastic volatility 2D grids [56], among others.
+3.6.2. The Design of $\Large{X}$: Feature Engineering
+In the example, we designed the feature $\boldsymbol{X}$ as the sum of two correlated Brownian increments. As discussed in Section 3.3, the choice was motivated by two key properties:
+  • Predictability: The sum of Brownian motion increments explains away a portion of the variance of the SDE.
+  • Marginal simulation: The sum of two Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
+Given these considerations, it is tempting to employ Gaussian sums as a generic recipe for the majority of option pricing problems handled by PEMC, particularly since these problems commonly involve stochastic differential equations driven by Brownian motion.
+Indeed, this is precisely the approach we adopt in our applications, demonstrating its robustness and effectiveness.
+In cases involving long simulation paths, one might consider segmenting the sum into smaller increments, though we observed that a single increment usually suffices to achieve significant variance reduction.
 However, the choice of Brownian increments, while effective, is certainly not the only one.
@@
 where $S_{t}$ is the asset price process and $T$ is the option maturity. A good choice of feature could be:
-$\bm{X}=(W_{T}^{S},\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}).$
+$\boldsymbol{X}=(W_{T}^{S},\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}).$
 The rationale behind the construction of $\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}$ term is grounded in the domain knowledge of the Heston model.
@@
 $S_{t}\approx S_{0}e^{(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}\text{ and }\min_{0\leq t\leq T}S_{t}\approx S_{0}e^{\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}.$
-This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\bm{X}]_{1}}-S_{0}e^{[\bm{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$.
-This corresponds directly to Property 1 in the construction of $\bm{X}$ in 2.3.
-For Property 2, the marginal distribution of $\bm{X}$ permits efficient and parallelizable simulation.
-In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\bm{X}]_{1},[\bm{X}]_{2})$ Karatzas and Shreve (1991).
+This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\boldsymbol{X}]_{1}}-S_{0}e^{[\boldsymbol{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$.
+This corresponds directly to Property 1 in the construction of $\boldsymbol{X}$ in 3.3.
+For Property 2, the marginal distribution of $\boldsymbol{X}$ permits efficient and parallelizable simulation.
+In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\boldsymbol{X}]_{1},[\boldsymbol{X}]_{2})$ [78].
 This result makes it straightforward to sample these components directly.
-2.6.3 Eliminating Data Storage Overheads
+3.6.3. Eliminating Data Storage Overheads
 In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively.
@@
 For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training.
-In this manner, we have successfully trained models in the applications using data on the order of $10^{7}–10^{8}$.
-3 Analysis and Performance
-In this section, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical considerations.
-All properties of PEMC discussed here pertain to the evaluation stage, where $\bm{\theta}$ is fixed.
-Henceforth, we omit the explicit dependence on $\bm{\theta}$ in the notation (e.g., writing $\bm{Y}$ in place of $\bm{Y}(\bm{\theta})$).
+In this manner, we have successfully trained models in the applications using data on the order of $10^{7}\textendash 10^{8}$.
+4. Analysis and Performance
+In this section, we provide a detailed and comprehensive analysis of the PEMC estimator.
+All properties of PEMC discussed here pertain to the evaluation stage, where $\boldsymbol{\theta}$ is arbitrary but fixed.
+Henceforth, we omit the explicit dependence on $\boldsymbol{\theta}$ in the notation (e.g., writing $\boldsymbol{Y}$ in place of $\boldsymbol{Y}(\boldsymbol{\theta})$).
 We begin by establishing the unbiasedness of our estimator.
-3.1 Bias Analysis
-Theorem 3.1 (Unbiasedness)
+4.1. Bias Analysis
+Theorem 1 (Unbiasedness).
 The PEMC estimator in equation (6), is unbiased:
-$\mathbb{E}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})\bigg{]}=\mathbb{E}[f(\bm{Y})]$
-Proof 3.2
-[…]
+$\mathbb{E}\bigg{[}\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})\bigg{]}=\mathbb{E}[f(\boldsymbol{Y})]$
+Proof.
+Proof First note that $g$ is pre-trained. The proof then follows since $\boldsymbol{X}_{i}$ and $\boldsymbol{\tilde{X}}_{j}$ have the same marginal distribution for any $i,j$. ∎
 Given the unbiased nature of PEMC, the analysis of PEMC as a Monte Carlo method primarily hinges on its variances and computational costs.
-3.2 Variance Analysis
+4.2. Variance Analysis
 To facilitate the analysis, we first introduce some necessary notation.
-Definition 3.3
-Given function $f$ that takes input $\bm{Y}$ and $g$ that takes input $\bm{X}$ (so both $f,g$ are considered fixed here), we denote
-$\sigma_{f-g}^{2}:=\mathrm{Var}(f(\bm{Y})-g(\bm{X})),\text{ }\sigma^{2}_{f}:=\mathrm{Var}(f(\bm{Y})),\text{ and }\sigma^{2}_{g}:=\mathrm{Var}(g(\bm{X})).$
-Similarly, we denote cost of generating a coupled sample $f(\bm{Y})-g(\bm{X})$ as $c_{f-g}$, the cost of generating $f(\bm{Y})$ as $c_{f}$ and the cost of generating $g(\bm{X})$ as $c_{g}$.
-Lemma 3.4
-Using the notations from 3.3, we have
-$\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))=\frac{\sigma^{2}_{f}+\sigma^{2}_{g}-\sigma^{2}_{f-g}}{2}.$
+Definition 2.
+Given function $f$ that takes input $\boldsymbol{Y}$ and $g$ that takes input $\boldsymbol{X}$ (so both $f,g$ are considered fixed here), we denote
+$\sigma_{f-g}^{2}:=\mathrm{Var}(f(\boldsymbol{Y})-g(\boldsymbol{X})),\text{ }\sigma^{2}_{f}:=\mathrm{Var}(f(\boldsymbol{Y})),\text{ and }\sigma^{2}_{g}:=\mathrm{Var}(g(\boldsymbol{X})).$
+Similarly, we denote cost of generating a coupled sample $f(\boldsymbol{Y})-g(\boldsymbol{X})$ as $c_{f-g}$, the cost of generating $f(\boldsymbol{Y})$ as $c_{f}$ and the cost of generating $g(\boldsymbol{X})$ as $c_{g}$.
+Lemma 1.
+Using the notations from 2, we have
+$\mathrm{Cov}(f(\boldsymbol{Y}),g(\boldsymbol{X}))=\frac{\sigma^{2}_{f}+\sigma^{2}_{g}-\sigma^{2}_{f-g}}{2}.$
 We make two remarks here.
-First, since we generally obtain $\bm{X}$ in the same process during the generation of $\bm{Y}$, thus typically in either sense of the cost, we have $c_{f-g}\approx c_{f}$.
-Second, the term cost is intentionally left ambiguous because its interpretation varies depending on the context.
-In some cases, cost refers to the sample size, while in others it represents the wall-clock time required to generate and evaluate the sample.
-Furthermore, in certain scenarios, both time and computational resources must be jointly considered and optimized, under parallelization and other speedup, making the concept of cost more instance dependent.
+First, since we generally obtain $\boldsymbol{X}$ in the same process during the generation of $\boldsymbol{Y}$, thus typically in either sense of the cost, we have $c_{f-g}\approx c_{f}$.
+Second, the term cost is intentionally left unspecified because its interpretation varies depending on the context.
+In some context, cost refers to the sample size, while in others it represents the wall-clock time required to generate and evaluate all the samples, making the concept of cost instance dependent.
 Therefore, for now, we use cost as an umbrella term to capture various meanings under different contexts.
-Lemma 3.5
+Lemma 2.
 The Variance of the PEMC estimator in equation (6) is
 $\mathrm{Var}(PEMC)=\frac{1}{n}\sigma_{f-{g}}^{2}+\frac{1}{N}\sigma_{g}^{2}.$
-Proof 3.6
-[…]
+Proof.
+Proof First note that $g$ is pre-trained.
+Moreover, in the evaluation phase, data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$ are generated independent of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$.
+∎
 Having established the unbiasedness and the variance form of the PEMC estimator, we could turn to inference.
@@
 Moreover, a range of inferential techniques may be employed—such as deriving high-probability bounds or other non-asymptotic guarantees—the asymptotic confidence interval construction suffices to illustrate the point.
-The approach follows closely from results in prediction-powered inference Angelopoulos et al.
-(2023).
-Theorem 3.7 (Asymptotic Confidence Intervals)
-In the set up of the evaluation algorithm 2, given the prediction model $g$, $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$ and $(\bm{\tilde{X}}_{j})_{j\in[N]}$, we define
-$\hat{\sigma}^{2}_{f-g}=\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}f(\bm{Y}_{i})-g(\bm{X}_{i})-\Bigl{(}\frac{1}{n}\sum_{i^{\prime}=1}^{n}f(\bm{Y}_{i^{\prime}})-g(\bm{X}_{i^{\prime}})\Bigr{)}\Bigr{)}^{2},$
+The approach follows closely from results in prediction-powered inference [6].
+Theorem 3 (Asymptotic Confidence Intervals).
+In the set up of the evaluation algorithm 2, given the prediction model $g$, $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$ and $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$, we define
+$\hat{\sigma}^{2}_{f-g}=\frac{1}{n}\sum_{i=1}^{n}\Bigl{(}f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i})-\Bigl{(}\frac{1}{n}\sum_{i^{\prime}=1}^{n}f(\boldsymbol{Y}_{i^{\prime}})-g(\boldsymbol{X}_{i^{\prime}})\Bigr{)}\Bigr{)}^{2},$
 and
-$\hat{\sigma}^{2}_{g}=\frac{1}{N}\sum_{j=1}^{N}\bigl{(}g(\bm{\tilde{X}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}g(\bm{\tilde{X}}_{j^{\prime}})\bigr{)}^{2},$
+$\hat{\sigma}^{2}_{g}=\frac{1}{N}\sum_{j=1}^{N}\bigl{(}g(\boldsymbol{\tilde{X}}_{j})-\frac{1}{N}\sum_{j^{\prime}=1}^{N}g(\boldsymbol{\tilde{X}}_{j^{\prime}})\bigr{)}^{2},$
 be the respective sample variance. Let $z_{1-\alpha/2}$ denote the $(1-\alpha/2)$-quantile of the standard normal distribution. Then, the interval
 $\Bigg{(}PEMC-z_{1-\alpha/2}\sqrt{\frac{\hat{\sigma}^{2}_{f-g}}{n}+\frac{\hat{\sigma}^{2}_{g}}{N}},\quad PEMC+z_{1-\alpha/2}\sqrt{\frac{\hat{\sigma}^{2}_{f-g}}{n}+\frac{\hat{\sigma}^{2}_{g}}{N}}\text{ }\Bigg{)},$
-is an asymptotically valid $1-\alpha$ confidence interval for $\mathbb{E}[f(\bm{Y})]$, as $n,N\to\infty$.
-Proof 3.8
-[…]
-3.3 Variance Reduction
+is an asymptotically valid $1-\alpha$ confidence interval for $\mathbb{E}[f(\boldsymbol{Y})]$, as $n,N\to\infty$.
+Proof.
+Proof Theorem 1 establishes PEMC is unbiased.
+Moreover, $\hat{\sigma}^{2}_{f-g},\hat{\sigma}^{2}_{g}$ are consistent estimates of ${\sigma}^{2}_{f-g},{\sigma}^{2}_{g}$, the results follows from Lemma 2 and the central limit theorem and Slutsky’s theorem.
+∎
+4.3. Variance Reduction
 Having analyzed the bias and variance of the PEMC estimator, we now turn to a fundamental practical question: Under a fixed cost budget—whether measured in terms of the number of samples, computational time, or other limited resources—when does PEMC outperform standard Monte Carlo (MC)?
@@
 This investigation provides a guideline for parameter selection and reveals conditions under which PEMC’s gains are maximized.
-Lemma 3.9
+Lemma 3.
 Assuming $n,N\in\mathbb{R}^{+}$ and $c_{g},c_{f-g}\in\mathbb{R}^{+}$, the optimal allocation between $n$ and $N$ for PEMC, for any positive budget $B$, follows as
 $\frac{n}{N}=\frac{\sigma_{f-g}}{\sigma_{g}}\cdot\sqrt{\frac{c_{g}}{c_{f-g}}}\quad.$ (8)
-Proof 3.10
+Proof.
 Proof By relaxing the constraints to $n,N\in\mathbb{R}^{+}$, the optimization problem
@@
 s.t. $nc_{f-g}+Nc_{g}\leq B.$
-[…]
-In practice, $c_{f-g}\gg c_{g}$ often holds which suggests one should set $N\gg n$ according to Lemma 3.9.
-While Lemma 3.9 relies on treating $n,N\in\mathbb{R}^{+}$ as continuous—ignoring the integer constraints $n,N\in\mathbb{N}^{+}$— this continuous approximation still provides valuable guidance when $n,N\geq 1$.
-In fact, in practice, we found the ratio suggested by Lemma 3.9 guides near-optimal parameter selection.
-However, the standard MC corresponds to the case $N=0$, a scenario not covered by Lemma 3.9, and thus requires separate consideration.
-Building on Lemma 3.9, we can estimate an upper bound on the variance reduction PEMC can achieve relative to standard MC under these idealized assumptions.
-Lemma 3.11
-Assume $c_{f-g}=c_{f}$. In the same setup as Lemma 3.9, the variance ratio between PEMC under the optimal allocation and standard MC follows as
+is jointly convex in $n,N\in\mathbb{R}^{+}$.
+The objective value also approaches infinity as $N\to 0$ or $n\to 0$.
+Consequently, the strict convexity ensures the existence of a global minimizer in the interior and directly solving the Lagrangian gives us the result.
+∎
+In practice, $c_{f-g}\gg c_{g}$ often holds which suggests one should set $N\gg n$ according to Lemma 3.
+While Lemma 3 relies on treating $n,N\in\mathbb{R}^{+}$ as continuous—ignoring the integer constraints $n,N\in\mathbb{N}^{+}$— this continuous approximation still provides valuable guidance when $n,N\geq 1$.
+In fact, in practice, we found the ratio suggested by Lemma 3 guides near-optimal parameter selection.
+However, the standard MC corresponds to the case $N=0$, a scenario not covered by Lemma 3, and thus requires separate consideration.
+Building on Lemma 3, we can estimate an upper bound on the variance reduction PEMC can achieve relative to standard MC under these idealized assumptions.
+Lemma 4.
+Assume $c_{f-g}=c_{f}$. In the same setup as Lemma 3, the variance ratio between PEMC under the optimal allocation and standard MC follows as
 $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}=\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}\bigg{(}1+\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{g}}{c_{f}}}\bigg{)}+\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}\bigg{(}\frac{\sigma_{f-g}}{\sigma_{g}}\cdot\sqrt{\frac{c_{g}}{c_{f}}}+\frac{c_{g}}{c_{f}}\bigg{)}.$ (9)
-Proof 3.12
-[…]
-At this point, a natural question arises: how can we gauge $\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}$ or $\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}$ in practice?
+Proof.
+Proof Based on the results of Lemma 3, we can deduce that this ratio again does not depend on the budget $B$, so we omit its dependence in the statement.
+Moreover, we can conveniently choose a budget of the form $B=n_{0}c_{f-g}+n_{0}\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{f-g}}{c_{g}}}c_{g}$ which gives an allocation of $n=n_{0}$ and $N=n_{0}\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{f-g}}{c_{g}}}$ for PEMC in accordance with Lemma 3, while comparing with $B/c_{f}$ samples for standard MC.
+The rest follows from Lemma 2 and the assumption $c_{f}=c_{f-g}$.
+∎
+Based on the result of Lemma 4, a natural question arises: how can we gauge $\frac{\sigma_{f-g}^{2}}{\sigma_{f}^{2}}$ or $\frac{\sigma_{g}^{2}}{\sigma_{f}^{2}}$ in practice?
 How are these ratios tied to the quality of the predictive model $g$ in PEMC?
@@
 To illustrate this, we first consider an ideal scenario where $g$ is trained to the optimum, i.e., the true regression function.
-Lemma 3.13
-Suppose $c_{f-g}=c_{f}$, $f(\bm{Y})$ is square-integrable and $g=\mathbb{E}[f(\bm{Y})\,|\,\bm{X}]$.
+Lemma 5.
+Suppose $c_{f-g}=c_{f}$, $f(\boldsymbol{Y})$ is square-integrable and $g=\mathbb{E}[f(\boldsymbol{Y})\,|\,\boldsymbol{X}]$.
 Define $\rho:=\mathrm{corr}(f,g)$ and $c:=\frac{c_{g}}{c_{f}}$.
-Then we have $\rho=\frac{\sigma_{g}}{\sigma_{f}}$ and the $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}$ in Lemma 3.11 becomes
+Then we have $\rho=\frac{\sigma_{g}}{\sigma_{f}}$ and the ratio $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}$ in Lemma 4 reduces to
 $r(\rho,c):=$ $(1-\rho^{2})\bigg{(}1+\frac{\rho}{\sqrt{1-\rho^{2}}}\cdot\sqrt{c}\bigg{)}+\rho^{2}\bigg{(}\frac{\sqrt{1-\rho^{2}}}{\rho}\cdot\sqrt{c}+c\bigg{)}.$ (10)
-Proof 3.14
-Proof When $g=\mathbb{E}[f(\bm{Y})\,|\,\bm{X}]$ is the true regressor, i.e.,
-[…]
-[…]
-Lemma 3.13 says that the variance reduction of PEMC relative to standard MC can be approximated by the function $r(\rho,c)$ in (8), where $\rho=\sigma_{g}/\sigma_{f}$ captures the predictive power of $g(\bm{X})$ for $f(\bm{Y})$, and $c=c_{g}/c_{f}$ quantifies the relative cost of evaluating $g(\bm{X})$ versus generating full samples for $f(\bm{Y})$.
-In Figure 3.3, we visualize the variance reduction function $r(\rho,c)$ for $\rho,c\in(0,1)$.
-On the left is a contour plot of $r(\rho,c)$ where the red line traces the curve $r(\rho,c)=1$, corresponding to the “break-even” boundary where PEMC’s variance matches that of MC.
+Proof.
+Proof When $g=\mathbb{E}[f(\boldsymbol{Y})\,|\,\boldsymbol{X}]$ is the true regressor, i.e.,
+$g=\operatorname*{arg\,min}_{h\text{ measurable}}\mathbb{E}[(f(\boldsymbol{Y})-h(\boldsymbol{X}))^{2}],$
+and $f$ is square-integrable, it follows that $\mathbb{E}[(f-g)h]=0$ for all square-integrable $h$.
+Plug in $h=g$, we obtain $\mathrm{Cov}(f-g,g)=0$.
+Consequently we have $\rho:=\mathrm{corr}(f,g)=\frac{\sigma_{g}}{\sigma_{f}}$ and $\sigma_{f}^{2}=\sigma_{f-g}^{2}+\sigma^{2}_{g}$, which further implies $\sigma_{g}=\rho^{2}\sigma_{f}^{2}$ and $\sigma_{f-g}^{2}=(1-\rho^{2})\sigma^{2}_{f}$.
+The rest follows from $c_{f-g}=c_{f}$.
+∎
+Lemma 5 says that the variance reduction of PEMC relative to standard MC can be approximated by the function $r(\rho,c)$ in (10), where $\rho=\sigma_{g}/\sigma_{f}$ captures the predictive power of $g(\boldsymbol{X})$ for $f(\boldsymbol{Y})$, and $c=c_{g}/c_{f}$ quantifies the relative cost of evaluating $g(\boldsymbol{X})$ versus generating full samples for $f(\boldsymbol{Y})$.
+In Figure 1, we visualize the variance reduction function $r(\rho,c)$ for $\rho,c\in(0,1)$.
+On the left is a contour plot of $r(\rho,c)$ where the red line traces the curve $r(\rho,c)=1$, corresponding to the “break-even" boundary where PEMC’s variance matches that of MC.
 Regions below this line, $\{(\rho,r)\in[0,1]^{2}|r(\rho,c)<1\}$, represent regimes of $\rho$ and $c$ where PEMC achieves variance reduction, i.e., where $\rho$ is sufficiently large and $c$ is sufficiently small.
 On the right, the graph of $r(\rho,c)$ as a function of $\rho$ when $c=0.001$ provides a clear benchmark: a correlation $\rho=0.5$ yields approximately 22.2% variance reduction, while $\rho=0.7$ yields about 45.8%.
-In our PEMC applications, we design feature $\bm{X}$ and predictive model $g$ so that, for most $\bm{\theta}\in\Theta$ during evaluation, we can safely gauge that $c$ falls between $10^{-2}$ to $10^{-3}$ and $\rho$ exceeds 0.5.
-In these regions, Lemma 3.9 suggests choosing $N/n$ in the range of 5 to 20 for near-optimal performance.
+In our PEMC applications, we design feature $\boldsymbol{X}$ and predictive model $g$ so that, for most $\boldsymbol{\theta}\in\Theta$ during evaluation, we can safely gauge that $c$ falls between $10^{-2}$ to $10^{-3}$ and $\rho$ exceeds 0.5.
+In these regions, Lemma 3 suggests choosing $N/n$ in the range of 5 to 20 for near-optimal performance.
 In our experiments, a ratio of $N=10n$ proved effective.
-[…]
+Figure 1.
 Variance Reduction Function $r(\rho,c)$.
-Left figure: A contour map of the variance reduction ratio $r(\rho,c)$ as a function of the correlation $\rho=\sigma_{g}/\sigma_{f}$ and relative cost $c=c_{g}/c_{f}$ The red curve indicates the curve $r(\rho,c)=1$ and the regimes where PEMC attains variance reduction over MC lie below the line.
-Right figure: A graph of $r(\rho,c)$ as a function of $\rho$ when $c=10^{-3}$.
-3.3.1 Learning Theory Estimates
-The statements in Lemma 3.9 assume that $g$ represents the true regression function $\mathbb{E}[f|\bm{X}]$.
+Left: A contour map of the variance reduction ratio $r(\rho,c)$ as a function of the correlation $\rho=\sigma_{g}/\sigma_{f}$ and relative cost $c=c_{g}/c_{f}$.
+The red curve indicates the level set $r(\rho,c)=1$; the regimes where PEMC outperforms standard MC lie below the curve.
+Right: A graph of $r(\rho,c)$ as a function of $\rho$ when $c=10^{-3}$.
+4.3.1. Learning Theory Estimates
+The statements in Lemma 3 assume that $g$ represents the true regression function $\mathbb{E}[f|\boldsymbol{X}]$.
 In practice, however, $g$ is obtained through a learning procedure applied to finite training samples, and thus will only approximate the true conditional expectation.
-Consequently, any practical implementation of PEMC must account for the approximation and estimation errors inherent in the training process.
-Moreover, as the parameter $\bm{\theta}$ varies, the correlation $\rho$ between $g(\bm{X})$ and $f(\bm{Y})$, the complexity of the underlying model and payoff function also changes.
-To rigorously quantify these effects, one can invoke tools from statistical learning theory, such as uniform convergence bounds, VC-dimension, or Rademacher complexities (Vapnik 1998, Bartlett and Mendelson 2005, Mohri et al.
-2018).
+Consequently, any practical implementation of PEMC must account for the approximation and estimation errors inherent in the ML training process, and this error also varies with $\boldsymbol{\theta}$.
+To rigorously quantify these effects, one can invoke tools from statistical learning theory, such as uniform convergence bounds, VC-dimension, or Rademacher complexities [126, 13, 94].
 These quantities, by relating the complexity of the function class used to represent $g$ and the available training sample size $N_{\text{train}}$, lead to error bounds that ensure, with high probability, that the trained predictor $g$ is close to the true regression function within a controlled margin.
-Although numerous, more refined approaches exist, we avoid delving deeply into them so as not to distract from our main theme.
-Instead, we adopt a standard, off-the-shelf, almost ‘placeholder’-type result detailed in the Appendix, that illustrates the typical form of guarantees one could expect from PEMC.
-Lemma 3.15
+We avoid delving deeply into them so as not to distract from our main theme.
+Instead, we adopt a standard, off-the-shelf and ‘placeholder’-type result, whose proof is detailed in the Appendix, that illustrates the typical form of guarantees one could expect from PEMC.
+Lemma 6.
 Under the regularity conditions specified in the Appendix, for any $\epsilon,\delta>0$, there exists a sufficiently large sample size $N_{\text{train}}$ and a suitably chosen neural network class from which we select and train a predictor ${g}$ on $N_{\text{train}}$ samples, such that
@@
 for a randomly evaluated $\theta\sim\Theta$ with probability at least $1-\delta$.
-The simplified version captures the essential idea: under suitable regularity conditions, one can achieve the result of Lemma 3.9 within an arbitrarily small margin, as both the sample size $N_{\text{train}}$ and the complexity of the hypothesis class increase.
-3.4 Control Variate Coefficient
+4.4. Control Variate Coefficient
 In this section we explore a connection between the PEMC framework and traditional Control Variate (CV) methods. In (6), note that we could introduce a free parameter $a$ to create a variant:
-$PEMC(a):=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-ag(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}ag(\bm{\tilde{X}}_{j})$ (9)
-The introduction of $a$ is analogous to CV approaches, where the parameter $a$ is chosen to minimize variance.
+$PEMC(a):=\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-ag(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}ag(\boldsymbol{\tilde{X}}_{j})$ (11)
+The introduction of $a$ is analogous to CV approaches, where the parameter $a$ does not introduce bias but can be tuned to minimize variance.
 In PEMC, this viewpoint may appear somewhat redundant since $ag$ is also a valid prediction model, which suggests $a$ would be chosen implicitly during the training of the prediction model.
 Nevertheless, suppose $g$ is fixed.
-Then $\frac{1}{n}\sum_{i=1}^{n}-g(\bm{X}_{i})+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$ can be viewed as a readily available, zero-mean control variate.
+Then $\frac{1}{n}\sum_{i=1}^{n}-g(\boldsymbol{X}_{i})+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})$ can be viewed as a readily available, zero-mean control variate for $\frac{1}{n}\sum_{i=1}^{n}f(\boldsymbol{Y}_{i})$.
 We can then choose $a$ which minimizes the variance.
 The optimal $a^{*}$ then is:
-$a^{*}=\frac{\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))}{(n/N+1)\mathrm{Var}(g(\bm{X}))}.$
-In the ideal scenario, we have $N\gg n$ and $g(\bm{X})=\mathbb{E}[f(\bm{Y})|\bm{X}]$, which would imply $\mathrm{Cov}(f(\bm{Y})-g(\bm{X}),g(\bm{X}))=0$ and $\mathrm{Cov}(f(\bm{Y}),g(\bm{X}))=\mathrm{Var}(g(\bm{X}))$.
-This would further imply $a^{*}=\frac{1}{1+n/N}\approx 1$.
-Consequently, we can justify PEMC’s default choice of $a=1$, which saves us the trouble of estimating $a^{*}$, and gives PEMC a degree of robustness.
-3.5 Variance Reduction of PEMC: An Example on Asian Options
+$a^{*}=\frac{\mathrm{Cov}(f(\boldsymbol{Y}),g(\boldsymbol{X}))}{(n/N+1)\mathrm{Var}(g(\boldsymbol{X}))}.$
+In the ideal scenario, we have $N\gg n$ and $g(\boldsymbol{X})=\mathbb{E}[f(\boldsymbol{Y})|\boldsymbol{X}]$, which would imply $\mathrm{Cov}(f(\boldsymbol{Y})-g(\boldsymbol{X}),g(\boldsymbol{X}))=0$ and $\mathrm{Cov}(f(\boldsymbol{Y}),g(\boldsymbol{X}))=\mathrm{Var}(g(\boldsymbol{X}))$.
+This leads to $a^{*}=\frac{1}{1+n/N}\approx 1$, the default choice of PEMC, i.e., $a=1$.
+5. Applications
+After establishing the theoretical foundations of PEMC and demonstrating its effectiveness in controlled environments, we now apply it to real-world financial scenarios where the complexity of models often renders standard variance reduction techniques difficult or unavailable.
+In this section, we first examine PEMC in practice through two complex options pricing problems.
+In particular, we examine variance swaps pricing under stochastic local volatility models [46, 43, 56], and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework [34].
+In both cases, the path-dependent nature of the contracts and the complexity of the model dynamics make traditional control variates difficult.
+5.1. Variance Swaps in stochastic local vol models
+Variance swaps are financial derivatives that enable investors to trade future realized volatility against current implied volatility [14].
+Unlike traditional options, which derive their value from the underlying asset’s price, variance swaps are based on the variance of the asset’s returns over a specified period [42].
+This unique structure allows for pure exposure to volatility, making variance swaps valuable tools for risk management and speculative strategies [56, 108].
+In this subsection, we focus on using PEMC to price variance swaps under a stochastic local volatility (SLV) model [111], as closed-form formulas exist for variance swaps in stochastic volatility models such as Heston [134], but not for SLV [125] which provides greater flexibility through a data-intensive, non-parametric specification of the volatility surface [56].
+SLV models emerged as a hybrid approach that combines the market-implied local volatility surface with stochastic volatility dynamics [63], providing practitioners with greater calibration flexibility and more accurate price reproduction across strike-maturity ranges [111].
+Unlike the Heston model’s parsimonious parameter set, SLV requires handling a full volatility surface discretized on a dense 2D grid, effectively making the parameter space $\boldsymbol{\Theta}$ for PEMC high-dimensional.
+To efficiently process this grid-structured volatility data, we adopt a Convolutional Neural Network (CNN) architecture [83] for PEMC, effectively feeding the local vol surface as an image into predictive modeling.
+The CNNs are particularly well-suited for this task as they naturally exploit the spatial relationships in the volatility surface, similar to their success in image processing tasks [64, 81].
+Under the SLV model, we adopt the following SDEs for simulating the asset price $\{S_{t}\}_{t}$:
+$\mathrm{d}S_{t}$ $=rS_{t}\mathrm{d}t+\sigma(S_{t},t)e^{\nu_{t}}S_{t}\mathrm{d}W_{t}^{S},$ (12)
+$\mathrm{d}\nu_{t}$ $=\kappa(\eta_{t}-\nu_{t})\mathrm{d}t+\delta\mathrm{d}W_{t}^{\nu},$ (13)
+with $\langle\mathrm{d}W_{t}^{S},\mathrm{d}W_{t}^{\nu}\rangle=\rho dt$ and $\eta_{t}:=-\frac{\delta^{2}}{2\kappa}(1+e^{-2\kappa t}).$ Here $\sigma(\cdot,\cdot):\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is 2D function representing the (interpolated) local volatility surface and $\exp(\nu_{t})$ is a stochastic multiplier with $\exp(\nu_{0})=1$ and $\mathbb{E}[e^{2\nu_{t}}]=1$.
+In practice, local volatility surfaces are calibrated to and stored as discrete two-dimensional grids indexed by asset prices (spot) and time.
+During simulation, these discrete values are interpolated as needed to obtain volatilities at arbitrary price-time points [37].
+For our PEMC implementation, we treat this discrete grid as part of our input parameter $\boldsymbol{\theta}_{\text{model}}$, reflecting how the market-calibrated surfaces would be used in practice.
+While local volatility models [46, 43] and their calibration to market data constitute an extensive research area in their own right, our focus here is on the PEMC implementation.
+Thus, we assume a calibrated local volatility has been given for evaluation, regardless of the method used to obtain it.
+Following practical conventions, we store $\sigma(\cdot,\cdot)$ on a $|\mathcal{S}|\times|\mathcal{T}|$ grid, where $\mathcal{S}$ contains $|\mathcal{S}|$ equally-spaced price points in $[S_{\text{surface}}^{\min},S_{\text{surface}}^{\max}]$ and $\mathcal{T}$ contains $|\mathcal{T}|$ equally-spaced time points in $[t_{\text{surface}}^{\min},t_{\text{surface}}^{\max}]$.
+At each point on the grid, we store the value of local volatility according to [34]:
+$\sigma_{base}^{2}(x,t)=\frac{\sum_{i=0}^{2}p_{i}\tau_{i}e^{-x^{2}/(2t\tau_{i}^{2})-t\tau^{2}_{i}/8}}{\sum_{i=0}^{2}(p_{i}/\tau_{i})e^{-x^{2}/(2t\tau_{i}^{2})-t\tau_{i}^{2}/8}},\quad\text{with}~{}~{}~{}p_{0}:=1-p_{1}-p_{2},~{}~{}x:=\log(S_{t}/S_{0}),$ (14)
+for $p_{0}=0.3,p_{1}=0.5,p_{2}=0.2,\tau_{0}=0.4,\tau_{1}=0.3,\tau_{2}=0.6$ as in Figure 2 in [34].
+While this analytical form (14) is used in our data-generating process, it is important to note that PEMC treats the surface as any market-calibrated volatility surface - accessing it only through its discrete grid values.
+When sampling $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ to produce $\sigma$ grid, we add a $\mathcal{N}(0,\xi^{2})$ noise independently to all the $|\mathcal{S}|\times|\mathcal{T}|$ points in the grid, on top of their baseline value $\sigma_{base}$ (14).
+A path is then generated by Euler’s scheme:
+$\nu_{t+\Delta t}$ $\leftarrow\nu_{t}+\kappa(\eta_{t}-\nu_{t})\Delta t+\delta\Delta W^{\nu}_{t},$
+$S_{t+\Delta t}$ $\leftarrow S_{t}\exp\Big{(}\big{(}r-\tfrac{1}{2}\tilde{\nu}_{t+\Delta t}^{2}\big{)}\Delta t+\tilde{\nu}_{t+\Delta t}\Delta W^{S}_{t}\Big{)},\quad\text{where}~{}~{}\tilde{\nu}_{t+\Delta t}=\widehat{\sigma}(S_{t},t)e^{\nu_{t+\Delta t}},$
+with $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}(\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$, and $\widehat{\sigma}(S_{t},t)$ obtained by interpolations of the grid $\sigma$.
+The input $(\boldsymbol{\theta},\boldsymbol{X})$ can be represented as
+$\text{feature}_{i}=\{\underbrace{\{\sigma^{2}(s,t)\}_{s\in\mathcal{S},t\in\mathcal{T}}}_{\text{surface info }},\underbrace{r,\delta,\kappa,\rho,\mu}_{\boldsymbol{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0}}_{\boldsymbol{\theta}_{\text{simulation}}},\underbrace{K}_{\boldsymbol{\theta}_{\text{payoff}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}(\boldsymbol{\theta})}\}.$
+The parameters space $\boldsymbol{\Theta}$, as well as the evaluation $\boldsymbol{\theta}$, is summarized in Table 1.
+Table 1. Parameter Setup in SLV Model
+mode $\xi$ $S_{0}$ $r$ $\kappa$ $\lambda$ $\rho$ $T$, $\Delta t$
+$N_{\text{train}}=3,000,000$ $0.02$ $[50,150]$ $0.02$ $[1.5,4.5]$ $[0.1,1.0]$ $[-0.2,-0.9]$ $1$, $1/252$
+evaluation 00 $100$ $0.02$ $3.0$ $0.5$ $-0.5$
+To handle the 2D grid of high-dimensional volatility surface data, we design a two-branch neural network architecture, which is illustrated in Figure 2.
+The first branch processes the discretized volatility surface $\sigma^{2}(x,t)$ using a CNN architecture inspired by VGG [118], which has become a standard choice for image processing tasks and is readily available in modern deep learning packages like PyTorch [103].
+This branch then consists of two 2D convolutional layers interspersed with ReLU activations, followed by a MaxPool2d operation for dimensionality reduction.
+The surface features are then flattened through a fully connected layer to produce an embedding.
+The second branch handles the remaining model parameters in $(\boldsymbol{\theta},\boldsymbol{X})$ through a series of fully connected layers with dropout regularization, batch normalization, and ReLU activations, ultimately producing another embedding.
+Finally, the two separate embeddings are then fed into a “Synthesizer" module, which combines the information through additional fully connected layers with dropout and ReLU activations to produce the final prediction.
+This architecture choice is motivated by the proven effectiveness of CNNs in handling grid-structured data [83, 64], and particularly the VGG architecture’s success in extracting hierarchical features while maintaining relative simplicity [118].
+The details of NN architecture are summarized in Table 2.
+Figure 2. Neural Network architecture for the SLV Model.
+Table 2. Neural Network Architecture Parameters
+CNN Branch Feed-forward Branch
+kernel size: 3, stride: 1, padding: 1
+max pooling (kernel: 2, stride: 2, padding: 0)
+hidden dim: 512
+output dim: 128
+Finally, we evaluate PEMC’s performance in the SLV setting across sample sizes $n\in\{1000,2000,$ 4000,8000,10000,20000}4000,8000,10000,20000\}4000 , 8000 , 10000 , 20000 } with $N=10n$, benchmarking against a ground truth computed from $5\times 10^{7}$ MC samples.
+The results, based on 200 independent experiments and shown in Table 3, reveal that PEMC’s effectiveness persists even in this more complex setting.
+Despite the added complexity of handling high-dimensional volatility surfaces, PEMC achieves a 30-40% reduction in mean squared error compared to standard MC, demonstrating its robustness as a variance reduction technique across different model frameworks.
+Table 3. RMSE from 200 Experiments for Pricing Variance Swap under the SLV Model
+Method $n=1000$ $n=2000$ $n=4000$ $n=6000$ $n=8000$ $n=10000$ $n=20000$
+Monte Carlo (MC) 0.0206 0.0145 0.0101 0.0075 0.0064 0.0065 0.0047
+PEMC 0.0130 0.0088 0.0061 0.0055 0.0040 0.0040 0.0027
+5.2. Swaptions in HJM Models
+Interest rate derivatives play a crucial role in financial markets, with swaptions being particularly important instruments for managing interest rate risk [29].
+A swaption gives its holder the right to enter into an interest rate swap at a future date, providing flexibility in hedging future interest rate exposures [73].
+The pricing of these instruments, however, presents significant computational challenges due to the high-dimensional nature of interest rate modeling [5].
+In this subsection, we demonstrate PEMC’s application to swaption pricing under the Heath-Jarrow-Morton (HJM) framework [65].
+The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models.
+For illustration purposes, we focus on a one-factor specification with exponential volatility structure [60], though the framework readily extends to multi-factor cases.
+A swaption is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date.
+In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite.
+Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$.
+The value of this swap at time $t^{\prime}_{0}$ is:
+$V_{t^{\prime}_{0}}=C\left(R\sum_{l=1}^{n_{p}}B(t^{\prime}_{0},t^{\prime}_{l})\Delta t^{\prime}+B(t^{\prime}_{0},t^{\prime}_{n_{p}})-1\right),$
+where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}F$.
+A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$.
+The payoff of the swaption is simply
+$\max(0,V_{t^{\prime}_{0}}),$
+[…]
+$df(t,T)=\mu(t,T)\,dt+\sigma(t,T)^{\top}dW(t),$
+where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion.
+In contrast to short-rate models (e.g., Vasicek [127] or Cox-Ingersoll-Ross [39]), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates [65].
+The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures [29, 5].
+However, the model’s generality also leads to the need for sophisticated numerical methods for simulation [60].
+A key property of the HJM model is the no-arbitrage condition [60], which specifies the drift completely by the volatility:
+[…]
+Thus, in the HJM framework, the model is fully specified by defining the initial forward rate curve $f(0,T)$ and the structure of the volatility $\sigma(t,T)$.
+In our experiment we used a simple one factor HJM for illustration.
+5.2.1. PEMC for HJM
+Just as with the local volatility surface case, in practice $\sigma(t,T)$ cannot be predefined parametrically and must be calibrated from market data of caps, floors, and swaptions, yielding a discrete grid of values.
+However, for demonstration purposes, we employ a classical exponential decay specification as our baseline model inspired from [60]:
+$\sigma_{base}(t,T)=\sigma_{0}\exp(-\alpha_{\sigma}(T-t))$ (16)
+with $\sigma_{0}$ and $\alpha_{\sigma}$ as part of $\boldsymbol{\theta}$. Similarly, for a baseline initial forward curve, we use:
+$f_{base}(0,T)=f_{0}+c_{f}(1-\exp(-\alpha_{f}T)).$ (17)
+with $f_{0},c_{f}$, and $\alpha_{f}$ part of the $\boldsymbol{\theta}$.
+While this analytical form serves as our data-generating process, PEMC accesses it only through its discrete grid values with added noise.
+This approach mirrors our treatment of the local volatility surface in the previous section, where we used the parametric form in [34] solely as a realistic baseline for generating training data.
+Indeed, while HJM and these analytical forms (16)-(17) are formulated in continuous time, in practice we need to implement numerical discretization.
+Following the scheme in [60], one discretizes both the time axis and the set of maturities i.e., time steps $\mathcal{T}_{t}=\{t_{0},t_{1},\ldots,t_{N_{T}}\}$ and maturity points $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{M}}\}$ forming a time-maturity grid.
+In our experiment, for simplicity, we assume they share one grid $\mathcal{T}$ of size $|\mathcal{T}|$.
+Then, when sampling $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, one first sample $\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}$, then one sample the grids $\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}$ and $\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}$ from (16) and (17) with added noise $\mathcal{N}(0,(\frac{\sigma_{0}}{2(t_{j}+5)})^{2})$ and $\mathcal{N}(0,(\frac{1}{100(t_{j}+5)})^{2})$ respectively, on top of their baseline values.
+The noise level $\xi=\frac{\sigma_{0}}{2(t+5)}$.
+Paths are then generated using the discretization scheme described in [60] (e.g., ${f}(t_{i},t_{j})={f}(t_{i-1},t_{j})\;+\;{\mu}(t_{i-1},t_{j})(t_{i}-t_{i-1})\;+\;\sqrt{t_{i}-t_{i-1}}{\sigma}(t_{i-1},t_{j})\mathcal{N}(0,1)$ where ${\mu}(t_{i-1},t_{j})$ is the discretized drift term determined through (15)).
+We refer interested readers to [60] for the complete derivation and implementation details of this standard simulation scheme.
+For PEMC, our the parameter $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ includes the volatility parameters $(\sigma_{0},\alpha_{\sigma})$ in (16) and initial forward curve parameters $(f_{0},c,\alpha_{f})$ in (17).
+The swaption’s fixed rate is set as $R=\exp(-\frac{\sum_{i}^{n-1}f(0,t^{\prime}_{i})}{T_{\text{final}}-t^{\prime}_{0}})$, reflecting the common practice where swap rates are typically determined in reference to the prevailing forward rate curve rather than being arbitrarily chosen.
+Here the parameters are notional amount $C$, start time $t^{\prime}_{0}$, payment interval $\Delta t^{\prime}$, and number of payments $n_{p}$.
+The simulation uses time step $\Delta t$ up to final maturity $T_{\text{final}}$.
+During training, parameters are sampled uniformly from the ranges specified in Table 4 where the evaluation $\boldsymbol{\theta}$ is also listed.
+Table 4. Parameter Setup in HJM Model
+Mode $\sigma_{0}$ $\alpha_{\sigma}$ $f_{0}$ $c_{f}$ $\alpha_{f}$ $R$ $C$ $t^{\prime}_{0}$ $\Delta t^{\prime}$ $n_{p}$ $\Delta t$ $T^{*}$
+$N_{\text{train}}=3,000,000$ $[0.01,0.03]$ $[0.001,0.9]$ $[0.01,0.03]$ $[0.01,0.05]$ $[0.001,0.9]$ $\exp(-\frac{\sum_{i=0}^{n-1}f(0,t^{\prime}_{i})}{T^{*}-t^{\prime}_{0}})$ $100$ $5$ $1$ $20$ $1/52$ $25$
+evaluation $0.0015$ $100$ $0.02$ $3.0$ $0.5$
+The input for PEMC is, similar as before:
+$\text{feature}=\{\underbrace{\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}}_{\text{volatility structure}},\underbrace{\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}}_{\boldsymbol{\theta}_{\text{model}}},\underbrace{\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}}_{{\text{initial forward curve}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}}\}.$
+To handle both the two-dimensional grid of volatility structure $\sigma(t,T)$ and the one dimensional grid of initial forward curve $f(0,T)$, we design a three-branch neural network architecture, illustrated in Figure 3.
+The first branch, labeled“2D Function Encoder", processes the 2d volatility structure grid using a CNN architecture with two 2D convolutional layers, each followed by batch normalization.
+The branch concludes with an average pooling operation and produces an embedding.
+The second branch processes the initial forward curve $f(0,T)$ grid through a “1D Function Encoder" utilizing 1D convolutional layers - a natural choice for sequential data [83] - followed by batch normalization and average pooling to produce another embedding.
+The third branch handles the remaining input through fully connected layers with batch normalization.
+Finally, these three separate embeddings are then fed into a “Synthesizer" module that combines the information through multiple fully connected layers with batch normalization, ultimately producing the final prediction.
+This architecture effectively leverages both the spatial structure of the volatility surface through 2D CNNs [118], the sequential nature of the forward curve through 1D CNNs [101], and the scalar parameters through standard deep learning techniques [64].
+The complete network architecture is detailed in Table 5.
+Figure 3. Neural network architecture for modeling Swaption payoff.
+2D function branch parameters 1D function branch parameters Vector feature branch parameters
+Kernel size: (1, 3) Stride: (1, 3) Padding: 0 AvgPool2d kernel size: (2, 2) AvgPool2d stride: (2, 2) AvgPool2d padding: 0 Kernel size: 10 Stride: 3 Padding: 0 AvgPool1d kernel size: 2 AvgPool1d stride: 2 AvgPool1d padding: 0 Hidden dim: 512 Output dim: 128
+Feed-forward Synthesizer Parameters
+Hidden dim: 128
+Output dim: 1
+Table 5. Hyper-parameter setup for the neural network
+Finally, following the same evaluation methodology, we assess PEMC’s performance in the HJM swaption pricing setting across sample sizes $n\in\{1000,3000,5000,7000,9000,11000\}$ with $N=10n$, using $5\times 10^{7}$ MC samples to establish the ground truth value.
+As presented in Table 6, the outcomes from 200 independent experiments confirm that PEMC remains highly effective within the context of interest rate derivatives.
+Despite the added complexity of managing both volatility structures and forward rate curves, PEMC achieves a 45-50% reduction in RMSE to standard Monte Carlo methods.
+This performance aligns with the variance reduction levels observed in our earlier examples.
+The boxplot is shown in Figure 4.
+The consistent effectiveness of PEMC across various financial instruments and modeling frameworks further underscores its versatility as a robust variance reduction technique.
+Table 6. Root Mean Squared Error from 200 Experiments
+Method $n=1000$ $n=3000$ $n=5000$ $n=7000$ $n=9000$ $n=11000$
+Monte Carlo (MC) 0.0096 0.0062 0.0048 0.0039 0.0035 0.0029
+PEMC 0.0055 0.0028 0.0024 0.0019 0.0018 0.0015
+Figure 4. Boxplots for HJM Experiments.
+5.3. Discussions, Extensions and More Examples
+We conclude with several observations and potential extensions of the PEMC framework.
+5.3.1. Evaluation Metric
+First, while training the neural network estimator $g$ using mean squared error (MSE) loss is common, it is not always clear how to interpret the resulting MSE score.
+Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem.
+This ambiguity makes it challenging to determine when the network is sufficiently trained.
+To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$.
+If the network approximates this expectation well, then the sample average of $g(\boldsymbol{X})$ should be close to the sample average of $f(\boldsymbol{Y})$ over a given dataset.
+One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages.
+If $\mathbb{E}[g(\boldsymbol{X})]\approx\mathbb{E}[f(\boldsymbol{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation.
+Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in machine learning workflows.
+More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard Monte Carlo.
+5.3.2. XVA, Greeks and Quasi-Monte Carlo
+Beyond pricing exotic options, this approach naturally extends to other computationally intensive Monte Carlo settings in quantitative finance.
+Notably, adjustments like Credit valuation adjustment (CVA) and more complex XVAs often require vast simulation runs, sometimes taking days to complete [62, 4].
+Integrating PEMC with a well-trained neural network can cut these computational times drastically while preserving accuracy.
+Moreover, in the pricing of exotic derivatives and other path-dependent instruments, automatic differentiation (AD) techniques are often used to compute Greeks efficiently.
+By leveraging modern deep learning frameworks—such as PyTorch or TensorFlow—that support automatic differentiation natively, we can easily differentiate the trained network $g$ with respect to input parameters [59].
+This enables quick and accurate sensitivity analysis without resorting to lattice-structure approximations [133, 132] or nested simulations, which could lead to another direction of extension for PEMC.
+Finally, although we have focused on standard Monte Carlo sampling, combining PEMC with Quasi Monte Carlo (QMC) techniques may offer further variance reduction benefits [33].
+However, it is important to note that QMC methods introduce a low-discrepancy bias, thus forgoing the unbiasedness property of pure Monte Carlo.
+Balancing unbiasedness with the additional variance reduction from QMC is an intriguing direction for future research.
+5.3.3. Ambulance Diversion Policies Evaluation
+During the early COVID-19 surge, New York City’s EMS system faced a drastic spatial shift in emergency calls, pushing certain hospitals to crisis-level overload [41, 45].
+In response, a load-balancing rule was devised to divert non-critical ambulance patients away from the closest hospital if that facility nears capacity, balancing travel time against queue length through an optimization-based assignment approach [45].
+A stochastic simulation based on historical EMS data [41] was used to evaluate how load-balancing reduces peak occupancy and hospital congestion—showcasing how simulation can improve patient outcomes in EMS networks.
+The simulation workflow spans nested simulation system.
+We focus here on a representative subproblem—the two-hospital ambulance diversion— to illustrate PEMC’s advantage.
+The discrete-event simulation models two emergency departments, ED-1 and ED-2, each with physicians, patients, and arrival streams whose rates $\lambda_{h,d}$ vary by hour $h$ and day $d$.
+Ingredients include nonhomogeneous Poisson process with hourly arrival/service rates $\lambda_{h,d}$, e.g., $\text{Arrivals}(t)\sim\text{NHPP}\bigl{(}\lambda(t)\bigr{)}$; Triage levels and service times where each patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution; Priority queueing and threshold diversion where physicians are modeled as resources and patients forms priority queue ordered by triage.
+A threshold $\tau$ controls ambulance diversion: if the queue $Q$ in one hospital exceeds $\tau$, new ambulance arrivals divert to the other hospital, incurring additional travel time $\Delta$ and so on.
+The evaluation needed is how weekly mortality counts depends on the diversion threshold $\tau$.
+By varying $\tau$ in the simulation, we obtain mortality counts (or rates) as a function of policy choices, enabling downstream decisions.
+For PEMC predictor we employ a random forest estimator—well-suited for context [35, 105].
+We gather a training size $10^{5}$, and sample $\boldsymbol{\theta}$ from a $\boldsymbol{\Theta}$ calibrated from EMS data.
+The $\boldsymbol{X}$ are readily simulated components for the nested MC—drawing on Weibull, Exponential, Gamma, and Poisson distributions with $\dim(\boldsymbol{X})=12$and is also parallelizable.
+The results are summarized Figure 5 and Table7.
+Further details are left in Appendix E.
+Figure 5. Ambulance diversion policies evaluation
+Table 7. MSE and MAE for mortality evaluation at various $\tau$.
+Threshold MSE $\downarrow$ MAE $\downarrow$
+MC PEMC MC PEMC
+$n=0$ 10.974 4.142 2.765 1.587
+$n=20$ 9.535 3.266 2.497 1.416
+$n=40$ 9.174 3.122 2.428 1.376
+6. Conclusion
+In this paper, we introduced Prediction-Enhanced Monte Carlo (PEMC), a general framework that uses machine learning–based predictors as flexible control variates.
+By integrating inexpensive, highly parallelizable simulations as features, PEMC preserves the unbiasedness and error quantification of classical Monte Carlo while delivering substantial variance and runtime reductions in settings where traditional variance reduction techniques are infeasible.
+Looking ahead, several promising research directions arise.
+First, meta-learning approaches could enable PEMC to automatically learn effective control variates across related tasks, reducing manual tuning burdens [53].
+Second, recent advances in generative modeling—such as consistency models and diffusion-based methods [70, 120]—suggest the possibility of generating advanced proposal distributions or learned variates, though integration into finance requires new theoretical developments.
+Third, robust regression techniques from causal inference and distribution-shift literature could enhance PEMC’s stability under model misspecification or nonstationary data [87, 20, 48, 77, 68, 6, 109, 17, 123].
+Incorporating ideas from doubly robust estimators and orthogonal statistical learning may yield hybrid methods that further safeguard against bias.
+Finally, PEMC’s versatility invites applications beyond standard derivatives pricing: accelerating robust valuation of exotic options [99, 124], exploring equilibrium asset pricing with transaction costs [95], and solving mean field game equilibria [98].
+More ambitiously, PEMC could tackle complex financial control problems—such as optimal control of debt management [25, 26], Stackelberg equilibria in stochastic games [28, 27], and bank salvage modeling via impulse controls [38].
+Acknowledgement: We would like to thank [acknowledge individuals, organizations, or institutions] for their valuable input and support throughout this research.
+Appendix
+Appendix A Learning Theory Proofs
+In this part, we go through the pipeline for a result in the form of Lemma 6 in Section 4.3.1.
+We begin by formally introducing the key concepts associated with the learning framework under consideration.
+Definition 4.
+Let $g_{0}$ denote the true regression function, defined by
+$g_{0}(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta})):=$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))\,\bigg{|}\,(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}$
+$=$ $\operatorname*{arg\,min}_{g\text{ measurable}}\mathbb{E}_{\boldsymbol{\theta}\sim\Theta,(\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{Y}(\boldsymbol{\theta}))\sim\text{risk neutral measure}(\boldsymbol{\theta})}\biggl{(}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))-g(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\biggr{)}^{2}.$
+Note that equality exploits the well-known property of the squared loss minimizer: it is the conditional expectation of the target variable given the input features.
+The notable part is that, this formulation inherently incorporates the sampling over $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$.
+This is because, in PEMC we generate training data by first drawing $\boldsymbol{\theta}$ from $\boldsymbol{\Theta}$ and then sampling $\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{Y}(\boldsymbol{\theta})$ from the corresponding risk-neutral measure indexed by $\boldsymbol{\theta}$.
+However, since the result holds pointwise for each fixed $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, it also holds in the aggregate setting where $\boldsymbol{\theta}$ is (uniformly) random.
+For simplicity, we do not delve into technical measurability considerations here.
+Next, we introduce the hypothesis class and define the best-in-class predictor.
+The hypothesis class $\mathcal{G}$ plays a pivotal role in the learning framework, encapsulating the set of candidate functions from which the predictor $g$ is selected.
+Definition 5.
+Let $\mathcal{G}$ be the hypothesis class induced by the NN model family. Define $g^{*}$ as the best-in-class function satisfying
+$g^{*}(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta})):=$ $\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{\boldsymbol{\theta}\sim\Theta,(\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{Y}(\boldsymbol{\theta}))\sim\text{risk neutral measure}(\boldsymbol{\theta})}\biggl{(}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))-g(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\biggr{)}^{2}$
+$=$ $\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{\boldsymbol{\theta}\sim\Theta,\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}\biggl{(}(g_{0}-g)(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\biggr{)}^{2}.$
+where the second line follows again from the definition of $g_{0}$.
+Considering a model $g$ trained with $N_{\text{train}}$ samples and is held fixed during the evaluation of the expectation, as is standard when discussing generalization error, the approximation error is defined as
+$\epsilon_{a}^{\mathcal{G}}:=\mathbb{E}_{\boldsymbol{\theta}\sim\Theta,\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}\biggl{(}(g_{0}-g^{*})(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\biggr{)}^{2},$
+and the statistical error from $g$ obtained from training on $N_{\text{train}}$ samples:
+$\epsilon_{e}^{N_{\text{train}}}:=\mathbb{E}_{\boldsymbol{\theta}\sim\Theta,\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}\biggl{(}(g^{*}-g)(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\biggr{)}^{2}.$
+Finally, define the total error as
+$\epsilon_{\text{total}}:=\mathbb{E}_{\boldsymbol{\theta}\sim\Theta,\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}\biggl{(}(g_{0}-g)(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\biggr{)}^{2}.$
+It follows that
+$\epsilon_{\text{total}}\leq 2\epsilon_{a}^{\mathcal{G}}+2\epsilon_{e}^{N_{\text{train}}}.$ (18)
+The interaction between $\epsilon_{a}^{\mathcal{G}}$ and $\epsilon_{e}^{N_{\text{train}}}$ essentially captures the bias-variance tradeoff.
+By choosing more expressive model classes (e.g., richer neural network architectures) and increasing the training set size, we can jointly reduce $\epsilon_{a}^{\mathcal{G}}$ and $\epsilon_{e}^{N_{\text{train}}}$.
+Thus, under reasonable conditions, $\epsilon_{\text{total}}$ can be made arbitrarily small if both $\epsilon_{a}^{\mathcal{G}}$ and $\epsilon_{e}^{N_{\text{train}}}$.
+In the following sections, we analyze $\epsilon_{a}^{\mathcal{G}}$ and $\epsilon_{e}^{N_{\text{train}}}$ respectively.
+Appendix B Approximation Error $\epsilon_{a}^{\mathcal{G}}$
+Note that this best-in-class function $g^{*}$ depends on the distribution of $\boldsymbol{\theta}\sim\Theta$, meaning the notion of optimality is distribution-dependent on $\boldsymbol{\theta}$.
+This is also why we choose a distribution $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ that sufficiently covers the space.
+In this paper, $\mathcal{G}$ is induced by our choice of neural network architecture and training procedure.
+The complexity of $\mathcal{G}$ significantly impacts both approximation and statistical errors.
+Neural networks are renowned for their flexibility and expressiveness, serving as universal approximators.
+According to the Universal Approximation Theorem [40, 71], neural networks with a single hidden layer containing a sufficient number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary accuracy.
+Modern extensions of this theorem provide more nuanced insights into how network depth and architecture influence approximation capabilities [131, 92, 117, 40, 71].
+Typical universal universal approximation theorems deals with compact input space in $\mathbb{R}^{n}$ and point-wise convergence, in our context, we do not restrict $\boldsymbol{X}$ to live in a compact space and we only need $\mathscr{L}^{2}$ convergence, so the theorem from [104] or [86] is sufficient, which gives us
+Lemma 7.
+Suppose $g_{0}:\boldsymbol{\Theta}\times\mathbb{R}^{dim(\boldsymbol{X})}\to\mathbb{R}$ is in $\mathscr{L}^{2}(\mathbb{P})$ where $\mathbb{P}$ is the probability measure that governs the joint distribution of $\boldsymbol{\theta}\sim\Theta,(\boldsymbol{X}(\boldsymbol{\theta}),\boldsymbol{Y}(\boldsymbol{\theta}))\sim\text{risk neutral measure}(\boldsymbol{\theta})$.
+Then, for any $\epsilon>0$, we can find a class of NN $\mathcal{G}$ such that
+$\epsilon_{a}^{\mathcal{G}}\leq\epsilon.$
+Proof.
+Proof See [104] or [86]. ∎
+Appendix C Statistical Error $\epsilon_{e}^{N_{\text{train}}}$
+To quantify the capacity of $\mathcal{G}$, one could employ common complexity measures such as the Vapnik-Chervonenkis (VC) dimension [126] or Rademacher complexity [12].
+These measures provide bounds on the generalization error by capturing the richness of the hypothesis class.
+Modern studies have further refined our understanding of neural network complexity.
+For instance, norm-based capacity controls [11], PAC-Bayesian bounds [47], and local Rademacher complexities [13] offer more nuanced insights.
+In particular, most neural network architectures are known to have finite complexity measures, such as finite VC dimension or other capacity metrics, when considered as a hypothesis class with a fixed number of parameters [7].
+Consequently, when empirical risk minimization (ERM) is performed over these networks with controlled complexity, standard statistical learning theory guarantees apply, yielding generalization bounds that typically decrease on the order of $O(1/\sqrt{N_{\text{train}}})$ where $N_{\text{train}}$ is the sample size [11, 97, 8] and the constant in the rate depends on the complexity measure.
+Our PEMC prediction model can be considered as a ERM estimator.
+Thus, these results ensure that, given sufficiently large training sets and appropriate capacity constraints (e.g., weight regularization or architectural choices), neural networks can achieve low statistical error $\epsilon_{e}^{N_{\text{train}}}$.
+Finally, there is an optimization error arising from the discrepancy between the empirical risk minimizer within $\mathcal{G}$ and the final trained model $g$.
+This error accounts for situations where the optimization algorithm does not perfectly identify the empirical risk minimizer.
+However, we do not consider this optimization error, effectively assuming the presence of an ideal “oracle” for optimization.
+Consequently, theoretical analyses typically treat our $g$ as readily available empirical minimizer.
+Lemma 8.
+Suppose $g$ is the empirical risk minimizer of Algorithm 1 and the neural network class $\mathcal{G}$ has a finite VC-dimension or Rademacher complexity.
+Then, for any $\epsilon>0$, there exists $N_{\text{train}}$ such that
+$\epsilon_{e}^{N_{\text{train}}}\leq\epsilon.$
+Appendix D Proof of Lemma 6
+Lemma 7 and Lemma 8 allows us to control $\epsilon_{\text{total}}$ using (18), which allows us to prove Lemma 6 with the help of the following technical definition and lemma.
+Definition 6.
+Define $\epsilon_{\text{total}}(\boldsymbol{\theta}):=\mathbb{E}_{\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}\big{(}g(\boldsymbol{X}(\boldsymbol{\theta}))-g_{0}(\boldsymbol{X}(\boldsymbol{\theta}))\big{)}^{2}$.
+As a result, we have $\mathbb{E}_{\boldsymbol{\theta}\sim\boldsymbol{\Theta}}\epsilon_{\text{total}}(\boldsymbol{\theta})=\epsilon_{\text{total}}$, which allows us to establish the following lemma.
+Lemma 9.
+Let $G$ be a second moment bound for $g$, i.e.,
+$\mathbb{E}_{\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}g^{2}(\boldsymbol{X}(\boldsymbol{\theta}))\leq G.$
+uniformly for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Then, for each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$,
+$|\sigma_{f}^{2}-\sigma_{g}^{2}-\sigma_{f-g}^{2}|\leq 2\sqrt{\epsilon_{\text{total}}(\boldsymbol{\theta})}\sqrt{G}.$
+Proof.
+Proof Fix any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. From the variance decomposition, we have
+$\mathrm{Var}(f)=\mathrm{Var}(g)+\mathrm{Var}(f-g)+2\,\mathrm{Cov}(f-g,g),$
+which implies
+$|\sigma_{f}^{2}-\sigma_{g}^{2}-\sigma_{f-g}^{2}|\leq 2|\mathrm{Cov}(f-g,g)|.$
+Conditioning on $\boldsymbol{X}$ and utilizing the properties of $g_{0}$, we find that $\mathrm{Cov}(f-g,g)=\mathrm{Cov}(g_{0}-g,g)$.
+Applying the Cauchy–Schwarz inequality yields the desired inequality.
+∎
+We can now prove Lemma 6.
+Proof.
+Proof of Lemma 6 Besides the assumptions in Lemma 7, Lemma 8 and Lemma 9, we further assume
+$g\leq\mathbb{E}_{\boldsymbol{X}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}g^{2}(\boldsymbol{X}(\boldsymbol{\theta})),\mathbb{E}_{\boldsymbol{Y}(\boldsymbol{\theta})\sim\text{risk neutral measure}(\boldsymbol{\theta})}f^{2}(\boldsymbol{Y}(\boldsymbol{\theta}))\leq G,$ (19)
+for some $0<g<G<\infty$, uniformly for all $\theta\in\Theta$.
+Then, based on Lemma 7, 8 and Markov inequality, we can find $\mathcal{G}$ and $N_{\text{train}}$ such that, with probability at least $1-\delta$, the randomly sampled $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ satisfies
+$\epsilon_{\text{total}}(\boldsymbol{\theta})\leq\epsilon,$
+which, together with Lemma 9 gives
+$|\sigma_{f}^{2}-\sigma_{g}^{2}-\sigma_{f-g}^{2}|\leq O(\epsilon),$
+with the constant in $O$ not dependent on $\boldsymbol{\theta}$.
+Then, for at least $1-\delta$ fraction of $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, as one shrinks $\epsilon\to 0$, $\epsilon_{\text{total}}(\boldsymbol{\theta})\leq\epsilon$ and $|\sigma_{f}^{2}-\sigma_{g}^{2}-\sigma_{f-g}^{2}|\leq O(\epsilon)$ implies $\sigma_{g}^{2}\to\sigma_{g_{0}}^{2}$ and $\sigma_{f-g_{0}}^{2}\to\sigma_{f-g}^{2}$ uniformly for all such $\boldsymbol{\theta}$.
+Since $\sigma_{f}^{2}=\sigma_{g_{0}}^{2}+\sigma_{f-g_{0}}^{2}$, this further implies $\frac{\sigma_{g}}{\sigma_{f}}\to\frac{\sigma_{g_{0}}}{\sigma_{f}}$ and $\frac{\sigma_{f-g}}{\sigma_{f}}\to\frac{\sigma_{f-g_{0}}}{\sigma_{f}}$, given the boundedness of (19).
+Thus, $\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}$ in (9) $\to$ $r(\rho,c)$ in (10), or equivalently, with probability at least $1-\delta$, the randomly sampled $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ satisfies
+$\frac{\mathrm{Var}(PEMC)}{\mathrm{Var}(MC)}=r(\rho,c)+O(\epsilon),$
+where the constant in $O$ does not depend on $\boldsymbol{\theta}$. This concludes the proof. ∎
+D.1. Variance Reduction of PEMC: An Example on Asian Options
 We revisit the Asian option example and conduct experiments using the arithmetic Asian call option under the Geometric Brownian Motion (GBM) model.
 We choose the GBM model instead of the Heston model (1) to facilitate a direct comparison between PEMC, standard Monte Carlo (MC), and the traditional Control Variate (CV) method on the extent of variance reduction.
-This is because the geometric Asian option, which serves as a well-established control variate, is analytically tractable within the GBM framework but lacks a closed-form solution under the Heston model (Dingec et al.
-2015).
+This is because the geometric Asian option, which serves as a well-established control variate, is analytically tractable within the GBM framework but lacks a closed-form solution under the Heston model [44].
 This experiment aims to demonstrate the typical level of variance reduction that PEMC achieves.
-Our focus is strictly limited to variance reduction, because the GBM model actually allows for highly efficient and parallelized generation of both asset paths and corresponding payoffs, making the simulation of $\bm{Y}$ computationally straightforward.
+Our focus is strictly limited to variance reduction, because the GBM model actually allows for highly efficient and parallelized generation of both asset paths and corresponding payoffs, making the simulation of $\boldsymbol{Y}$ computationally straightforward.
 In other words, one could not fully leverage PEMC’s strengths in GBM model and this toy experiment only serves to provide insights into PEMC’s effectiveness in a controlled setting where a known CV exists.
 Practical applications and actual use-cases of PEMC will be provided in Section 5.
-3.5.1 Experimental Setup
+D.1.1. Experimental Setup
 We implement the PEMC procedure detailed in Section 3.5, utilizing a neural network (NN) as the predictive model. The asset price follows the Geometric Brownian Motion (GBM) model:
@@
 The payoff calculation $P_{A}(\{S_{t}\}_{t})=(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0)^{+}$ with $n_{D}=252$, where simulations are carried out using daily time increments (i.e., $\Delta t=1$).
-The parameter space $\bm{\theta}:=(r,S_{0},\sigma,K)\in\bm{\Theta}$ is: $r\in[0.01,0.03]$, $S_{0}\in[80,120]$, $\sigma\in[0.05,0.25]$, and $K\in[90,110]$.
-For our feature $\bm{X}$, we experiment with the whole sum of Brownian increments $W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim(\bm{X})=1$.
-We also tried a more granular representation where we partition 252 Brownian increments evenly into 14 parts, each summing 18 consecutive increments: $[\bm{X}]_{i}=\sum_{j=18(i-1)+1}^{18i}\Delta W^{S}_{j}$ for $i\in[14]$, yielding $\dim(\bm{X})=14$.
-The neural network is trained on a dataset comprising $N_{\text{train}}=1.28\times 10^{6}$ samples, generated by uniformly sampling $\bm{\theta}\sim\bm{\Theta}$.
-The neural network architecture consists of three components: a $\bm{\theta}$ network branch, a $\bm{X}$ network branch, and a combined network.
-This architecture incorporates several modern deep learning practices, including Batch Normalization (Ioffe and Szegedy 2015), Rectified Linear Unit (ReLU) activations (Nair and Hinton 2010), and skip connections (He et al.
-2016) to enhance training stability and model performance.
-The $\bm{\theta}$ network branch processes the 4-dimensional input $\bm{\theta}$ using two fully connected layers with Batch Normalization and ReLU activation functions, outputting a 10-dimensional feature vector for subsequent processing.
-The $\bm{X}$ network branch uses two fully connected layers with width $\max(32,2\dim(\bm{X}))$ neurons.
+The parameter space $\boldsymbol{\theta}:=(r,S_{0},\sigma,K)\in\boldsymbol{\Theta}$ is: $r\in[0.01,0.03]$, $S_{0}\in[80,120]$, $\sigma\in[0.05,0.25]$, and $K\in[90,110]$.
+For our feature $\boldsymbol{X}$, we experiment with the whole sum of Brownian increments $W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim(\boldsymbol{X})=1$.
+We also tried a more granular representation where we partition 252 Brownian increments evenly into 14 parts, each summing 18 consecutive increments: $[\boldsymbol{X}]_{i}=\sum_{j=18(i-1)+1}^{18i}\Delta W^{S}_{j}$ for $i\in[14]$, yielding $\dim(\boldsymbol{X})=14$.
+The neural network is trained on a dataset comprising $N_{\text{train}}=1.28\times 10^{6}$ samples, generated by uniformly sampling $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$.
+The neural network architecture consists of three components: a $\boldsymbol{\theta}$ network branch, a $\boldsymbol{X}$ network branch, and a combined network.
+This architecture incorporates several modern deep learning practices, including Batch Normalization [75], Rectified Linear Unit (ReLU) activations [96], and skip connections [64] to enhance training stability and model performance.
+The $\boldsymbol{\theta}$ network branch processes the 4-dimensional input $\boldsymbol{\theta}$ using two fully connected layers with Batch Normalization and ReLU activation functions, outputting a 10-dimensional feature vector for subsequent processing.
+The $\boldsymbol{X}$ network branch uses two fully connected layers with width $\max(32,2\dim(\boldsymbol{X}))$ neurons.
 Finally, the combined network integrates outputs from both branches through concatenation and employs skip connections featured in ResNet architectures to preserve information flow.
 The concatenated features pass through two additional fully connected layers with Batch Normalization and ReLU activation to produce the final prediction for PEMC.
-The network is implemented using PyTorch (Paszke et al.
-2019) and trained using the Adam optimizer (Kingma and Ba 2014) with a learning rate of $1\times 10^{-3}$, incorporating dropout layers (Srivastava et al.
-2014) with a rate of 0.5 after each hidden layer to prevent overfitting.
-3.5.2 Evaluation
+The network is implemented using PyTorch [103] and trained using the Adam optimizer [80] with a learning rate of $1\times 10^{-3}$, incorporating dropout layers [122] with a rate of 0.5 after each hidden layer to prevent overfitting.
+D.1.2. Evaluation
 For comparison, we implement PEMC, MC and CV method with geometric Asian option
 $P_{G}(\{S_{t}\}_{t})=((\prod_{i=1}^{n_{D}}S_{t_{i}})^{\frac{1}{{n_{D}}}}-K,0)^{+},$
-for the evaluation at $\bm{\theta}=(r,S_{0},\sigma,K)=(0.02,100,0.2,100)$. The CV estimator is given by
-$\text{CV}=\frac{1}{n}\sum_{i=1}^{n}\left(P_{A}(\{S_{t}^{(i)}\}_{t})-P_{G}(\{S_{t}^{(i)}\}_{t})\right)+P_{G}^{\text{exact}}(\bm{\theta}),$
+for the evaluation at $\boldsymbol{\theta}=(r,S_{0},\sigma,K)=(0.02,100,0.2,100)$. The CV estimator is given by
+$\text{CV}=\frac{1}{n}\sum_{i=1}^{n}\left(P_{A}(\{S_{t}^{(i)}\}_{t})-P_{G}(\{S_{t}^{(i)}\}_{t})\right)+P_{G}^{\text{exact}}(\boldsymbol{\theta}),$
 […]
-Table 1: Root Mean Squared Error from 300 Experiments
+Table 8. Root Mean Squared Error from 300 Experiments
 Method $n=1000$ $n=4000$ $n=9000$
 Monte Carlo (MC) 0.2376 0.1173 0.0854
-PEMC ($\dim\bm{X}=1$) 0.1509 0.0809 0.0481
-PEMC ($\dim\bm{X}=14$) 0.0781 0.0397 0.0261
+PEMC ($\dim\boldsymbol{X}=1$) 0.1509 0.0809 0.0481
+PEMC ($\dim\boldsymbol{X}=14$) 0.0781 0.0397 0.0261
 Geometric CV 0.0099 0.0051 0.0036
-[…]
+Figure 6.
 Performance of Estimators for Asian Options.
-Comparison of estimator across Monte Carlo (MC), PEMC, and Geometric Control Variate (CV) based on 300 experiments.
-Top: Mean squared errors (MSE) plot as a function of $n$ for (left) PEMC with $\dim\bm{X}=1$ and (right) PEMC with $\dim\bm{X}=14$, both compared against MC and Geometric CV.
+The Comparison of estimator across Monte Carlo (MC), PEMC, and Geometric Control Variate (CV) is based on 300 experiments.
+Top: Mean squared errors (MSE) plot as a function of $n$ for (left) PEMC with $\dim\boldsymbol{X}=1$ and (right) PEMC with $\dim\boldsymbol{X}=14$, both compared against MC and Geometric CV.
 Bottom: Corresponding boxplots of the 300 estimates.
@@
 The experimental results demonstrate the superior performance of the PEMC framework compared to the standard Monte Carlo (MC) estimator.
-In particular, in all sample sizes, the PEMC framework delivers visible variance reduction over MC: PEMC’s basic variant with $\dim\bm{X}=1$ achieves a 30-40% reduction in root mean squared error (RMSE) relative to MC, while the more sophisticated variant with $\dim\bm{X}=14$ attains an impressive 65-70% reduction in RMSE over MC.
+In particular, in all sample sizes, the PEMC framework delivers visible variance reduction over MC: PEMC’s basic variant with $\dim\boldsymbol{X}=1$ achieves a 30-40% reduction in root mean squared error (RMSE) relative to MC, while the more sophisticated variant with $\dim\boldsymbol{X}=14$ attains an impressive 65-70% reduction in RMSE over MC.
 However, the Geometric CV emerges as the most efficient estimator by far, with RMSE an order of magnitude smaller than MC.
@@
 However, PEMC is a flexible framework designed to enhance general simulation baselines, including sophisticated ones, not just standard Monte Carlo.
+In practice, efficient MC techniques such as Quasi-Monte Carlo (QMC) methods [33, 2] and their variants—including Langevin Monte Carlo (LMC) [90], Orthogonal Monte Carlo (OMC), and Near-Orthogonal Monte Carlo (NOMC) [89]—can be integrated with PEMC for further variance reduction.
 In particular, when a control variate method is available, such as in our current example, PEMC can be built upon it to potentially achieve even greater variance reduction.
-To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\bm{\theta})$ (Geometric CV baseline), and let the model $g$ learn
-$g(\bm{\theta},\bm{X}(\bm{\theta}))\approx$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\bm{\theta})\bigg{|}(\bm{\theta},\bm{X}(\bm{\theta}))\bigg{]}$
-$=$ $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})\bigg{|}(\bm{\theta},\bm{X}(\bm{\theta}))\bigg{]}+P_{G}^{\text{exact}}(\bm{\theta}).$ (11)
-This formulation suggests we can train $g$ on the difference of $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})$ as labels and then add back the closed-form $P_{G}^{\text{exact}}(\bm{\theta})$ to $g$.
-We implemented this approach, which we call “Boost PEMC,” using the basic PEMC variant where $\bm{X}=W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim\bm{X}=1$.
-As shown in Table 2 and Figure 3.5.2, this Boost PEMC indeed achieves variance reduction over the sophisticated Geometric CV estimator, delivering a 35-40% reduction in RMSE, remarkably consistent with the relative performance gain we observed when applying PEMC to the MC baseline in Figure 3.5.2.
-Table 2: Root Mean Squared Error from 300 Experiments
+To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\boldsymbol{\theta})$ (Geometric CV baseline), and let the model $g$ learn
+$g(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\approx$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\boldsymbol{\theta})\bigg{|}(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}$
+$=$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})\bigg{|}(\boldsymbol{\theta},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}+P_{G}^{\text{exact}}(\boldsymbol{\theta}).$ (21)
+This formulation suggests we can train $g$ on the difference of $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})$ as labels and then add back the closed-form $P_{G}^{\text{exact}}(\boldsymbol{\theta})$ to $g$.
+We implemented this approach, which we call “Boost PEMC," using the basic PEMC variant where $\boldsymbol{X}=W_{T}^{S}:=\sum_{j=1}^{252}\Delta W^{S}_{j}$ with $\dim\boldsymbol{X}=1$.
+As shown in Table 9 and Figure 7, this Boost PEMC indeed achieves variance reduction over the sophisticated Geometric CV estimator, delivering a 35-40% reduction in RMSE, remarkably consistent with the relative performance gain we observed when applying PEMC to the MC baseline in Figure 6.
+Table 9. Root Mean Squared Error from 300 Experiments
 Method $n=1000$ $n=4000$ $n=9000$
-Boost PEMC ($\dim\bm{X}=1$) 0.0065 0.0031 0.0021
+Boost PEMC ($\dim\boldsymbol{X}=1$) 0.0065 0.0031 0.0021
 Geometric CV 0.0099 0.0051 0.0036
-[…]
+Figure 7.
 Performance of CV-based Estimators for Asian Options.
@@
 In the following section, we go beyond this over-simplified setting and apply PEMC to practical applications in complex derivative pricing.
-4 Applications
-After establishing the theoretical foundations of PEMC and demonstrating its effectiveness in controlled environments, we now apply it to real-world financial scenarios where the complexity of models often renders standard variance reduction techniques difficult or unavailable.
-In this section, we examine PEMC in practice through two complex pricing problems.
-In particular, we examine variance swaps pricing under stochastic (local) volatility models (Dupire et al.
-1994, Derman and Kani 1994, Gatheral 2006), and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework (Carmona 2007).
-In both cases, the path-dependent nature of the contracts and the complexity of the model dynamics make traditional control variates difficult.
-4.1 Variance Swaps Pricing
-Variance swaps are financial derivatives that enable investors to trade future realized volatility against current implied volatility (Bayer et al.
-1999).
-Unlike traditional options, which derive their value from the underlying asset’s price, variance swaps are based on the variance of the asset’s returns over a specified period (Derman and Fantazzini 1999).
-This unique structure allows for pure exposure to volatility, making variance swaps valuable tools for risk management and speculative strategies (Gatheral 2006, Protter 2010).
-In this subsection, we present an application of the PEMC framework in pricing a variance swaps, with the underlying asset paths assumed to be generated according to a Heston model (1) and a Stochastic Local Volatility (SLV) model (Ren et al.
-2007), respectively.
-These models represent two distinct approaches to stochastic volatility modeling in practical derivatives pricing: the Heston model offers a parsimonious parametric form with relatively few parameters to calibrate, while the SLV model provides greater flexibility through a data-intensive, non-parametric specification of the volatility surface.
-By examining both cases, we demonstrate PEMC’s versatility across the spectrum of model complexity.
-4.1.1 Variance swaps under a Heston model
-Under the Heston model (1), and the procedure follows from the illustration in Section 2.5.
-Given $\bm{\theta}_{\text{model}}:=(r,\kappa,\eta,\delta,\rho)$, the asset price $\{S_{t}\}_{t}$ is simulated by Euler’s scheme with fixed $T$ and $\Delta t$:
-$\nu_{t+\Delta t}$ $\leftarrow\kappa(\eta-\nu_{t})\Delta t+\delta\sqrt{\nu_{t}}\Delta W^{\nu}_{t},$
-$S_{t+\Delta t}$ $\leftarrow S_{t}\exp\Big{(}\big{(}r-\tfrac{1}{2}\nu_{t+\Delta t}^{2}\big{)}\Delta t+\sqrt{\nu_{t+\Delta t}}\Delta W^{S}_{t}\Big{)}.$
-where $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
-For each asset path $\{S_{t}\}_{t}$, the payoff is
-$f_{\bm{\theta}_{\text{payoff}}}(\{S_{t}\}_{t}):=\sqrt{\frac{252\times\sum_{t=1}^{T/\Delta t}\log(S_{t}/S_{t-1})^{2}}{T/\Delta t}}-K.$ (12)
-Note that due to the linear form of strike $K$ in the payoff, it suffices to effectively removes strike $K$ from training.
-We also conveniently omit any discount factor.
-The parameters space $\bm{\Theta}$, as well as the evaluation $\bm{\theta}$, is summarized in Table 4.
-Table 3: Parameter Setup in Heston Model
-mode $S_{0}$ $r$ $\nu_{0}$ $\kappa$ $\eta$ $\delta$ $\rho$ $T$, $\Delta t$
-$N_{\text{train}}=3,000,000$ $[50,150]$ $[0.01,0.05]$ $[0.1,0.375]^{2}$ $[1.5,4.5]$ $[0.1,0.3]^{2}$ $[0.1,1.0]$ $[-0.2,-0.9]$ $1$, $1/252$
-evaluation $100$ $0.02$ $0.25^{2}$ $3.0$ $0.2^{2}$ $0.6$ $-0.4$
-The feature $\bm{X}$ is constructed by
-$W_{T}^{S}:=\sum_{j=1}^{T/\Delta t}\Delta W^{S}_{j\Delta t}\quad\text{ and }\quad W_{T}^{\nu}:=\sum_{j=1}^{T/\Delta t}\Delta W^{\nu}_{j\Delta t},$
-which are jointly distributed as $\sqrt{T}\cdot\mathcal{N}(0,[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$ just as in Section 2.5.
-For the prediction model training in PEMC algorithm 1, a basic Residual Network (ResNet)-style architecture (He et al.
-2016) is adopted and trained on $N_{\text{train}}=3\times 10^{6}$ samples.
-The network consists of two Multi-Layer Perceptron (MLP) blocks connected through skip connections.
-The first MLP block processes $(\bm{\theta},\bm{X})$ together and pass them through hidden layers with dimension 512, eventually producing a 256-dimensional intermediate representation.
-This output is then fed into the second MLP block, which maintains the same hidden dimension of 256 throughout its layers.
-The complete architecture specification is provided in Table 4.
-At evaluation stage (i.e., Algorithm 2), we use $\bm{\theta}$ in Table 4 and set $N=10n$ as before.
-Table 4: Hyper-parameter Setup for Neural Network
-1st MLP block 2nd MLP Block
-hidden dim: 512; output dim: 256 hidden dim: 256
-Following our evaluation methodology from Section 3.5, we first establish a ground truth value by averaging $5\times 10^{7}$ MC samples.
-We then implement both standard MC and PEMC estimators (with $N=10n$) using sample sizes $n\in\{1000,2000,4000,8000,10000,20000\}$, repeating each experiment 200 times to assess sampling variability.
-As shown in Table 5, PEMC achieves a 30-50% reduction in root mean squared error compared to standard MC, consistent with the variance reduction levels we observed in Section 3.5.
-This demonstrates that PEMC’s effectiveness as a variance reduction technique carries over to more complex models and derivative products.
-Table 5: RMSE from 200 Experiments for Pricing Variance Swap under the Heston Model
-Method $n=1000$ $n=2000$ $n=4000$ $n=6000$ $n=8000$ $n=10000$ $n=20000$
-Monte Carlo (MC) 0.0996 0.0659 0.0517 0.0391 0.0366 0.0290 0.0218
-PEMC 0.0599 0.0387 0.0263 0.0233 0.0213 0.0179 0.0126
-4.1.2 Variance swaps under an SLV model
-Having demonstrated PEMC’s effectiveness for the Heston model in Section 4.1.1, we turn to the more sophisticated Stochastic Local Volatility (SLV) framework.
-While parametric models like Heston offer mathematical elegance and computational efficiency, they often lack the flexibility to fully capture market dynamics (Gatheral 2006).
-SLV models emerged as a hybrid approach that combines the market-implied local volatility surface with stochastic volatility dynamics (Guyon and Henry-Labordère 2014), providing practitioners with greater calibration flexibility and more accurate price reproduction across strike-maturity ranges (Ren et al.
-2007).
-The increased modeling power of SLV, however, comes with significant computational challenges.
-It is known that SLV has no closed-form formula for even vanilla options (Tataru and Fisher 2010).
-Moreover, unlike the Heston model’s parsimonious parameter set, SLV requires handling a full volatility surface discretized on a dense 2D grid, effectively turning our parameter space $\bm{\Theta}$ for PEMC into a high-dimensional object.
-To efficiently process this grid-structured volatility data, we adopt a Convolutional Neural Network (CNN) architecture (LeCun et al.
-1998) for our PEMC implementation, effectively feeding the local volatility surface as an image into PEMC.
-The CNNs are particularly well-suited for this task as they naturally exploit the spatial relationships in the volatility surface, similar to their success in image processing tasks (He et al.
-2016, Krizhevsky et al.
-2012).
-Under the SLV model, we adopt the following SDEs for simulating the asset price $\{S_{t}\}_{t}$:
-$\mathrm{d}S_{t}$ $=rS_{t}\mathrm{d}t+\sigma(S_{t},t)e^{\nu_{t}}S_{t}\mathrm{d}W_{t}^{S},$ (13)
-$\mathrm{d}\nu_{t}$ $=\kappa(\eta_{t}-\nu_{t})\mathrm{d}t+\delta\mathrm{d}W_{t}^{\nu},$ (14)
-with $\langle\mathrm{d}W_{t}^{S},\mathrm{d}W_{t}^{\nu}\rangle=\rho$ and $\eta_{t}:=-\frac{\delta^{2}}{2\kappa}(1+e^{-2\kappa t}).$ Here $\sigma(\cdot,\cdot):\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{+}$ is 2D function representing the local volatility surface and $\exp(\nu_{t})$ is a stochastic multiplier with $\exp(\nu_{0})=1$ and $\mathbb{E}[e^{2\nu_{t}}]=1$.
-In practice, local volatility surfaces are calibrated to and stored as discrete two-dimensional grids indexed by asset prices (spot) and time.
-During simulation, these discrete values are interpolated as needed to obtain volatilities at arbitrary price-time points (Coleman et al.
-2001).
-For our PEMC implementation, we treat this discrete grid as part of our input parameter $\bm{\theta}_{\text{model}}$, reflecting how the market-calibrated surfaces would be used in practice.
-While local volatility models (Dupire et al.
-1994, Derman and Kani 1994) and their calibration to market data constitute an extensive research area in their own right, our focus here is on the PEMC implementation where, a calibrated local volatility has been given for evaluation, regardless of the calibration method used to obtain it.
-Following practical conventions but also simplifying for the sake of illustration, we store $\sigma(\cdot,\cdot)$ on a $|\mathcal{S}|\times|\mathcal{T}|$ grid, where $\mathcal{S}$ contains $|\mathcal{S}|$ equally-spaced price points in $[S_{\text{surface}}^{\min},S_{\text{surface}}^{\max}]$ and $\mathcal{T}$ contains $|\mathcal{T}|$ equally-spaced time points in $[t_{\text{surface}}^{\min},t_{\text{surface}}^{\max}]$.
-At each point on the grid, we store the value of local volatility according to Carmona (2007):
-$\sigma_{base}^{2}(x,t)=\frac{\sum_{i=0}^{2}p_{i}\tau_{i}e^{-x^{2}/(2t\tau_{i}^{2})-t\tau^{2}_{i}/8}}{\sum_{i=0}^{2}(p_{i}/\tau_{i})e^{-x^{2}/(2t\tau_{i}^{2})-t\tau_{i}^{2}/8}},\quad\text{with}~{}~{}~{}p_{0}:=1-p_{1}-p_{2},~{}~{}x:=\log(S_{t}/S_{0}),$ (15)
-for $p_{0}=0.3,p_{1}=0.5,p_{2}=0.2,\tau_{0}=0.4,\tau_{1}=0.3,\tau_{2}=0.6$ as in Figure 2 in Carmona (2007).
-While this analytical form (15) is used in our data-generating process, it is important to note that PEMC treats the surface as any market-calibrated volatility surface - accessing it only through its discrete grid values.
-When sampling $\bm{\theta}\sim\bm{\Theta}$ to produce $\sigma$ grid, we add a $\mathcal{N}(0,\xi^{2})$ noise independently to all the $|\mathcal{S}|\times|\mathcal{T}|$ points in the grid, on top of their baseline value $\sigma_{base}$ (15).
-A path is then generated by Euler’s scheme:
-$\nu_{t+\Delta t}$ $\leftarrow\nu_{t}+\kappa(\eta_{t}-\nu_{t})\Delta t+\delta\Delta W^{\nu}_{t},$
-$S_{t+\Delta t}$ $\leftarrow S_{t}\exp\Big{(}\big{(}r-\tfrac{1}{2}\tilde{\nu}_{t+\Delta t}^{2}\big{)}\Delta t+\tilde{\nu}_{t+\Delta t}\Delta W^{S}_{t}\Big{)},\quad\text{where}~{}~{}\tilde{\nu}_{t+\Delta t}=\widehat{\sigma}(S_{t},t)e^{\nu_{t+\Delta t}},$
-with $(\Delta W^{S}_{t},\Delta W^{\nu}_{t})\stackrel{{\scriptstyle\text{i.i.d}}}{{\sim}}\sqrt{\Delta t}\cdot\mathcal{N}(\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}])$, and $\widehat{\sigma}(S_{t},t)$ obtained by interpolations of the grid $\sigma$.
-The input $(\bm{\theta},\bm{X})$ can be represented as
-$\text{feature}_{i}=\{\underbrace{\{\sigma^{(s,t)}\}_{s\in\mathcal{S},t\in\mathcal{T}}}_{\text{surface info }},\underbrace{r,\delta,\kappa,\rho,\mu}_{\bm{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0}}_{\bm{\theta}_{\text{simulation}}},\underbrace{K}_{\bm{\theta}_{\text{payoff}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}(\bm{\theta})}\}.$
-The parameters space $\bm{\Theta}$, as well as the evaluation $\bm{\theta}$, is summarized in Table 6.
-Table 6: Parameter Setup in SLV Model
-mode $\xi$ $S_{0}$ $r$ $\kappa$ $\lambda$ $\rho$ $T$, $\Delta t$
-$N_{\text{train}}=3,000,000$ $0.02$ $[50,150]$ $0.02$ $[1.5,4.5]$ $[0.1,1.0]$ $[-0.2,-0.9]$ $1$, $1/252$
-evaluation 00 $100$ $0.02$ $3.0$ $0.5$ $-0.5$
-To handle the 2D grid of high-dimensional volatility surface data, we design a two-branch neural network architecture, which is illustrated in Figure 1.
-The first branch processes the discretized volatility surface $\sigma^{2}(x,t)$ using a CNN architecture inspired by VGG (Simonyan and Zisserman 2014), which has become a standard choice for image processing tasks and is readily available in modern deep learning packages like PyTorch (Paszke et al.
-2019).
-This branch then consists of two 2D convolutional layers interspersed with ReLU activations, followed by a MaxPool2d operation for dimensionality reduction.
-The surface features are then flattened through a fully connected layer to produce an embedding.
-The second branch handles the remaining model parameters in $(\bm{\theta},\bm{X})$ through a series of fully connected layers with dropout regularization, batch normalization, and ReLU activations, ultimately producing another embedding.
-Finally, the two separate embeddings are then fed into a “Synthesizer” module, which combines the information through additional fully connected layers with dropout and ReLU activations to produce the final prediction.
-This architecture choice is motivated by the proven effectiveness of CNNs in handling grid-structured data (LeCun et al.
-1998, He et al.
-2016), and particularly the VGG architecture’s success in extracting hierarchical features while maintaining relative simplicity (Simonyan and Zisserman 2014).
-The details of NN architecture are summarized in Table 7.
-Figure 1: Neural Network architecture for the SLV Model.
-Table 7: Neural Network Architecture Parameters
-CNN Branch Feed-forward Branch
-kernel size: 3, stride: 1, padding: 1
-max pooling (kernel: 2, stride: 2, padding: 0)
-hidden dim: 512
-output dim: 128
-Using the same experimental setup as before, we evaluate PEMC’s performance in the SLV setting across sample sizes $n\in\{1000,2000,4000,8000,10000,20000\}$ with $N=10n$, benchmarking against a ground truth computed from $5\times 10^{7}$ MC samples.
-The results, based on 200 independent experiments and shown in Table 8, reveal that PEMC’s effectiveness persists even in this more complex setting.
-Despite the added complexity of handling high-dimensional volatility surfaces, PEMC achieves a 30-40% reduction in mean squared error compared to standard MC, demonstrating its robustness as a variance reduction technique across different model frameworks.
-Table 8: RMSE from 200 Experiments for Pricing Variance Swap under the SLV Model
-Method $n=1000$ $n=2000$ $n=4000$ $n=6000$ $n=8000$ $n=10000$ $n=20000$
-Monte Carlo (MC) 0.0206 0.0145 0.0101 0.0075 0.0064 0.0065 0.0047
-PEMC 0.0130 0.0088 0.0061 0.0055 0.0040 0.0040 0.0027
-4.2 Swaptions in HJM Models
-Interest rate derivatives play a crucial role in financial markets, with swaptions being particularly important instruments for managing interest rate risk (Brigo and Mercurio 2006).
-A swaption gives its holder the right to enter into an interest rate swap at a future date, providing flexibility in hedging future interest rate exposures (Hull 2018).
-The pricing of these instruments, however, presents significant computational challenges due to the high-dimensional nature of interest rate modeling (Andersen and Piterbarg 2010).
-In this subsection, we demonstrate PEMC’s application to swaption pricing under the Heath-Jarrow-Morton (HJM) framework (Heath et al.
-1992).
-The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models.
-For illustration purposes, we focus on a one-factor specification with exponential volatility structure (Glasserman 2013b), though the framework readily extends to multi-factor cases.
-This model choice represents a balance between analytical tractability and practical relevance, allowing us to demonstrate PEMC’s effectiveness in a high-dimensional setting while maintaining computational feasibility.
-4.2.1 Swaptions and HJM Model
-A swaption is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date.
-In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite.
-Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$.
-The value of this swap at time $t^{\prime}_{0}$ is:
-$V_{t^{\prime}_{0}}=C\left(R\sum_{l=1}^{n_{p}}B(t^{\prime}_{0},t^{\prime}_{l})\Delta t^{\prime}+B(t^{\prime}_{0},t^{\prime}_{n_{p}})-1\right),$
-where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}$.
-A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$.
-The payoff of the swaption is simly
-$\max(0,V_{t^{\prime}_{0}}),$
+Appendix E Ambulance Diversion
+Details can be further found in [41, 45].
+Some key ingredients are 1.
+Patient Arrivals: We simulate patient inflow to each hospital (A or B) according to a nonhomogeneous Poisson process with hourly rates $\lambda_{h,d}$.
+Specifically:
+$\text{Arrivals}(t)\sim\text{NHPP}\bigl{(}\lambda(t)\bigr{)},$
+where $\lambda(t)$ is determined by a piecewise schedule over each hour $h$ of day $d$.
+Many of these arrivals (e.g., 25%) come by ambulance, with the remainder as walk-ins.
+2.
+Triage Levels and Service Times: Each arriving patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution:
+$\Pr[L=\ell]=p_{\ell}\quad\text{or}\quad p_{\ell}^{\text{(crisis)}}.$
 […]
-$df(t,T)=\mu(t,T)\,dt+\sigma(t,T)^{\top}dW(t),$
-where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion.
-In contrast to short-rate models (e.g., Vasicek (Vasicek 1977) or Cox-Ingersoll-Ross (Cox et al.
-1985)), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates (Heath et al.
-1992).
-The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures (Brigo and Mercurio 2006, Andersen and Piterbarg 2010).
-However, the model’s generality also leads to the need for sophisticated numerical methods for simulation Glasserman (2013b).
-A key property of the HJM model is the no-arbitrage condition (Glasserman 2013b), which specifies the drift completely by the volatility:
+$\text{divert}(\tau)=\begin{cases}1,&\text{if }Q>\tau,\\ 0,&\text{otherwise}.\end{cases}$
+4. Mortality Function Each patient $i$ has a time-varying risk of death. The code uses two complementary representations: 1. Direct Mortality Probability:
 […]
-Thus, in the HJM framework, the model is fully specified by defining the initial forward rate curve $f(0,T)$ and the structure of the volatility $\sigma(t,T)$.
-In our experiment we used a simple one factor HJM for illustration.
-4.2.2 PEMC for HJM
-Just as with the local volatility surface case, in practice $\sigma(t,T)$ cannot be predefined parametrically and must be calibrated from market data of caps, floors, and swaptions, yielding a discrete grid of values.
-However, for demonstration purposes, we employ a classical exponential decay specification as our baseline model inspired from Glasserman (2013b):
-$\sigma_{base}(t,T)=\sigma_{0}\exp(-\alpha_{\sigma}(T-t))$ (17)
-with $\sigma_{0}$ and $\alpha_{\sigma}$ as part of $\bm{\theta}$. Similarly, for a baseline initial forward curve, we use:
-$f_{base}(0,T)=f_{0}+c_{f}(1-\exp(-\alpha_{f}T)).$ (18)
-with $f_{0},c_{f}$, and $\alpha_{f}$ part of the $\bm{\theta}$.
-While this analytical form serves as our data-generating process, PEMC accesses it only through its discrete grid values with added noise.
-This approach mirrors our treatment of the local volatility surface in the previous section, where we used the parametric form in Carmona (2007) solely as a realistic baseline for generating training data.
-Indeed, while HJM and these analytical forms (17)-(18) are formulated in continuous time, in practice we need to implement numerical discretization.
-Following the scheme in Glasserman (2013b), one discretizes both the time axis and the set of maturities i.e., time steps $\mathcal{T}_{t}=\{t_{0},t_{1},\ldots,t_{N_{T}}\}$ and maturity points $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{M}}\}$ forming a time-maturity grid.
-In our experiment, for simplicity, we assume they share one grid $\mathcal{T}$ of size $|\mathcal{T}|$.
-Then, when sampling $\bm{\theta}\sim\bm{\Theta}$, one first sample $\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}$, then one sample the grids $\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}$ and $\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}$ from (17) and (18) with added noise $\mathcal{N}(0,(\frac{\sigma_{0}}{2(t_{j}+5)})^{2})$ and $\mathcal{N}(0,(\frac{1}{100(t_{j}+5)})^{2})$ respectively, on top of their baseline values.
-The noise level $\xi=\frac{\sigma_{0}}{2(t+5)}$.
-Paths are then generated using the discretization scheme described in Glasserman (2013b) (e.g., ${f}(t_{i},t_{j})={f}(t_{i-1},t_{j})\;+\;{\mu}(t_{i-1},t_{j})(t_{i}-t_{i-1})\;+\;\sqrt{t_{i}-t_{i-1}}{\sigma}(t_{i-1},t_{j})\mathcal{N}(0,1)$ where ${\mu}(t_{i-1},t_{j})$ is the discretized drift term determined through (16)).
-We refer interested readers to Glasserman (2013b) for the complete derivation and implementation details of this standard simulation scheme.
-For PEMC, our the parameter $\bm{\theta}\sim\bm{\Theta}$ includes the volatility parameters $(\sigma_{0},\alpha_{\sigma})$ in (17) and initial forward curve parameters $(f_{0},c,\alpha_{f})$ in (18).
-The swaption’s fixed rate is set as $R=\exp(-\frac{\sum_{i}^{n-1}f(0,t^{\prime}_{i})}{T_{\text{final}}-t^{\prime}_{0}})$, reflecting the common practice where swap rates are typically determined in reference to the prevailing forward rate curve rather than being arbitrarily chosen.
-Here the parameters are notional amount $C$, start time $t^{\prime}_{0}$, payment interval $\Delta t^{\prime}$, and number of payments $n_{p}$.
-The simulation uses time step $\Delta t$ up to final maturity $T_{\text{final}}$.
-During training, parameters are sampled uniformly from the ranges specified in Table 9 where the evaluation $\bm{\theta}$ is also listed.
-Table 9: Parameter Setup in HJM Model
-Mode $\sigma_{0}$ $\alpha_{\sigma}$ $f_{0}$ $c_{f}$ $\alpha_{f}$ $R$ $C$ $t^{\prime}_{0}$ $\Delta t^{\prime}$ $n_{p}$ $\Delta t$ $T^{*}$
-$N_{\text{train}}=3,000,000$ $[0.01,0.03]$ $[0.001,0.9]$ $[0.01,0.03]$ $[0.01,0.05]$ $[0.001,0.9]$ $\exp(-\frac{\sum_{i=0}^{n-1}f(0,t^{\prime}_{i})}{T^{*}-t^{\prime}_{0}})$ $100$ $5$ $1$ $20$ $1/52$ $25$
-evaluation $0.0015$ $100$ $0.02$ $3.0$ $0.5$
-The input for PEMC is, similar as before:
-$\text{feature}=\{\underbrace{\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}}_{\text{volatility structure}},\underbrace{\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}}_{\bm{\theta}_{\text{model}}},\underbrace{\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}}_{{\text{initial forward curve}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}}\}.$
-To handle both the two-dimensional grid of volatility structure $\sigma(t,T)$ and the one dimensional grid of initial forward curve $f(0,T)$, we design a three-branch neural network architecture, illustrated in Figure 2.
-The first branch, labeled“2D Function Encoder”, processes the 2d volatility structure grid using a CNN architecture with two 2D convolutional layers, each followed by batch normalization.
-The branch concludes with an average pooling operation and produces an embedding.
-The second branch processes the initial forward curve $f(0,T)$ grid through a “1D Function Encoder” utilizing 1D convolutional layers - a natural choice for sequential data (LeCun et al.
-1998) - followed by batch normalization and average pooling to produce another embedding.
-The third branch handles the remaining input through fully connected layers with batch normalization.
-Finally, these three separate embeddings are then fed into a “Synthesizer” module that combines the information through multiple fully connected layers with batch normalization, ultimately producing the final prediction.
-This architecture effectively leverages both the spatial structure of the volatility surface through 2D CNNs (Simonyan and Zisserman 2014), the sequential nature of the forward curve through 1D CNNs (Oord et al.
-2016), and the scalar parameters through standard deep learning techniques (He et al.
-2016).
-The complete network architecture is detailed in Table 10.
-Figure 2: Neural network architecture for modeling Swaption payoff.
-2D function branch parameters 1D function branch parameters Vector feature branch parameters
-Kernel size: (1, 3) Stride: (1, 3) Padding: 0 AvgPool2d kernel size: (2, 2) AvgPool2d stride: (2, 2) AvgPool2d padding: 0 Kernel size: 10 Stride: 3 Padding: 0 AvgPool1d kernel size: 2 AvgPool1d stride: 2 AvgPool1d padding: 0 Hidden dim: 512 Output dim: 128
-Feed-forward Synthesizer Parameters
-Hidden dim: 128
-Output dim: 1
-Table 10: Hyper-parameter setup for the neural network
-Finally, following the same evaluation methodology, we assess PEMC’s performance in the HJM swaption pricing setting across sample sizes $n\in\{1000,3000,5000,7000,9000,11000\}$ with $N=10n$, using $5\times 10^{7}$ MC samples to establish the ground truth value.
-As presented in Table 11, the outcomes from 200 independent experiments confirm that PEMC remains highly effective within the context of interest rate derivatives.
-Despite the added complexity of managing both volatility structures and forward rate curves, PEMC achieves a 45-50% reduction in RMSE to standard Monte Carlo methods.
-This performance aligns with the variance reduction levels observed in our earlier examples.
-The boxplot is shown in Figure 4.2.2.
-The consistent effectiveness of PEMC across various financial instruments and modeling frameworks further underscores its versatility as a robust variance reduction technique.
-Table 11: Root Mean Squared Error from 200 Experiments
-Method $n=1000$ $n=3000$ $n=5000$ $n=7000$ $n=9000$ $n=11000$
-Monte Carlo (MC) 0.0096 0.0062 0.0048 0.0039 0.0035 0.0029
-PEMC 0.0055 0.0028 0.0024 0.0019 0.0018 0.0015
-Boxplots for HJM Experiments.
-4.3 Discussions and Extensions
-We conclude with several practical observations and potential extensions of the PEMC framework.
-4.3.1 Evaluation Metric
-First, while training the neural network estimator $g$ using mean squared error (MSE) loss is common, it is not always clear how to interpret the resulting MSE score.
-Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem.
-This ambiguity makes it challenging to determine when the network is sufficiently trained.
-To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$.
-If the network approximates this expectation well, then the sample average of $g(\bm{X})$ should be close to the sample average of $f(\bm{Y})$ over a given dataset.
-One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages.
-If $\mathbb{E}[g(\bm{X})]\approx\mathbb{E}[f(\bm{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation.
-Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in machine learning workflows.
-More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard Monte Carlo.
-4.3.2 XVA, Greeks and Quasi-Monte Carlo
-Beyond pricing exotic options, this approach naturally extends to other computationally intensive Monte Carlo settings in quantitative finance.
-Notably, adjustments like Credit valuation adjustment (CVA) and more complex XVAs often require vast simulation runs, sometimes taking days to complete (Green et al.
-2015, Andersen et al.
-2017).
-Integrating PEMC with a well-trained neural network can cut these computational times drastically while preserving accuracy.
-Moreover, in the pricing of exotic derivatives and other path-dependent instruments, automatic differentiation (AD) techniques are often used to compute Greeks efficiently.
-By leveraging modern deep learning frameworks—such as PyTorch or TensorFlow—that support automatic differentiation natively, we can easily differentiate the trained network $g$ with respect to input parameters (Giles and Glasserman 2006).
-This enables quick and accurate sensitivity analysis without resorting to finite-difference approximations or nested simulations, which could lead to another direction of extension for PEMC.
-Finally, although we have focused on standard Monte Carlo sampling, combining PEMC with Quasi Monte Carlo (QMC) techniques may offer further variance reduction benefits (Caflisch 1998).
-However, it is important to note that QMC methods introduce a low-discrepancy bias, thus forgoing the unbiasedness property of pure Monte Carlo.
-Balancing unbiasedness with the additional variance reduction from QMC is an intriguing direction for future research.
-5 Conclusion
-In this paper, we introduced the Prediction-Enhanced Monte Carlo (PEMC) framework, a novel approach that leverages machine learning-based predictors to serve as flexible, data-driven control variates.
-By doing so, we preserve the unbiasedness of Monte Carlo estimators while significantly enhancing their efficiency in complex settings where traditional variance reduction techniques such as control variates are difficult or impossible to construct.
-We demonstrated the practical value of PEMC through applications: pricing variance swaps under both a Heston and a Stochastic Local Volatility model, and pricing swaptions under the Heath-Jarrow-Morton framework.
-These examples, each introducing their own modeling and computational challenges, highlight PEMC’s versatility and effectiveness.
-While we have demonstrated the power of the PEMC framework on an options pricing application, we believe PEMC represents a new frontier for variance reduction, blending modern machine learning methodologies with classical Monte Carlo simulation.
-A promising avenue for future research is to explore the connection between PEMC and meta-learning, which seeks to develop models capable of learning how to construct effective control variates from related problems.
-This approach reduces the need for problem-specific tuning, as demonstrated by Finn et al.
-(2017).
-Additionally, the recent surge in consistency models and diffusion-based generative techniques in machine learning (Ho et al.
-(2020), Song et al.
-(2021)) opens up new possibilities.
-While their direct application to pricing and control variates remains largely unexplored, these flexible generative models could, in theory, produce advanced proposal distributions or control variates.
-Nonetheless, significant methodological and theoretical gaps must be bridged before such methods can be integrated seamlessly and rigorously into financial computation pipelines.
-Another intriguing direction lies in leveraging recent advancements in causal inference and robust statistics, particularly methods designed to handle shifts in underlying data distributions (see Li and Lam (2020), Bodnar et al.
-(2022), Ellickson et al.
-(2023), Kang and Schafer (2007), Hernán and Robins (2020)).
-For instance, studies such as Angelopoulos et al.
-(2023), Li and Lam (2020), Quionero-Candela et al.
-(2009), Bickel et al.
-(2009), Sugiyama and Kawanabe (2009) propose robust regression techniques under covariate shift.
-These advances suggest that methodologies initially developed for causal inference—designed to ensure stability against model misspecification and distributional drift—could inform the creation of more versatile and robust variance reduction schemes.
-Incorporating ideas from doubly robust estimators into the ML-based control variate framework could pave the way for hybrid methods like PEMC to better handle uncertain or dynamic environments.
+[…]
+$\min_{\tau}\;\mathbb{E}\Bigl{[}\sum_{\text{patients }i}\mathbf{1}\bigl{\{}\text{wait}_{i}>\text{death\_time}_{i}\bigr{\}}\Bigr{]},$
+subject to feasible travel times and doctor schedules.
+To validate the proposed load-balancing rule, we develop a discrete-event simulation reflecting real-time dynamics of New York City’s EMS system.
+Patient arrivals are modeled by a nonhomogeneous Poisson process, $\lambda(t)$, calibrated to historical data from the pandemic’s peak.
+Upon arrival, each patient is assigned a triage level $\ell\in\{1,\dots,5\}$ via a multinomial distribution that shifts according to COVID surges.
+Service times in the emergency department follow exponential distributions with rate $\mu_{\ell}$.
+At each event, if the queue length at a chosen hospital exceeds a threshold, the ambulance is diverted to an alternative facility, trading off increased travel against avoiding overcrowding.
+By iterating this simulation over a range of thresholds and arrival scenarios, we optimize a load-balancing policy that significantly reduces hospital overloading and shortens ED wait times.
+Specifically, our discrete-event simulation (spanning a one-week horizon) envisions two hospitals (A and B) under nonhomogeneous Poisson arrivals—with "normal" or "crisis" modes dictating time-varying rates—and assigns each patient a triage level from ${1,2,3,4,5}$.
+Service rates $(\mu_{\ell}^{A},\mu_{\ell}^{B})$ differ by hospital and triage $\ell$, while death times follow a Weibull distribution parameterized by $(\alpha,\beta)$.
+Our decision variable is a threshold $\tau$, triggering ambulance diversions whenever either hospital’s queue exceeds $\tau$.
+We optimize $\tau$ to minimize total mortality, measured by the fraction of patients whose wait time surpasses their Weibull survival limit.
+Experiments show that, particularly in crisis mode, adaptive thresholding significantly reduces overloading and mortality compared to the naive "always-nearest-hospital" policy.
+Let $\boldsymbol{\lambda}=(\lambda_{h,d})$ denote the matrix of arrival rates across hours and days.
+Patients arrive with triage levels $L\in\{1,\dots,5\}$ (where 1 is most critical), drawn from multinomial probabilities $\{p_{\ell},p_{\ell}^{\text{crisis}}\}$.
+Within each ED, service times for triage $\ell$ follow exponential distributions $\exp(\mu_{\ell})$, and mortality follows a logistic model: $\mathbb{P}(\text{death}\mid L=\ell,W=w)=\frac{1}{1+\exp(-(\beta_{0,\ell}+\beta_{1,\ell}w))}$.
+Notably, 25% of arrivals come by ambulance.
+If an ED’s queue length $Q$ exceeds a threshold $\tau$, any new ambulances divert to the other ED, incurring extra travel time.
+In a data-adaptive version, $\tau$ is updated each shift using predicted arrival rates.
+All model inputs—arrival rates, triage probabilities, and service rates—come from empirical data [130], while the logistic mortality parameters $\beta_{0,\ell}$, $\beta_{1,\ell}$ are tuned (with minor per-patient perturbations) to capture heterogeneous risk profiles.
+In this example, we construct the feature of
+[…]
+[…]
+[…]
+[…]
+[…]
+The blue features are $\boldsymbol{\theta}$ while the purple ones are $\boldsymbol{X}$.
+The simulation starts by first sampling the number of patients, which follows a Poisson distribution with parameter of $\lambda T$, denoted by $N_{p}$.
+hopsital1_max_patience and hopsital2_max_patience can be simulated with max of weilbull distribution, and max_patience can be computed by taking the max of these quantities (which has closed-form density function, i.e., maximum of Weibull’s).
+total_service_time can be simulated by first sampling service time for $N_{p}$ times with $\exp(\lambda_{\text{service}})$ and summing up (which has closed-form density function, i.e., sum of exponential is Gamma ).
+total_death_time can be simulated by taking first sampling individual using inversion sampling from $F^{-1}(u)$, where $F^{-1}:$
+$F^{-1}_{t,B,\nu}(u)=\begin{cases}0,&\text{if }u\leq\text{mortality\_func}(0,t,B,\nu)\\[10.0pt] \infty,&\text{if }u\geq K[t]\\[10.0pt] -\frac{\ln\left(\left(\frac{K[t]-A[t]}{u-A[t]}\right)^{(\nu+0.25t)}-1\right)}{3t(B+5-t)}+\frac{(-2+2.5a)t}{(B+5-t)},&\text{otherwise}.\end{cases}$
+$K=\{1:1.0,2:0.9,3:0.05,4:0.02,5:0.01\}$
+$A=\{1:0.6,2:0.1,3:0.0,4:0.0,5:0.0\}$
+[…]
+where $B\sim\text{Unif}(2.5,3.5)$, $\nu\sim\text{Unif}(1.5,2.5)$, and the keys of $K$ and $A$ are the triage levels.
+Triage count of each level can be simulated by firstly sampling from multi-nominal distribution with probabilities from triage 1 to 5 as $\{0.1098,0.2761,0.4596,0.1297,0.0248\}$, and then counting the number of each category.
+The quantity we are estimating is the mortality.
+The PEMC predictor is a random forest.
+The amount of training data is $10^{5}$.
+During the pre-training stage, we collect the training data by the sampling procedure in Table 10.
+Variable Distribution
+$\tau$ $\text{Unif}\{0,50\}$
+hosp2_doctor_shift_counts_1 $\text{Unif}\{1,3\}$
+hosp2_doctor_shift_counts_2 $\text{Unif}\{1,3\}$
+hosp2_doctor_shift_counts_3 $\text{Unif}\{2,6\}$
+hosp2_doctor_shift_counts_4 $\text{Unif}\{1,4\}$
+hosp2_doctor_shift_counts_5 $\text{Unif}\{2,5\}$
+hosp2_doctor_shift_counts_6 $\text{Unif}\{1,3\}$
+crisis $\text{Unif}[1,2]$
+Table 10. Training Data Sampling Distributions
+During the evaluating stage, we fix ciris $=1.25$, hosp2_doctor_shift_counts $=[2,2,4,2,4,1]$ and a set of thresholds in $\{1,...,40\}$.
+Threshold MSE $\downarrow$ MAE $\downarrow$
+MC PEMC MC PEMC
+0 10.974 4.142 2.765 1.587
+4 10.749 3.899 2.751 1.500
+8 10.882 4.080 2.786 1.554
+12 10.243 3.609 2.638 1.466
+16 9.777 3.326 2.555 1.422
+20 9.535 3.266 2.497 1.416
+24 9.291 3.181 2.456 1.394
+28 9.236 3.192 2.445 1.403
+32 9.212 3.145 2.434 1.392
+36 9.175 3.120 2.427 1.382
+40 9.174 3.122 2.428 1.376
+Table 11. Error analysis for mortality estimation with varied thresholds.

### v3

@@
 For many complex simulation tasks spanning areas such as healthcare, engineering, and finance, Monte Carlo (MC) methods are invaluable due to their unbiased estimates and precise error quantification.
-Nevertheless, Monte Carlo simulations often become computationally prohibitive, especially for nested, multi-level, or path-dependent evaluations lacking effective variance reduction techniques.
+Nevertheless, MC simulations often become computationally prohibitive, especially for nested, multi-level, or path-dependent evaluations lacking effective variance reduction techniques.
 While machine learning (ML) surrogates appear as natural alternatives, naïve replacements typically introduce unquantifiable biases.
-We address this challenge by introducing Prediction-Enhanced Monte Carlo (PEMC), a framework that leverages modern ML models as learned predictors, using cheap and parallelizable simulation as features, to output unbiased evaluation with reduced variance and runtime.
-PEMC can also be viewed as a “modernized" view of control variates, where we consider the overall computation-cost-aware variance reduction instead of per-replication reduction, while bypassing the closed-form mean function requirement and maintaining the advantageous unbiasedness and uncertainty quantifiability of Monte Carlo.
-We illustrate PEMC’s broad efficacy and versatility through three examples: first, equity derivatives such as variance swaps under stochastic local volatility models; second, interest rate derivatives such as swaption pricing under the Heath-Jarrow-Morton (HJM) interest-rate model.Finally, we showcase PEMC in a socially significant context—ambulance dispatch and hospital load balancing—where accurate mortality rate estimates are key for ethically sensitive decision-making.
-Across these diverse scenarios, PEMC consistently reduces variance while preserving unbiasedness, highlighting its potential as a powerful enhancement to standard Monte Carlo baselines.
-Equal Contribution. 1Morgan Stanley. 2Department of IEOR, Columbia University. The Work was done during the internship at Morgan Stanley in the Machine Learning Research group.
-     Corresponding author: fengpei.li@morganstanley.com
+We address this challenge by introducing Prediction-Enhanced Monte Carlo (PEMC), a framework that leverages modern ML models as learned predictors, using cheap and parallelizable simulations as features, to output unbiased evaluations with reduced variance and runtime.
+As a result, PEMC eliminates the closed-form-mean requirement that constrains classical control-variate methods, preserves unbiasedness and explicit confidence intervals of MC, and achieves scheme-wide variance reduction.
+Our theoretical analysis quantifies the optimal allocation between expensive evaluations and large batches of cheap, parallelizable feature draws.
+Across three representative applications—variance-swap pricing under stochastic-local-volatility, swaption pricing under Heath–Jarrow–Morton models, and ambulance diversion policy evaluation—we show that PEMC reduces root-mean-squared error by 30–55% relative to standard MC at similar computational cost.
+Equal Contribution. 1Morgan Stanley. 2Department of IEOR, Columbia University. The work was done during empolyment at Morgan Stanley.
+    Corresponding author: fl2412@columbia.edu
 […]
@@
 Monte Carlo (MC) is a fundamental tool for evaluating complex stochastic models where closed-form solutions are unavailable.
-Its main strengths—unbiasedness and rigorous uncertainty quantification—make it essential in high-stakes, high-complexity applications like healthcare resource allocation and financial risk management, where fairness or regulatory compliance are crucial.
+Its main strengths—unbiasedness and rigorous uncertainty quantification through variance estimation and confidence intervals—make it essential in high-stakes, high-complexity applications such as healthcare resource allocation and financial risk management, where fairness or regulatory compliance are crucial.
 For example, transparent and bias-aware methods in healthcare are vital to ensure equitable outcomes and prevent disparities in care access [113, 77, 42].
-Similarly, Monte Carlo estimation is widely used in financial risk management to implement hedging strategies, rebalance portfolios, and establish reliable error bounds, as even small biases in financial simulations can accumulate over time and lead to significant errors [30].
+Similarly, financial risk assessments rely on MC estimation to support hedging, re-balancing decisions, and valid error bounds, as minor biases in financial simulations can accumulate [30].
 Despite these advantages, classical MC faces significant challenges in nested, multilevel, or path-dependent simulations that are difficult to parallelize.
-Complex stochastic dynamics can make each sample path extremely costly to generate—sometimes taking hours for one sample (e.g., credit valuation adjustment [62])—and effective variance reduction methods, such as control variates, are often unavailable [9].
+Complex stochastic dynamics can make each sample path extremely costly to generate—sometimes taking hours for one sample (e.g., to estimate credit valuation adjustments in over-the-counter contracts, see [63])—and effective variance reduction methods are often unavailable [9].
 Consequently, MC’s inherent $\mathcal{O}(1/\sqrt{n})$ convergence rate often prohibits real-time decision-making with adequate precision.
-This is also problematic in exotic option pricing, where sample paths are generated sequentially under complex dynamics, making parallelization unavailable [60].
-The issue is further exacerbated when many evaluation tasks need to be repeatedly conducted in a timely manner.
-In application domains such as option pricing in computational finance, such a computational barrier has been pointed out prominently, e.g., [73, Chapter 21] or [30, Chapter 4].
+This is the case, for example, in the context of pricing of exotic options, where sample paths need to be generated sequentially under complex dynamics, making parallelization not easily achievable [61].
+These computational challenges are critical in financial derivative pricing, as highlighted in the literature, for example, in [74, Chapter 21] and [30, Chapter 4].
+Similar computational bottlenecks arise in healthcare simulation, where modeling complex patient flows and hospital operations can require extensive computation time but transparent and bias-aware methods are crucial to ensure equitable outcomes [113, 77, 42].
 Conversely, machine learning (ML) based predictive modeling has emerged as a fast alternative for approximating MC evaluations, including pricing financial derivatives [75, 53, 15].
 Neural networks and other ML techniques can learn complex mappings and, once trained, can generate predictions at negligible marginal cost - what practitioners often call “fast evaluation" [73].
-However, the black-box nature and the lack of inherent statistical control lead to unquantifiable biases, rendering them unsuitable for tasks requiring rigorous error quantification, e.g., high-stakes applications in risk management [74, 32].
+However, their black-box nature and lack of inherent statistical control hinder the quantification of errors, rendering them unsuitable for tasks requiring rigorous reliability assessments, such as high-stakes applications in risk management [75, 32].
 Thus, MC and ML represent opposite ends of the spectrum: MC is reliable and uncertainty quantifiable but intrinsically slow, while ML is computationally fast but falls short in statistical guarantees.
 Given this trade-off between the speed of ML and the reliability of MC, a natural question arises: can we design a method that combines both advantages while avoiding their respective pitfalls?
-In this paper, we answer affirmatively by introducing the Prediction-Enhanced Monte Carlo (PEMC) framework.
+In this paper, we provide a positive answer by introducing the Prediction-Enhanced Monte Carlo (PEMC) framework.
 PEMC leverages predictive ML modeling to replace a large portion of costly simulations with cheap/parallelizable samples from ML models, reducing variance on a scheme-wide level.
 Moreover, despite relying on black-box predictive models, PEMC maintains unbiasedness, balancing computational savings with rigorous statistical guarantees.
-1.1. Contribution
-The PEMC innovates in two aspects.
+1.1. Our Contributions
+PEMC innovates in two aspects.
 First, PEMC integrates ML predictive models into any MC baseline while preserving unbiasedness and error quantification of MC.
-Second, we show how the predictive modeling of ML combined with cheap, parallelizable samples will result in a scheme-wide reduction of variance and costly samples, enhancing real-time efficiency for complex MC simulations.
-PEMC method can also be viewed as a modernized variant of the control variate (CV) technique, a classical approach that reduces Monte Carlo variance via auxiliary outputs whose known mean offsets random fluctuations [60, 9].
-Unfortunately, the classical CV requirement of a closed-form mean of the a correlated auxiliary output easily becomes prohibitive in many modern applications or complicated stochastic systems.
-In such cases, there simply is no readily available analytic expression for the auxiliary variable’s expectation, effectively barring the use of a CV.
-The PEMC framework bridges the above gap and substantially enlarge the scope of CV to many ML candidates.
+Specifically, we show how the predictive modeling of ML combined with cheap, parallelizable samples will result in a reduction of variance and costly samples, enhancing real-time efficiency for complex MC simulations.
+Second, PEMC can also be viewed as a modernized variant of the control variate (CV) technique, a classical approach that reduces MC variance via auxiliary outputs whose known mean offsets random fluctuations [61, 9, 98].
+Unfortunately, the classical CV approach, which requires a closed-form expression for the mean of a correlated auxiliary output, often becomes unavailable in many modern applications or complex stochastic systems, effectively barring the use of CV.
+To explain concretely the second point above, our PEMC framework bridges ML with MC by substantially enlarging the scope of CV to many ML candidates.
 The key idea is the following: In the traditional CV framework, the known mean of CV is used to achieve a per-replication variance reduction while maintaining unbiasedness, i.e., the variance is reduced even if we run just one simulation replication.
-This is a powerful ability, but as mentioned above it is achieved at the expense of severely narrowing the applicability of CV by requiring closed-form mean function.
+While this is a powerful feature, it comes at the cost of significantly limiting CV’s applicability, due to the requirement of a closed-form expression for the mean.
 However, if the CV itself is computationally cheap in the sense of being estimatable with fast simulation, then we can afford to run plenty of additional MC separately from the original MC to estimate its mean.
-This would preserve unbiasedness and overall estimation variance relative to the total computation cost, we could achieve a scheme-wide efficiency improvement.
-That is, in the PEMC framework, we consider scheme-wide instead of per-replication variance reduction, through which we relax the crucial requirement of known CV mean, and this in turn allows us to substantially expand the CV methodology to cover modern ML models.
+This would preserve unbiasedness and overall estimation variance relative to the total computation cost, as long as we consider efficiency improvement at the scheme-wide level.
+That is, in the PEMC framework, we consider total cost-aware instead of per-replication variance reduction, through which we relax the crucial requirement of known CV mean, and this in turn allows us to substantially expand the CV methodology to cover modern ML models.
 The alignment of our framework with the use of ML predictions prompts us to regard PEMC as a modernized view of CV.
-We demonstrate PEMC’s efficacy across multiple MC applications.
-First, we consider two exotic financial derivative-pricing scenarios: (i) variance swap pricing under a stochastic local volatility model, a fundamental instrument for volatility trading due to its payout on realized variance; and (ii) swaption pricing within the Heath-Jarrow-Morton (HJM) interest rate framework, a sophisticated yet widely used derivative by financial institutions for hedging and speculation.
-Both instruments are highly liquid, exchange-traded products but pose significant computational challenges due to the infeasibility of traditional variance reduction techniques, as they lack analytic or easily computable expected values.
-Additionally, we illustrate PEMC’s broad applicability through an ambulance dispatch and hospital load-balancing simulation, where unbiased mortality rate estimates critically inform ethical and efficient resource allocation.
-These examples collectively highlight PEMC’s capability to leverage machine learning predictions to effectively manage computational complexity while maintaining unbiased estimation and reliable error quantification.
 1.2. Literature Review
-The literature on enhancing Monte Carlo (MC) simulations with machine learning (ML) approximators is extensive, yet many existing ML-based methods for control variates (CV) and variance reduction impose restrictive conditions that limit their practical applicability.
+The literature on enhancing MC simulations with ML approximators is extensive, yet many existing ML-based methods for CV and variance reduction impose restrictive conditions that appear to limit their practical applicability.
 Approaches such as reproducing Stein kernels [102, 89], Neural CV derived from Stein’s identity [132, 83], regularized least squares for CV construction [109, 124, 86], adaptive CV schemes [67, 68, 80], and $\mathscr{L}^{2}$ function approximation frameworks [94] have demonstrated theoretical promise with potentially rapid convergence rates.
-Recent theoretical work by [18] provides important insights into when regression-adjusted control variates can achieve optimal performance, showing that their effectiveness depends critically on function smoothness.
-However, these methods typically require knowledge of the control variate mean or impose specific structural assumptions that can be difficult to check.
-As a result, while these existing methods offer strong theoretical foundations, their applicability in complex financial modeling scenarios can be sometimes limited.
-Multilevel Monte Carlo (MLMC) [58], another prominent variance reduction approach, has shown effectiveness through coupling simulations at different discretization levels but is often challenging to implement, requiring careful design for stable coupling [107, 19, 112, 57, 115].
-Our Prediction-Enhanced Monte Carlo (PEMC) framework addresses these limitations comprehensively by eliminating the necessity of known CV means, enabling the direct use of modern ML predictors as control variates, while rigorously preserving unbiasedness.
-In a different line of literature, the work of Emsermann and Simon [49] on quasi control variates and the work of [102] on control variates using Estimated Means (CVEMs) share notable similarities with our approach, as it relies on an estimated mean (rather than a priori known, and thus not a traditional control variate) of a quantity that is expected to exhibit a strong correlation with the random variable of interest.
-However, our methodology distinguishes itself in several critical ways.
-First, our primary motivation centers on integrating modern ML capabilities to train a suitable control variate, rather than relying on a fixed quantity observed to be correlated with the target (which may not always be obvious or even available, especially for our options pricing examples).
-In other words, we can create an effective control variate from an easily estimable quantity, even if its naive form does not exhibit strong correlation with the target.
-This approach circumvents the challenges of identifying a closed-form mean and strictly relying on high correlation-based pairing.
-Second, instead of adopting an adaptive algorithm as they do, we focus on a pre-training approach, demonstrating that our performance advantages would not emerge under their adaptive scheme.
-Third, our analysis is not anchored exclusively in asymptotic arguments.
-We provide a formulation that can be interpreted in both finite-sample and asymptotic regimes, offering more direct and flexible theoretical guarantees.
-In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively even when auxiliary means are unknown.
-Unlike previously discussed methods, PEMC aligns closely with the recently proposed Prediction-Powered Inference (PPI) approach [6], particularly in its hybrid methodology that integrates classical, unbiased statistical principles with modern machine learning (ML) techniques.
-This hybrid strategy is becoming increasingly important in evaluating generative models such as large language models (LLMs), where rapid, interpretable, and unbiased metrics are essential [135, 22, 51].
-PPI-type approaches combine sparse human annotations with abundant ML-generated labels, facilitating accurate model performance evaluation even in low-label scenarios.
-Additionally, these methodologies intersect with a rich literature in causal inference, specifically doubly robust estimators [10, 55, 113] and orthogonal statistical learning [36, 54].
+Recent theoretical work by [18] provides important insights into when regression-adjusted CVs can achieve optimal performance, showing that their effectiveness depends critically on function smoothness.
+However, these methods typically require knowledge of the CV mean, in turn restricting the classes of usable CV, or impose specific structural assumptions that can be difficult to check.
+As a result, while these existing methods offer strong theoretical foundations, their applicability in complex modeling scenarios can be undermined.
+In a different line of literature, the works of [50] on quasi CVs and [104] on CVs using Estimated Means (CVEMs) is related to our approach, because they rely on an estimated mean (rather than a priori known, and thus not a traditional CV) of a quantity that is expected to exhibit a strong correlation with the random variable of interest.
+There are noticeable differences between our methodology and theirs.
+To begin with, our primary objective is to integrate modern ML capabilities into MC, so that our estimates can retain the error-quantifiability of the latter approach while leveraging the strengths of ML.
+To this end, a core component of our methodology regards what is needed, and how to train a suitable CV, rather than relying on a prefixed CV candidate, which may not always be obvious or even available.
+Moreover, our investigation includes the use of a pre-training approach in place of the adaptive algorithms employed in previous works, and provides a framework applicable to both finite-sample and asymptotic regimes, unlike the exclusively asymptotic focus of prior studies.
+Within the literature of MC variance reduction, PEMC also relates to Multilevel Monte Carlo (MLMC) methods, originally introduced and popularized by [59], and also the idea of multi-fidelity modeling popular in scientific computation [106].
+MLMC reduces variance by leveraging a hierarchy of discretizations (e.g., coarser and finer time steps in SDE simulations) and coupling them to achieve computational efficiency over naive Monte Carlo approaches [58].
+The key idea is to couple simulations across resolution levels so that their differences exhibit reduced variance—or even enable unbiased estimators [115, 19, 118].
+Along a similar vein, in multi-fidelity modeling, computational simulation models of different levels of resolution, which in turn bring in different accuracies and computational demands, are jointly utilized and concatenated, with input model parameters treated as random variables that allow coupling between the model levels [106].
+Note that, despite the strong theoretical foundations, MLMC can be difficult to implement in practice, since constructing effective couplings, ensuring numerical stability, and tuning parameters to realize the theoretical complexity benefits are nontrivial tasks [110].
+PEMC adopts the coupling paradigm of MLMC and multi-fidelity modeling but innovates by using ML predictors as CV.
+Outside the MC literature, our PEMC framework aligns most closely with the recently proposed Prediction-Powered Inference (PPI) approach [6].
+PPI is a statistical framework that enables researchers to construct valid confidence intervals and p-values when combining a small dataset with gold-standard labels and a large dataset with ML predictions.
+The key challenge PPI addresses is how to leverage abundant but potentially biased ML predictions alongside scarce but trustworthy ground-truth data for statistical inference.
+Such PPI-type hybrid strategy is becoming increasingly important in evaluating generative models such as large language models (LLMs) [138, 22, 52], where sparse human annotations is combined with abundant ML-generated labels to facilitate accurate model performance evaluation even in low-label scenarios.
+Additionally, the PPI methodologies also intersect with a rich literature in causal inference, specifically in doubly robust estimators [10, 56, 116] and orthogonal statistical learning [36, 55].
 Doubly robust methods construct additional robustness layers by pairing sophisticated predictive models with auxiliary estimators, ensuring that minor errors in one model component do not significantly bias the overall estimation.
-This philosophy closely mirrors the logic of control variates, and by extension, the core principles of our PEMC framework.
-Traditional variance reduction and numerical approximation techniques also have a long and established history in computational finance and stochastic modeling.
-One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by [58].
-MLMC exploits a hierarchy of discretizations (e.g., finer and coarser time steps in an SDE simulation) to reduce variance at a fraction of the cost of a naive Monte Carlo approach.
-Its basic premise involves coupling simulations at different resolution levels so that differences between levels have lower variance, ultimately leading to significant computational savings.
-Despite its solid theoretical foundations, MLMC can be challenging to implement in practice: constructing suitable couplings between the fine and coarse approximations, ensuring stability, and tuning parameters to achieve the promised complexity gains are all nontrivial tasks.
-Nevertheless, the class of MLMC methods and their adoption of a coupling perspective provide extremely valuable insights into how PEMC manages the relationship between the target and the predictor [107, 19, 112, 57, 115].
-In computational finance, the emergence of machine learning (ML) has also sparked significant research efforts aimed at enhancing the speed and accuracy of financial computations, including option pricing.
+This philosophy closely mirrors the logic of CV, and by extension, the core principles of our PEMC framework.
+While PPI replaces expensive gold-standard labels with cheaper ML predictions and auto-evaluation, PEMC focuses on replacing expensive, path-dependent and non-parallelizable MC samples with cheap, efficiently simulated, and highly parallelizable MC samples.
+Finally, our work is also related to a branch of literature in computational finance, which has focused on harnessing ML to enhance the speed and accuracy of derivatives pricing.
 By using techniques such as neural network surrogates, these methods can approximate pricing functions or the Greeks at a fraction of the computational cost of traditional numerical schemes [75, 32, 122, 117, 85].
-Early work by hutchinson1994nonparametric [74], for instance, used neural networks to relax stringent modeling assumptions, while more recent approaches [32, 119, 114, 84] demonstrate that deep architectures can effectively handle high-dimensional and path-dependent pricing problems.
+Early work by hutchinson1994nonparametric [75], for instance, use neural networks to relax stringent modeling assumptions, while more recent approaches [32, 122, 117, 85] demonstrate that deep architectures can effectively handle high-dimensional and path-dependent pricing problems.
 However, these purely data-driven approximations typically lack the rigorous error controls afforded by more classical methods.
-As a result, biases introduced during training and potential overfitting can lead to systematic mispricings, particularly under market regimes not well-represented in the training data.
-Such issues underscore the need for frameworks that combine the computational advantages of ML with robust theoretical guarantees to ensure reliable and unbiased pricing outcomes.
-On the other hand, within option pricing, even though both MC and PDE-based methods are widely used, it is known that PDE methods generally cannot be used to price options (i.e., payoff depends on the price path).
+In the context of option pricing, even though both MC and partial differential equation (PDE)-based methods are widely used, it is known that PDE methods generally cannot be used to price options (i.e., payoff depends on the price path).
 While special cases allow for PDE formulations through state-space augmentation or boundary condition adjustments, these remain as exceptions rather than the rule [119].
 The rest of the paper is organized as follows.
-In Section 3, we provide a comprehensive background on PEMC.
-In Section 4, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical illustrations.
-In Section 5, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the HJM framework.
-We conclude the paper in Section 6 with discussion on broader implications and potential directions for future research.
-2. Prediction-Enhanced Monte Carlo
+In Section 2, we motivate and introduce PEMC.
+In Section 3, we provide more details on our framework via a running example of Asian option pricing.
+In Section 4, we present theoretical results and their practical implications.
+In Section 5, we apply PEMC to production-grade exotic option pricing, including variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the HJM framework, as well as an ambulance diversion problem.
+We conclude the paper in Section 6 with discussions on broader implications and potential directions for future research.
+2. The Prediction-Enhanced Monte Carlo Framework
 We provide an overview and motivational explanation of our PEMC framework. Suppose we want to produce an unbiased evaluation of the quantity
@@
 for arbitrary, potentially high-dimensional parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{d_{\theta}}$, where $\Theta$ is the parameter space.
-Here, $\boldsymbol{Y}\in\mathbb{R}^{d_{y}}$ represents a high-dimensional random vector governed by $\mathbb{E}_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}}$ is the evaluation function $\mathbb{R}^{d_{y}}\to\mathbb{R}$.
+Here, $\boldsymbol{Y}\in\mathbb{R}^{d_{y}}$ represents a high-dimensional random vector whose probability distribution is parameterized by $\mathbb{E}_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}}$ is the evaluation function $\mathbb{R}^{d_{y}}\to\mathbb{R}$.
 In general, both $f_{\boldsymbol{\theta}}$ and $\mathbb{E}_{\boldsymbol{\theta}}$ can depend on $\boldsymbol{\theta}$.
-The goal is to produce an unbiased estimate of $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, while the primary bottleneck stems from costly and time-consuming simulation of $\boldsymbol{Y}$.
-When $\boldsymbol{Y}$ can be simulated and $f$ evaluated, then standard MC amounts to generating many $f(\boldsymbol{Y}_{i})$, $i=1,\ldots,n$ and outputting their average.
-While this approach is demanding when $f$ and $\boldsymbol{Y}$ are expensive, it is unbiased and its statistical error can be quantified straightforwardly via standard confidence intervals.
-To reduce computation, one can approximate $\mu(\boldsymbol{\theta})$ with an ML regression model, say $\hat{\mu}(\boldsymbol{\theta})$, which can be evaluated more cheaply than simulating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$’s.
+The goal is to produce an unbiased estimate of $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and the primary bottleneck stems from costly simulation of $\boldsymbol{Y}$.
+2.1. Existing Challenges of MC, ML, and CV
+When $\boldsymbol{Y}$ can be simulated and $f_{\boldsymbol{\theta}}$ evaluated, then standard MC amounts to generating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$, $i=1,\ldots,n$ and outputting their average.
+This approach is unbiased and its statistical error can be quantified straightforwardly via standard confidence intervals.
+However, it is demanding when evaluating $f_{\boldsymbol{\theta}}$ and simulating $\boldsymbol{Y}$ is expensive.
+To reduce computation, one can approximate $\mu(\boldsymbol{\theta})$ with a ML regression model, say $\hat{\mu}(\boldsymbol{\theta})$, which can be evaluated more cheaply than simulating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$’s.
 This natural approach, however, would face challenges at multiple levels when reliability and error quantification is important: From a theoretical standpoint, ML generalization bounds typically assess training errors only within the model class, i.e., $\hat{\mu}(\boldsymbol{\theta})-\mu^{*}(\boldsymbol{\theta})$ where $\mu^{*}(\cdot)$ is the best in-class model that can be different from $\mu(\cdot)$.
@@
 None of these advantages are available for ML despite its computational edge.
-An idea to leverage the respective advantages of MC and ML is to use control variate (CV), a classical variance reduction method.
+An idea to leverage the respective advantages of MC and ML is to use CV which, as mentioned in Section 1, is a classical variance reduction method.
 CV aims to reduce the error of each MC run, consequently the required replication size, by utilizing information from “auxiliary” simulation outputs.
-Specifically, in generating $f(\boldsymbol{Y}_{i})$, suppose we can also generate auxiliary output $\boldsymbol{X}_{i}$ and evaluate $g(\boldsymbol{X}_{i})$ for some function $g$.
-Then we can use $f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i})+\mathbb{E}[g(\boldsymbol{X})]$ as an unbiased replicate for $\mathbb{E}[f(\boldsymbol{Y})]$.
-In particular, if $f(\boldsymbol{Y})$ and $g(\boldsymbol{X})$ are highly correlated, this replicate will have a smaller variance than $f(\boldsymbol{Y}_{i})$.
-Naturally, one can think of using ML to train a good predictor $g(\boldsymbol{X})$ for $f(\boldsymbol{Y})$, which would be highly correlated with $f(\boldsymbol{Y})$ and serve as an effective CV.
-However, taking aside the question of how to construct $\boldsymbol{X}$ and $g$, the CV method requires knowledge of $\mathbb{E}[g(\boldsymbol{X})]$, the closed-from mean on the auxiliary output.
-This knowledge is the key convertible into variance reduction.
+Specifically, in generating $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$, suppose we can also generate auxiliary output $\boldsymbol{X}_{i}$ and evaluate $g(\boldsymbol{\theta},\boldsymbol{X}_{i})$ for some function $g$.
+Then we can use $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})+\mathbb{E}[g(\boldsymbol{\theta},\boldsymbol{X})]$ as an unbiased replicate for $\mathbb{E}_{\boldsymbol{\theta}}[f_{\boldsymbol{\theta}}(\boldsymbol{Y})]$.
+In particular, if $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$ and $g(\boldsymbol{\theta},\boldsymbol{X})$ are highly correlated, this replicate will have a smaller variance than $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$.
+Naturally, one can think of using ML to train a good predictor $g(\boldsymbol{\theta},\boldsymbol{X})$ for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$, which would be highly correlated with $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$ and serve as an effective CV.
+However, taking aside the question of how to construct $\boldsymbol{X}$ and $g$, the CV method requires knowledge of $\mathbb{E}_{\boldsymbol{\theta}}[g(\boldsymbol{\theta},\boldsymbol{X})]$, the closed-from mean on the auxiliary output.
+This knowledge is a key requirement for CV to achieve variance reduction.
 Yet, it is seldom available and has apparently severely limited the historical applicability of CV [62] and arguably even the more recent ML-inspired approaches [132, 124, 89].
-We introduce Prediction-Enhanced Monte Carlo (PEMC), a framework that fulfills these objectives.
-PEMC leverages the classical control variate idea but generalizes its variance-reduction notion to a scheme-wide, modernized view, allowing for effective integration of MC with modern ML.
+2.2. Our Remedy and Requirements
+Our PEMC aims to leverage the CV idea, with a twist to circumvent the need on the closed-form mean knowledge in order to allow for effective integration of MC with modern ML.
+This requires crucially a scheme-wide cost-aware view on variance reduction.
 Concretely, PEMC aims to estimate (1) by
 $PEMC(\boldsymbol{\theta}):=\frac{1}{n}\sum_{i=1}^{n}\big{(}f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})\big{)}+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\theta},\boldsymbol{\tilde{X}}_{j}).$ (2)
-Here, $\boldsymbol{X}$ is an auxiliary variable, coupled with $\boldsymbol{Y}$, that serves as features for predicting $f(\boldsymbol{Y})$ and is also chosen to have a marginal distribution that is both efficiently simulatable and highly parallelizable, allowing large number $N$ independent samples of $\tilde{\boldsymbol{X}}$ with the same marginal as $\boldsymbol{X}$ to be generated at low cost.
-The function $g$ is a prediction model trained via squared-error loss using MC sample under $(\boldsymbol{X},\boldsymbol{Y})\sim\mathbb{P}_{\boldsymbol{\theta}}$ for various $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, which aims to approximate the function $g\bigl{(}\boldsymbol{\theta},\boldsymbol{X}\bigr{)}\approx\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}f_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{Y}\bigr{)}\big{|}\boldsymbol{\theta},\boldsymbol{X}\bigr{]}$.
-As we will see, the selection of an appropriate $\boldsymbol{X}$ is a pivotal component of PEMC, analogous to feature selection in common ML tasks.
-The key lies in selecting $\boldsymbol{X}$ to be both a feature for predicting $f(\boldsymbol{Y})$ and an efficiently simulatable, parallelizable random variable.
+Here, $\boldsymbol{X}$ is an auxiliary variable that is dependent on $\boldsymbol{Y}$.
+Suppose, for now, that we are given the $g$ function and the $\boldsymbol{X}$ variable.
+In equation (2), each pair of $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$ is coupled and simulated together.
+On the other hand, each $\tilde{\boldsymbol{X}}_{j}$ is simulated independently from all $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$’s, and has the same marginal distribution as $\boldsymbol{X}$.
+In this case, it is straightforward to see that (2) is an unbiased estimator of (1).
+On the other hand, in terms of cost efficiency, the total simulation cost consists of
+$[\text{cost of generating each\ }f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})]\times n+[\text{cost of generating each\ }g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j})]\times N$ (3)
+while the total variance is
+$\frac{Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))}{n}+\frac{Var(g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j}))}{N}$ (4)
+Suppose that $g$ and $\boldsymbol{X}$ are designed such that the cost of generating each $g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j})$ is relatively negligible compared to each original MC target replicate $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$.
+Then, the total cost (3) will become approximately $[\text{cost of evaluating each\ }f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})]\times n$, while the total variance (4) will become approximately $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))/n$.
+In this case, the cost-variance structure is the same as the conventional CV method: If $g(\boldsymbol{\theta},\boldsymbol{X})$ is a good predictor for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$, then $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))$ is smaller than $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i}))$ and thus we attain a variance reduction to the overall estimator.
+Like in the conventional CV method, to achieve a total reduction from $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i}))/n$ to $Var(f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i}))/n$, we generate $n$ samples of $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})$, i.e., variance reduction is achieved per replication.
+While the above depicts the idealized situation where we assume away the cost of generating $g(\boldsymbol{\theta},\tilde{\boldsymbol{X}}_{j})$, it provides guidance on what we need practically to elicit a cost-variance structure in (2), as revealed by (3) and (4), that is more favorable than naive MC.
+First, we need $g(\boldsymbol{\theta},\boldsymbol{X})$ to serve as a good predictor for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$, but also that $\boldsymbol{X}$ has a marginal distribution that is both efficiently simulatable and highly parallelizable, allowing a large number $N$ of independent samples of $\tilde{\boldsymbol{X}}_{j}$ with the same marginal as $\boldsymbol{X}$ to be generated at low cost.
+To this end, the selection of an appropriate $\boldsymbol{X}$ is a pivotal component of PEMC, analogous to feature selection in common ML tasks but with the additional need of being an efficiently simulatable, parallelizable random variable.
 Finding a good $\boldsymbol{X}$ often requires domain knowledge.
 In particular, given the whole path $\boldsymbol{Y}$, one can choose $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ to be some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ (e.g., ingredients for constructing $\boldsymbol{Y}$).
-For instance, for SDEs, one can define $\boldsymbol{X}$ as the sum of driving Brownian increments that generate $\boldsymbol{Y}$.
+For instance, for stochastic diferential equations (SDEs), one can define $\boldsymbol{X}$ as the sum of driving Brownian increments that generate $\boldsymbol{Y}$.
 This choice could capture a substantial portion of the SDE’s variance, making $\boldsymbol{X}$ a good feature for predicting $f(\boldsymbol{Y})$, while its Gaussianity allows it to be generated cheaply and in parallel.
-We explore this in detail in subsequent sections.
-Many implementation details remain unspecified in the basic description above.
-For instance, how exactly to train g, how to select the sample sizes $n$ and $N$, and how to adapt PEMC to a conditional setting where $\theta\mapsto\boldsymbol{X}(\theta)$, $\boldsymbol{Y}(\theta)$, and $\mathbb{E}_{\theta}$ all depend on some parameter $\theta$.
-The next sections will address these questions in depth, providing both theoretical justification and practical guidelines for deploying PEMC in diverse scenarios.
-3. Prediction-Enhanced Monte Carlo: Framework via Example
-In this section, we employ Asian option pricing as a running example on how to apply PEMC.
+Second, we require a function $g$ to convert $\boldsymbol{\theta}$ and $\boldsymbol{X}$ into a good predictor for $f_{\boldsymbol{\theta}}(\boldsymbol{Y})$.
+In this work, $g$ is pre-trained using a squared-error loss, with MC sample drawn under $(\boldsymbol{X},\boldsymbol{Y})\sim\mathbb{P}_{\boldsymbol{\theta}}$ for various $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, which aims to obtain a function $g\bigl{(}\boldsymbol{\theta},\boldsymbol{X}\bigr{)}\approx\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}f_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{Y}\bigr{)}\big{|}\boldsymbol{\theta},\boldsymbol{X}\bigr{]}$.
+Note that this pre-training can be resource-consuming and takes significant amount of efforts.
+However, once this pre-training is done, the function $g$ and our approach can be used to improve the efficiency in evaluating $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}$.
+For example, in the context of option pricing, the pre-trained $g$ can be stored as part of a model library tailored to a specific class of exotic options, enabling PEMC to be applied directly whenever new parameter configurations arise.
+3. An Illustrative Case Study: Pricing Asian Options with PEMC
+In this section, we employ Asian option pricing as a running example to explain the detailed steps and their implications in PEMC.
 Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period.
-As mentioned in [23] and [44], closed-form solutions for the pricing of arithmetic average Asian options are generally not available.
+As mentioned in [23] and [45], closed-form solutions for pricing arithmetic average Asian options are generally not available.
+However, when the asset price dynamic follows a geometric Brownian motion (GBM), there exists an excellent CV candidate, using the geometric average instead of the arithmetic average, whose mean is precisely known via the Black-Scholes formula.
+This latter case often serves as the “textbook" example to demonstrate the power of CV [61].
+As an important message, our case study here illustrates how PEMC can achieve a comparable level of improvements as the “textbook" CV, for evaluating arithmetic average Asian option price when the dynamics does not follow GBM.
+In other words, PEMC is capable to extend the scope of classical CV to potentially a much wider range of problems.
 3.1. Problem Setup
@@
 Option $\text{Option Price}(\boldsymbol{\theta})=\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}[f_{\text{payoff}(\boldsymbol{\theta})}(\boldsymbol{Y})],$ (5)
+where the risk neural measure is a probability measure where all financial securities are valued as if investors are indifferent to risk, and thus all assets are assumed to earn the risk-free rate [40].
+[^1] Under this measure, prices of financial securities are the discounted expected values of their future payoffs.
 For Asian option, we can specify the framework as:
@@
     Simulation Parameters $\boldsymbol{\theta}_{\text{simulation}}$: These parameters also specify $\mathbb{E}_{\text{risk neutral measure}}$.
-The distinction between $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are calibrated externally, but rather serve as hyperparameters of the Monte Carlo simulation itself (either customized or directly observed).
+The distinction between $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{model}}$ is somewhat artificial, primarily stemming from the convention that not all simulation parameters are calibrated externally, but rather serve as hyperparameters of the MC simulation itself (either customized or directly observed).
 For example, when one simulates the Heston model above using Euler discretization scheme, [^2] As mentioned in [45], Euler scheme was the primary method for simulation of Heston model until the exact simulation of [31] is developed.
@@
     Payoff Function Parameters $\boldsymbol{\theta_{\text{payoff}}}$: This parameter $\boldsymbol{\theta_{\text{payoff}}}$ parametrizes the payoff function $f_{\boldsymbol{\theta}_{\text{payoff}}}$.
-In options pricing, this specifies the details of the contract.
+In option pricing, this specifies the details of the contract.
 For example, given the full price path $\boldsymbol{Y}:=\{S_{t}\}_{t}$, the payoff function for an arithmetic Asian options takes the form
-    $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
+    $P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$ (7)
     where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
 Notice we have simply omitted the discount factor $e^{-rT}$.
-Note that the discrete sampling defied by the payoff does not necessarily determine the Euler discretization in time and in fact the latter is usually much more frequent for accuracy.
-Remark Having specified the distinct roles of $\boldsymbol{\theta}_{\text{model}}$, $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\boldsymbol{\theta}$ when no ambiguity arises.
-3.2. Prediction-Enhanced Monte Carlo: Motivation
-As we have seen, Asian options are path-dependent derivatives whose payoffs depend on the average prices of the underlying asset.
-Generally, closed-form solutions for these options are unavailable, prompting numerous studies to propose various numerical methods for their pricing.
-In particular, the arithmetic Asian option under the geometric Brownian motion (GBM) model is one of the rare cases where a closed-form geometric Asian option price exists and can serve as a control variate.
-Beyond simple models like GBM, closed-form CV solutions are unavailable [44].
-To address these challenges, the Prediction-Enhanced Monte Carlo (PEMC) framework aims to reduce the number of required Monte Carlo samples by generating a machine learning-based control variate that is both highly parallelizable and efficiently evaluable.
-Crucially, our framework preserves the unbiased nature of the Monte Carlo estimator, a property that [60] highlights as “fundamental to the wide applicability of Monte Carlo methods in finance,” achieving a remarkable elimination of any potential bias introduced by any black-box prediction model.
-In PEMC, we do not specify the training algorithm or model architecture for this prediction model, as our framework is compatible with any black-box model.
-However, we want to find some input feature $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ that is some typically low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ with:
+Note that the discrete sampling defined by the payoff does not necessarily determine the Euler discretization in time and in fact the latter is usually much more frequent for accuracy.
+Having specified the distinct roles of $\boldsymbol{\theta}_{\text{model}}$, $\boldsymbol{\theta}_{\text{simulation}}$ and $\boldsymbol{\theta}_{\text{payoff}}$, we will henceforth refer to them collectively as $\boldsymbol{\theta}$ when no ambiguity arises.
+3.2. Feature Engineering to Design $\Large{X}$
+We now follow the general ideas laid out in Section 2.2, starting with the creation of the variable $\boldsymbol{X}$ to couple with $\boldsymbol{Y}$.
+As described in Section 2.2, we want to find some input feature $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ that is some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ with:
   • Property 1: $\boldsymbol{X}$ can be regarded as a reasonable predictor for $f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y})$.
   • Property 2: In contrast to that of $\boldsymbol{Y}$, the marginal distribution of $\boldsymbol{X}$ allows for efficient and parallelizable simulation.
-The selection of an appropriate $\boldsymbol{X}$ is a pivotal component of our framework, analogous to feature selection in machine learning tasks.
-This process necessitates domain expertise.
-For now, we proceed under the assumption that a suitable $\boldsymbol{X}$ has been identified.
-3.3. Prediction-Enhanced Monte Carlo: Prediction Model
+To this end, we take $\boldsymbol{X}$ as the sum of two correlated Brownian increments driving (1), motivated by two key properties:
+  • Predictability: The sum of Brownian increments explains away a portion of the variance of the SDE.
+  • Marginal simulation: The sum of two Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
+More precisely, when sampling $\boldsymbol{Y}$, we collect $\boldsymbol{X}$ to be a two-dimensional vector consisting of
+$W_{T}^{S}:=\sum_{j=1}^{T/\Delta t}\Delta W^{S}_{j\Delta t}\quad\text{ and }\quad W_{T}^{\nu}:=\sum_{j=1}^{T/\Delta t}\Delta W^{\nu}_{j\Delta t},$
+simulated during each step of the Euler scheme with discretization scale $\Delta t$.
+The marginal of $\boldsymbol{X}$ is simply a two-dimensional Gaussian: $\frac{\boldsymbol{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
+Using Brownian increments, while natural, is not the only choice.
+As with most ML tasks, feature engineering that leverages domain expertise can yield better results.
+We illustrate with another example, the pricing of a floating strike lookback call option, under the Heston model (1).
+Here, the option payoff is given by:
+$f(\{S_{t}\}_{t})=\max(S_{T}-\min_{0\leq t\leq T}S_{t},0)$
+where $S_{t}$ is the asset price process and $T$ is the option maturity. A good choice of feature could be:
+$\boldsymbol{X}=(W_{T}^{S},\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}).$
+The rationale behind the construction of $\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}$ term is grounded in the domain knowledge of the Heston model.
+Specifically, since $\eta$ represents the mean-reverting level of variance process $\nu_{t}$, we may approximate $\nu_{t}\approx\eta$ (i.e., assume fixed volatility) and simplify the asset price process to a GBM:
+$S_{t}\approx S_{0}e^{(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}\text{ and }\min_{0\leq t\leq T}S_{t}\approx S_{0}e^{\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}.$
+This approximation leads to the expression $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\boldsymbol{X}]_{1}}-S_{0}e^{[\boldsymbol{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$.
+This corresponds directly to Property 1 in the construction of $\boldsymbol{X}$ in 3.3.
+For Property 2, the marginal distribution of $\boldsymbol{X}$ permits efficient and parallelizable simulation.
+In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\boldsymbol{X}]_{1},[\boldsymbol{X}]_{2})$ [79].
+This result makes it straightforward to sample these components directly.
+3.3. Building Prediction Model
 We now focus on the pre-training phase of our prediction model. The data generation and model training process can be outlined through the following key steps:
@@
 of $\boldsymbol{Y}(\boldsymbol{\theta}_{i})$ and its corresponding $\boldsymbol{X}(\boldsymbol{\theta}_{i}):=\phi(\boldsymbol{Y}(\boldsymbol{\theta}_{i}))$.
-It’s important to note that $\boldsymbol{X}$ and $\boldsymbol{Y}$ are coupled in this generation process.
+It is important to note that $\boldsymbol{X}$ and $\boldsymbol{Y}$ are coupled in this generation process.
   3. (3)
@@
   5. (5)
-    The machine learning model $g$ is then trained to minimize the mean squared error (MSE) loss and saved:
+    The ML model $g$ is then trained to minimize the MSE loss and saved:
     $\min_{g}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-g(\text{feature}_{i}))^{2}.$ (8)
@@
 18:end procedure
-Algorithm 1 Prediction Model Training
+Algorithm 1 Prediction Model Training in PEMC
 The essential steps are summarized in Algorithm 1.
@@
 This approach minimizes its impact on real-time pricing applications by effectively trading memory usage for computational speed during execution.
-This pre-training setup presents an ideal scenario for machine learning applications.
+This pre-training setup presents an ideal scenario for ML applications.
 It is important to also note that PEMC does not require $g$ to take specific forms.
@@
 For training NNs, various optimization techniques and architectural enhancements—such as the Adam optimizer [81], batch normalization [76], and skip connections [65]—can be employed to improve performance.
-In fact, recent advancements in machine learning technologies have made the training process remarkably accessible and efficient.
+In fact, recent advancements in ML technologies have made the training process remarkably accessible and efficient.
 The availability of open-source, user-friendly frameworks like PyTorch [105], TensorFlow [1], and JAX [24]—coupled with GPU acceleration and highly optimized C++ backends—has dramatically simplified the process of training models on large datasets.
 These frameworks support recent neural network architectures, such as Convolutional Neural Networks (CNNs) [84], Residual Networks (ResNets) [65], and Transformers [131], as well as various Stochastic Gradient Descent (SGD) optimization algorithms [21].
-In applications, we only utilized well-established and widely recognized neural network structures that are now considered common.
+In applications, we only utilized well-established and widely recognized NN structures that are now considered common.
 These models can typically be implemented using just a few lines of code, as they have been incorporated into existing packages and extensively optimized.
-Finally, regarding the choice loss function, we note that the goal of the prediction model $g$ is to produce an approximation of the underlying regression function:
-$g(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))$
-$\approx$ $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}\bigg{[}f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y}(\boldsymbol{\theta}))\bigg{|}(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X}(\boldsymbol{\theta}))\bigg{]}$
-As we will see, the quality of this approximation will impact the effectiveness of variance reduction, which we shall explore in detail in subsequent sections.
-3.4. Prediction-Enhanced Monte Carlo: Evaluation
-Having completed the training phase of PEMC, we now turn our attention to the evaluation procedure.
-This stage employs the pre-trained model to enhance the efficiency of Monte Carlo simulations, and we will demonstrate how control variates are cleverly constructed to achieve variance reduction.
-The evaluation process proceeds as follows:
+3.4. Evaluation
+Once we finish the pre-training phase, the prediction model is stored and used on-the-fly in the evaluation process as follows:
   1. (1)
-    A specific parameter $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})\in\boldsymbol{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing.
+    A specific parameter $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})\in\boldsymbol{\Theta}$ is given, e.g., calibrated from market data in real time, and needs to be used for pricing.
   2. (2)
     Generate $n$ pairs of $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$.
-Here we have suppressed $\boldsymbol{Y}$’s dependency on $\boldsymbol{\theta}$, as in the evaluation stage only one fixed $\boldsymbol{\theta}$ is considered.
+Here we have suppressed $\boldsymbol{Y}$’s dependency on $\boldsymbol{\theta}$ for convenience.
   3. (3)
     Based on the marginal distribution of $\boldsymbol{X}$, independently generate $N$ additional samples of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$, directly according to its marginal distribution.
-Importantly, these samples are independent of the previous data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$.
+These samples are independent of the previous data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$.
   4. (4)
@@
 6:end procedure
-Algorithm 2 Evaluation
-3.5. PEMC: An Example on Asian Options
-We now return to our example on arithmetic Asian option under the Heston model:
-$P_{A}(\{S_{t}\}_{t})=\max\left(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K,0\right)$
-where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
+Algorithm 2 Evaluation Procedure in PEMC
+3.5. Overall Methodology
+We now summarize our overall PEMC methodology applied to the Asian option under the Heston model, i.e., (7), with $\boldsymbol{\theta_{\text{payoff}}}$ specifying the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$.
 The Heston model is described in (1) which requires $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$.
-For this example, let’s say we use a neural network (NN) as the prediction model.
+Suppose we use an NN as the prediction model.
 Then, the PEMC framework could proceed as follows:
@@
   7. (7)
-    At evaluation, a specific $\boldsymbol{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4).
+    At evaluation, given a specific $\boldsymbol{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4).
 We also generate $N$ i.i.d.
@@
     Output $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(\textit{label}_{i}-NN_{\boldsymbol{\hat{w}}}(\textit{feature}_{i}))+\frac{1}{N}\sum_{j=1}^{N}NN_{\boldsymbol{\hat{w}}}(\tilde{\textit{feature}}_{j})$
-3.6. Implementation Details in PEMC
-Having walked through the Asian option example under the PEMC framework, we discuss several important implementation details.
-These considerations are critical for successfully applying PEMC in practice.
-3.6.1. Parameter Space $\Theta$
-When determining the parameter space $\boldsymbol{\Theta}$, a key consideration is how frequently the model will be updated.
-To ensure the performance of PEMC, it is desirable that the training data encompass all practical scenarios the model is expected to encounter within the designated update period.
-For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$ and the user of PEMC chooses to update the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses or confidence intervals about the bounds within which these parameters are likely to fluctuate over the coming month.
-This process is conceptually similar to the selection of uncertainty sets in robust optimization [16] or distributionally robust optimization [50], where the goal is to include realizations of parameters with high likelihood.
-However, unlike robust optimization approaches, which often favor data-dependent uncertainty sets, our focus here is also guided by the expertise of financial engineers.
-The training parameter space is designed to reflect realistic and practical scenarios derived from domain knowledge, rather than being strictly driven by statistical guarantees.
-In applications, we construct reasonable $\boldsymbol{\Theta}$ for practical models, such as forward curves in HJM modeling [60] or local stochastic volatility 2D grids [56], among others.
-3.6.2. The Design of $\Large{X}$: Feature Engineering
-In the example, we designed the feature $\boldsymbol{X}$ as the sum of two correlated Brownian increments. As discussed in Section 3.3, the choice was motivated by two key properties:
-  • Predictability: The sum of Brownian motion increments explains away a portion of the variance of the SDE.
-  • Marginal simulation: The sum of two Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
-Given these considerations, it is tempting to employ Gaussian sums as a generic recipe for the majority of option pricing problems handled by PEMC, particularly since these problems commonly involve stochastic differential equations driven by Brownian motion.
-Indeed, this is precisely the approach we adopt in our applications, demonstrating its robustness and effectiveness.
-In cases involving long simulation paths, one might consider segmenting the sum into smaller increments, though we observed that a single increment usually suffices to achieve significant variance reduction.
-However, the choice of Brownian increments, while effective, is certainly not the only one.
-As with most ML tasks, feature engineering that leverages domain expertise can yield better results.
-We illustrate with one such example, the pricing of a floating strike lookback call option, under the Heston model(1).
-The option payoff is given by:
-$f(\{S_{t}\}_{t})=\max(S_{T}-\min_{0\leq t\leq T}S_{t},0)$
-where $S_{t}$ is the asset price process and $T$ is the option maturity. A good choice of feature could be:
-$\boldsymbol{X}=(W_{T}^{S},\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}).$
-The rationale behind the construction of $\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W_{t}^{S}$ term is grounded in the domain knowledge of the Heston model.
-Specifically, one who is familiar with the role of $\eta$ as the mean-reverting level of variance process $\nu_{t}$ may approximate $\nu_{t}\approx\eta$ (i.e., assume fixed volatility) and simplify the asset price process to a geometric Brownian motion:
-$S_{t}\approx S_{0}e^{(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}\text{ and }\min_{0\leq t\leq T}S_{t}\approx S_{0}e^{\min_{0\leq t\leq T}(r-\eta/2)t+\sqrt{\eta}W^{S}_{t}}.$
-This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\boldsymbol{X}]_{1}}-S_{0}e^{[\boldsymbol{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$.
-This corresponds directly to Property 1 in the construction of $\boldsymbol{X}$ in 3.3.
-For Property 2, the marginal distribution of $\boldsymbol{X}$ permits efficient and parallelizable simulation.
-In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\boldsymbol{X}]_{1},[\boldsymbol{X}]_{2})$ [78].
-This result makes it straightforward to sample these components directly.
-3.6.3. Eliminating Data Storage Overheads
-In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively.
-By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage.
-This difference stems from two main considerations.
-First, the volume of training data required to achieve a well-performing prediction model can be large.
-Storing all of it would be both expensive and unnecessary.
-In PEMC, the training data can be produced directly via Monte Carlo simulation, ensuring an effectively unlimited supply.
-Second, this flexibility allows for a more efficient workflow.
-Data can be generated “on the fly” and processed in streaming fashion.
-For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training.
-In this manner, we have successfully trained models in the applications using data on the order of $10^{7}\textendash 10^{8}$.
+In Appendix B, we present numerical results on the performance comparisons of PEMC with standard MC and CV for Asian options under GBM.
+In this case, an excellent CV using the Black-Scholes formula is available under the standard CV framework for us to benchmark against.
+At the same time, in Section 5, we present results on more realistic and technically complex applications, including several production-grade exotic options that are vastly more complex than Asian options.
 4. Analysis and Performance
 In this section, we provide a detailed and comprehensive analysis of the PEMC estimator.
+All omitted proofs are delegated to Appendix A.
 All properties of PEMC discussed here pertain to the evaluation stage, where $\boldsymbol{\theta}$ is arbitrary but fixed.
 Henceforth, we omit the explicit dependence on $\boldsymbol{\theta}$ in the notation (e.g., writing $\boldsymbol{Y}$ in place of $\boldsymbol{Y}(\boldsymbol{\theta})$).
-We begin by establishing the unbiasedness of our estimator.
 4.1. Bias Analysis
@@
 The results are summarized Figure 5 and Table7.
-Further details are left in Appendix E.
+Further details are left in Appendix C.
 Figure 5. Ambulance diversion policies evaluation
@@
 Acknowledgement: We would like to thank [acknowledge individuals, organizations, or institutions] for their valuable input and support throughout this research.
-Appendix
-Appendix A Learning Theory Proofs
+Appendix A Proofs
+A.1. Proof of Lemma 3
+Proof.
+Proof By relaxing the constraints to $n,N\in\mathbb{R}^{+}$, the optimization problem
+$\underset{n>0,N>0}{\text{min}}\quad$ $\frac{1}{n}\sigma_{f-{g}}^{2}+\frac{1}{N}\sigma_{g}^{2}.$
+s.t. $nc_{f-g}+Nc_{g}\leq B,$
+is jointly convex in $n,N\in\mathbb{R}^{+}$.
+The objective value also approaches infinity as $N\to 0$ or $n\to 0$.
+Consequently, the strict convexity ensures the existence of a global minimizer in the interior and directly solving the Lagrangian gives us the result.
+∎
+A.2. Proof of Lemma 4
+Proof.
+Proof Based on the results of Lemma 3, we can deduce that this ratio again does not depend on the budget $B$, so we omit its dependence in the statement.
+Moreover, we can conveniently choose a budget of the form $B=n_{0}c_{f-g}+n_{0}\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{f-g}}{c_{g}}}c_{g}$ which gives an allocation of $n=n_{0}$ and $N=n_{0}\frac{\sigma_{g}}{\sigma_{f-g}}\cdot\sqrt{\frac{c_{f-g}}{c_{g}}}$ for PEMC in accordance with Lemma 3, while comparing with $B/c_{f}$ samples for standard MC.
+The rest follows from Lemma 2 and the assumption $c_{f}=c_{f-g}$.
+∎
+A.3. Proof of Lemma 5
+Proof.
+Proof When $g=\mathbb{E}[f(\boldsymbol{Y})\,|\,\boldsymbol{X}]$ is the true regressor, i.e.,
+$g=\operatorname*{arg\,min}_{h\text{ measurable}}\mathbb{E}[(f(\boldsymbol{Y})-h(\boldsymbol{X}))^{2}],$
+and $f$ is square-integrable, it follows that $\mathbb{E}[(f-g)h]=0$ for all square-integrable $h$.
+Plugging in $h=g$, we obtain $\mathrm{Cov}(f-g,g)=0$.
+Consequently we have $\rho:=\mathrm{corr}(f,g)=\frac{\sigma_{g}}{\sigma_{f}}$ and $\sigma_{f}^{2}=\sigma_{f-g}^{2}+\sigma^{2}_{g}$, which further implies $\sigma_{g}=\rho^{2}\sigma_{f}^{2}$ and $\sigma_{f-g}^{2}=(1-\rho^{2})\sigma^{2}_{f}$.
+The rest follows from $c_{f-g}=c_{f}$.
+∎
+A.4. Proof of Lemma 6
 In this part, we go through the pipeline for a result in the form of Lemma 6 in Section 4.3.1.
@@
 In the following sections, we analyze $\epsilon_{a}^{\mathcal{G}}$ and $\epsilon_{e}^{N_{\text{train}}}$ respectively.
-Appendix B Approximation Error $\epsilon_{a}^{\mathcal{G}}$
+A.4.1. Approximation Error $\epsilon_{a}^{\mathcal{G}}$
 Note that this best-in-class function $g^{*}$ depends on the distribution of $\boldsymbol{\theta}\sim\Theta$, meaning the notion of optimality is distribution-dependent on $\boldsymbol{\theta}$.
@@
 Proof See [107] or [87]. ∎
-Appendix C Statistical Error $\epsilon_{e}^{N_{\text{train}}}$
+A.4.2. Statistical Error $\epsilon_{e}^{N_{\text{train}}}$
 To quantify the capacity of $\mathcal{G}$, one could employ common complexity measures such as the Vapnik-Chervonenkis (VC) dimension [129] or Rademacher complexity [12].
@@
 $\epsilon_{e}^{N_{\text{train}}}\leq\epsilon.$
-Appendix D Proof of Lemma 6
+A.4.3. Proof of Lemma 6
 Lemma 7 and Lemma 8 allows us to control $\epsilon_{\text{total}}$ using (21), which allows us to prove Lemma 6 with the help of the following technical definition and lemma.
@@
 where the constant in $O$ does not depend on $\boldsymbol{\theta}$. This concludes the proof. ∎
-D.1. Variance Reduction of PEMC: An Example on Asian Options
+Appendix B Numerical Results on Asian Option Pricing
 We revisit the Asian option example and conduct experiments using the arithmetic Asian call option under the Geometric Brownian Motion (GBM) model.
@@
 Practical applications and actual use-cases of PEMC will be provided in Section 5.
-D.1.1. Experimental Setup
+B.1. Experimental Setup
 We implement the PEMC procedure detailed in Section 3.5, utilizing a neural network (NN) as the predictive model. The asset price follows the Geometric Brownian Motion (GBM) model:
@@
 The network is implemented using PyTorch [105] and trained using the Adam optimizer [81] with a learning rate of $1\times 10^{-3}$, incorporating dropout layers [125] with a rate of 0.5 after each hidden layer to prevent overfitting.
-D.1.2. Evaluation
+B.2. Evaluation
 For comparison, we implement PEMC, MC and CV method with geometric Asian option
@@
 In practice, efficient MC techniques such as Quasi-Monte Carlo (QMC) methods [33, 2] and their variants—including Langevin Monte Carlo (LMC) [91], Orthogonal Monte Carlo (OMC), and Near-Orthogonal Monte Carlo (NOMC) [90]—can be integrated with PEMC for further variance reduction.
-In particular, when a control variate method is available, such as in our current example, PEMC can be built upon it to potentially achieve even greater variance reduction.
+B.3. Using PEMC as a “Boost" to Known CV
+Lastly, we demonstrate how PEMC can be used to improve existing CV.
+This supports the creative use of PEMC for further efficiency enhancement even for problems that already possess some variance reduction approaches.
+In particular, when a CV with known mean is available, such as in this current example, PEMC can be built upon it to potentially achieve even greater variance reduction.
 To see how, in Algorithm 1, we change the label from $P_{A}(\{S_{t}\}_{t})$ (MC baseline) to $P_{A}(\{S_{t}\}_{t})-P_{G}(\{S_{t}\}_{t})+P_{G}^{\text{exact}}(\boldsymbol{\theta})$ (Geometric CV baseline), and let the model $g$ learn
@@
 In the following section, we go beyond this over-simplified setting and apply PEMC to practical applications in complex derivative pricing.
-Appendix E Ambulance Diversion
-Details can be further found in [41, 45].
-Some key ingredients are 1.
-Patient Arrivals: We simulate patient inflow to each hospital (A or B) according to a nonhomogeneous Poisson process with hourly rates $\lambda_{h,d}$.
+Appendix C Additional Details on the Ambulance Diversion Example
+We provide more details on the ambulance diversion policy evaluation example in Section LABEL:sec:EMS.
+Some key ingredients are as follows: 1) Patient Arrivals: We simulate patient inflow to each hospital (A or B) according to a nonhomogeneous Poisson process with hourly rates $\lambda_{h,d}$.
 Specifically:
@@
 Many of these arrivals (e.g., 25%) come by ambulance, with the remainder as walk-ins.
-2.
-Triage Levels and Service Times: Each arriving patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution:
+2) Triage Levels and Service Times: Each arriving patient is assigned a triage level $\ell\in\{1,2,3,4,5\}$ via a (possibly crisis-adjusted) multinomial distribution:
 $\Pr[L=\ell]=p_{\ell}\quad\text{or}\quad p_{\ell}^{\text{(crisis)}}.$
@@
 $\text{divert}(\tau)=\begin{cases}1,&\text{if }Q>\tau,\\ 0,&\text{otherwise}.\end{cases}$
-4. Mortality Function Each patient $i$ has a time-varying risk of death. The code uses two complementary representations: 1. Direct Mortality Probability:
+4) Mortality Function Each patient $i$ has a time-varying risk of death. The code uses two complementary representations: 1. Direct Mortality Probability:
 […]
 […]
-$\min_{\tau}\;\mathbb{E}\Bigl{[}\sum_{\text{patients }i}\mathbf{1}\bigl{\{}\text{wait}_{i}>\text{death\_time}_{i}\bigr{\}}\Bigr{]},$
+$\min\;\mathbb{E}\Bigl{[}\sum_{\text{patients }i}\mathbf{1}\bigl{\{}\text{wait}_{i}>\text{death\_time}_{i}\bigr{\}}\Bigr{]},$
 subject to feasible travel times and doctor schedules.
@@
 Table 11. Error analysis for mortality estimation with varied thresholds.
+Appendix D Further Implementation Details of PEMC
+We discuss several further implementation details that are important for successfully applying PEMC in practice.
+D.1. Parameter Space $\Theta$
+When determining the parameter space $\boldsymbol{\Theta}$, a key consideration is how frequently the model will be updated.
+To ensure the performance of PEMC, it is desirable that the training data encompass all practical scenarios the model is expected to encounter within the designated update period.
+For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$ and the user of PEMC chooses to update the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses or confidence intervals about the bounds within which these parameters are likely to fluctuate over the coming month.
+This process is conceptually similar to the selection of uncertainty sets in robust optimization [16] or distributionally robust optimization [51], where the goal is to include realizations of parameters with high likelihood.
+However, unlike robust optimization approaches, which often favor data-dependent uncertainty sets, our focus here is also guided by the expertise of financial engineers.
+The training parameter space is designed to reflect realistic and practical scenarios derived from domain knowledge, rather than being strictly driven by statistical guarantees.
+In applications, we construct reasonable $\boldsymbol{\Theta}$ for practical models, such as forward curves in HJM modeling [61] or local stochastic volatility 2D grids [57], among others.
+D.2. Eliminating Data Storage Overheads
+In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively.
+By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage.
+This difference stems from two main considerations.
+First, the volume of training data required to achieve a well-performing prediction model can be large.
+Storing all of it would be both expensive and unnecessary.
+In PEMC, the training data can be produced directly via MC simulation, ensuring an effectively unlimited supply.
+Second, this flexibility allows for a more efficient workflow.
+Data can be generated “on the fly” and processed in streaming fashion.
+For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training.
+In this manner, we have successfully trained models in the applications using data on the order of $10^{7}\textendash 10^{8}$.
+D.3. Evaluation Metrics
+While training the NN estimator $g$ using MSE loss is common, it is not always clear how to interpret the resulting MSE score.
+Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem.
+This ambiguity makes it challenging to determine when the network is sufficiently trained.
+To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$.
+If the network approximates this expectation well, then the sample average of $g(\boldsymbol{X})$ should be close to the sample average of $f(\boldsymbol{Y})$ over a given dataset.
+One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages.
+If $\mathbb{E}[g(\boldsymbol{X})]\approx\mathbb{E}[f(\boldsymbol{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation.
+Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in ML workflows.
+More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard MC.
+Appendix E Additional Details on Hyper-Parameters for Variance Swaps
+Tables 12 and 13 detail the hyper-parameter setups for the NNs presented in Section LABEL:sec:var_swap.
+Table 12. Neural Network Architecture Parameters
+CNN Branch Feed-forward Branch
+kernel size: 3, stride: 1, padding: 1
+max pooling (kernel: 2, stride: 2, padding: 0)
+hidden dim: 512
+output dim: 128
+2D function branch parameters 1D function branch parameters Vector feature branch parameters
+Kernel size: (1, 3) Stride: (1, 3) Padding: 0 AvgPool2d kernel size: (2, 2) AvgPool2d stride: (2, 2) AvgPool2d padding: 0 Kernel size: 10 Stride: 3 Padding: 0 AvgPool1d kernel size: 2 AvgPool1d stride: 2 AvgPool1d padding: 0 Hidden dim: 512 Output dim: 128
+Feed-forward Synthesizer Parameters
+Hidden dim: 128
+Output dim: 1
+Table 13. Hyper-parameter setup for the neural network
+Appendix F Review of HJM
+The HJM model directly describes the evolution of the entire forward rate curve, offering greater flexibility than traditional short-rate models.
+For illustration purposes, we focus on a one-factor specification with exponential volatility structure [61], though the framework readily extends to multi-factor cases.
+Consider the problem of pricing a swaption.
+This is a contract granting its holder the right, but not the obligation, to enter into an interest rate swap at a future date.
+In a standard interest rate swap, one party agrees to pay a fixed rate while receiving a floating rate, and the other party does the opposite.
+Consider a swap with $n_{p}$ fixed payment periods, each of length $\Delta t^{\prime}$, starting at time $t^{\prime}_{0}$ and ending at time $t^{\prime}_{n_{p}}=t^{\prime}_{0}+\sum_{l=1}^{n_{p}}\Delta t^{\prime}$.
+The value of this swap at time $t^{\prime}_{0}$ is:
+$V_{t^{\prime}_{0}}=C\left(R\sum_{l=1}^{n_{p}}B(t^{\prime}_{0},t^{\prime}_{l})\Delta t^{\prime}+B(t^{\prime}_{0},t^{\prime}_{n_{p}})-1\right),$
+where $C$ is the notional amount (contract size), $R$ is the fixed rate, and $B(t^{\prime}_{0},t^{\prime}_{l})$ is the discount factor from $t^{\prime}_{0}$ to $t^{\prime}_{l}$.
+A swaption provides the holder with the option to enter into this swap at $t^{\prime}_{0}$.
+The payoff of the swaption is simply
+$\max(0,V_{t^{\prime}_{0}}),$
+[…]
+$df(t,T)=\mu(t,T)\,dt+\sigma(t,T)^{\top}dW(t),$
+where $\mu(t,T)$ is the drift, $\sigma(t,T)$ is the volatility function of the forward rate, and $W(t)$ is a Brownian motion.
+In contrast to short-rate models (e.g., [130] or [39]), which only model the dynamics of the short-term interest rate, the HJM model directly models the dynamics of the entire term structure of interest rates [66].
+The HJM model is widely used in practice because of its flexibility in modeling interest rate derivatives like swaptions and its ability to incorporate complex volatility structures [29, 5].
+However, the model’s generality also leads to the need for sophisticated numerical methods for simulation [61].
+A key property of the HJM model is the no-arbitrage condition [61], which specifies the drift completely by the volatility:
+[…]
+Thus, in the HJM framework, the model is fully specified by defining the initial forward rate curve $f(0,T)$ and the structure of the volatility $\sigma(t,T)$.
+In our experiment we used a simple one factor HJM for illustration.